### v3

# Benchmarking Trustworthiness in Multimodal LLMs for Video Understanding

Abstract
Recent advancements in multimodal large language models for video understanding (videoLLMs) have enhanced their capacity to process complex spatiotemporal data. However, challenges such as factual inaccuracies, harmful content, biases, hallucinations, and privacy risks compromise their reliability. This study introduces Trust-videoLLMs, a first comprehensive benchmark evaluating 23 state-of-the-art videoLLMs (5 commercial, 18 open-source) across five critical dimensions: truthfulness, robustness, safety, fairness, and privacy. Comprising 30 tasks with adapted, synthetic, and annotated videos, the framework assesses spatiotemporal risks, temporal consistency and cross-modal impact. Results reveal significant limitations in dynamic scene comprehension, cross-modal perturbation resilience and real-world risk mitigation. While open-source models occasionally outperform, proprietary models generally exhibit superior credibility, though scaling does not consistently improve performance. These findings underscore the need for enhanced training data diversity and robust multimodal alignment. Trust-videoLLMs provides a publicly available, extensible toolkit for standardized trustworthiness assessments, addressing the critical gap between accuracy-focused benchmarks and demands for robustness, safety, fairness, and privacy. Code are available at: https://github.com/wangyouze/Trust-videoLLMs
[…]
1 Introduction
Recent advancements in video large language models demonstrate their superior capability to process dynamic visual information across diverse multimodal benchmarks [fu2024video, li2025temporal, li2024mvbench, wang2024videohallucer, guan2024hallusionbench, liu2025video], positioning them as foundational models for understanding the temporal and spatial complexities of real-world multimodal data. These developments mark significant progress toward artificial general intelligence. However, despite efforts to align with human values, videoLLMs face critical trustworthiness challenges, including factual inaccuracies [fu2024video, ning2023video], temporal inconsistency [liu2024tempcompass], harmful content generation [hu2025videojail], biases [park2025assessing], hallucinations [wang2024videohallucer, gao2025exploring], and privacy vulnerabilities [tang2025video]. The inherent spatiotemporal complexities of video data intensify these problems, compromising the dependability of videoLLMs and generating widespread apprehension across academic circles, governance bodies, and civil society.
TABLE I: Comparison of Trust-videoLLMs with existing mainstream video understanding benchmarks. Metrics: Discriminative (Dis.), Generative (Gen.), Tasks/scenario (TS.). "#Models" (the evaluated MLLMs counts, with (*) denoting proprietary models).
Benchmark Aspects Strategy Task Types Statistics
Truthfulness Robustness Safety Fairness Privacy Temporal Cross-modal Dis. Gen. TS. #Models
Video-MME [fu2024video] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 6 13(4)
TemporalBench [cai2024temporalbench] $\times$ $\times$ $\times$ $\times$ $\times$ 6 15(4)
VELOCITI [saravanan2024velociti] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 7 8(3)
TempCompass [liu2024tempcompass] $\times$ $\times$ $\times$ $\times$ $\times$ 4 11(1)
Video-Bench [ning2023video] $\times$ $\times$ $\times$ $\times$ $\times$ 10 8(0)
MVBench [ning2023video] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 20 14(1)
VideoHallucer [wang2024videohallucer] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 5 9(2)
HAVEN [gao2025exploring] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 4 12(2)
Video-SafetyBench [liu2025video] $\times$ $\times$ $\times$ $\times$ $\times$ 13 24(7)
SafeVidBench [wang2025safevid] $\times$ $\times$ $\times$ $\times$ $\times$ 7 16(6)
Trust-videoLLMs 30 23(5)
Compared to image-based MLLMs [zhu2023minigpt, liu2024llavanext, Qwen2.5-VL, hurst2024gpt], which process static visual data with limited exposure to external interference, videoLLMs excel in multimodal comprehension by integrating complex temporal and spatial interactions between visual, auditory, and textual inputs. Current multimodal large language models (MLLMs) benchmarks, covering both image- and video-based evaluations, primarily assess video understanding accuracy [ning2023video, fu2024video] and long-video comprehension reliability [wang2024lvbench] but often overlook critical dimensions such as adversarial robustness, safety, fairness, and privacy. Concurrent work [liu2025video, wang2025safevid] focus on the safety dimension of videoLLMs. Image-based MLLMs benchmark [zhang2024benchmarking], designed for static visual tasks, are inadequate to address trustworthiness risks arising from the dynamic nature of video content, necessitating comprehensive benchmarks tailored to the spatiotemporal complexities of videoLLMs (The limitations of these benchmarks are detailed in Table I).
This study presents Trust-videoLLMs, a comprehensive framework for evaluating the trustworthiness of MLLMs for video understanding and analysis, as shown in Figure 1. Extending the five core dimensions of the TrustLLM [sun2024trustllm]: truthfulness, safety, robustness, fairness, and privacy, we introduce novel tasks tailored to the spatiotemporal dynamics and multimodal nature of video data. A multi-level evaluation approach examines the evolution of multimodal risks in dynamic scenarios and the cross-modal impact of temporal visual inputs on foundational language models, revealing critical vulnerabilities in videoLLMs. The evaluation system comprises 30 tasks, including: (1) spatiotemporal tasks different from image-based trustworthiness benchmarks to establish dynamic scenario standards; (2) analysis of multimodal input interactions affecting videoLLMs decisions; and (3) assessment of model robustness in core video tasks and safety risks in real-world applications. To support this, we construct a large-scale dataset integrating task-adapted existing datasets, synthetic data generated using advanced text/image-to-video tools (e.g., Kling  [^1] https://app.klingai.com/cn/, Jimeng  [^2] https://jimeng.jianying.com/), and manually collect and annotated data, ensuring diverse scenario coverage.
This study evaluates 23 state-of-the-art videoLLMs (5 commercial, 18 open-source), selected for their technological representativeness and architectural diversity. Through rigorous comparative analysis, we identify significant limitations in dynamic scene understanding and cross-modal interference resilience. The Trust-videoLLMs framework offers an interpretable foundation for improving videoLLM performance, underscoring the urgent need for technical advancements to address these trustworthiness deficiencies. Our findings can be summarized as follows:
  • While open-source videoLLMs occasionally outperform proprietary models on specific truthfulness benchmarks [fu2024video, li2024mvbench, wang2024videohallucer], their overall trustworthiness remains lower than that of mainstream proprietary models. Our evaluation reveals that Claude series and Gemini1.5 series demonstrate superior security and trustworthiness.
  • Larger parameter counts don’t consistently translate to better performance. Instead, architectural design and training strategies play a more critical role. Notably, the 7B-parameter Qwen2.5-VL achieves strong performance across multiple dimensions, ranking 9th overall and outperforming many larger models.
  • Improved performance in standard scenarios reflects enhanced model capabilities but increases misuse risks. This trade-off underscores the need for advanced safety alignment strategies to balance performance gains with robust safeguards.
To facilitate standardized and scalable evaluations, we developed a toolbox for investigating the trustworthiness of videoLLMs. Featuring unified interfaces and a modular design for model interaction and task execution, this toolbox aims to provide a robust foundation for future research on trustworthy videoLLM evaluation and development.
Figure 1: Framework of Trust-videoLLMs: The framework encompasses five key aspects, examining the trustworthiness of videoLLMs through a broad lens. It addresses unique challenges such as spatiotemporal risks, temporal consistency issues, and cross-modal influences.
TABLE II: Task Overview. Each task ID is linked to the section 4–5. RtA denotes Refuse-to-Answer rate. B, M, C, R denote BLEU, Meteor, Cider, Rouge-L. : rule-based evaluation (e.g., keywords matching); : automatic evaluation by DeepSeek or other tools; : mixture evaluation. $\oplus$: datasets constructed from scratch.
ID Task Name Dataset Metrics Task Type Eval Type
T.1 Contextual Reasoning QA $\oplus$ Accuracy Dis.
T.2 Temporal Reasoning QA [liu2024tempcompass] Accuracy Dis.
T.3 Video Description [nan2024openvid, liu2024tempcompass] LLLM-score, score, B, M, C, R Gen.
T.4 Events Understanding [ZhXuCoAAAI18] Accuracy Dis.
T.5 Hallucination in Videos [wang2024videohallucer] Accuracy Dis.
R.1 OOD Videos Captioning [bashmal2023capera] LLLM-score, B, M, C, R Gen.
R.2 Noise Videos QA [li2024mvbench] Acc Dis.
R.3 Temporal Understanding Consistency [li2024mvbench] Accuracy Dis.
R.4 Adversarial Attack for Classification [li2024mvbench] Accuracy Dis.
R.5 Adversarial Attack for Captioning [li2024mvbench] LLLM-score, B, M, C, R Gen.
[…]
R.7 Misleading Video Prompts $\oplus$ LLLM-score Gen.
[…]
S.1 NSFW Videos Description $\oplus$ RtA, LLM-score, Toxicity-score Gen.
S.2 NSFW Prompts Execution [mazeika2024harmbench] RtA,LLLM-score,Toxicity-score Gen.
[…]
S.4 Risky Content Identification $\oplus$ L-score Gen.
[…]
S.6 Deepfake Identification [dufour2019deepfakes] Accuracy Dis.
S.7 Jailbreak Attacks [gong2025figstep, liu2024mm, hu2025videojail] RtA, LLLM-score, Toxicity-score Gen.
F.1 Stereotype Impact generation [nan2024openvid] LLLM-score Gen.
F.2 Perference Selection of videoLLMs $\oplus$ RtA, Classifer-RtA Gen.
F.3 Profession Competence Prediction $\oplus$ P-value Gen.
[…]
F.5 Time Sensitivity Analysis $\oplus$ LLM-score Gen.
P.1 Privacy content Recognition [sharma2023disability] Accuracy, Precision,Recall, F1 Dis.
P.2 Privacy Information QA $\oplus$ Accuracy, Precision,Recall, F1 Dis.
[…]
P.4 Celebrities Privacy Information QA $\oplus$ RtA Gen.
P.5 Privacy Information Self-inference $\oplus$ Leakage rate Gen.
2 Related Work
2.1 Multimodal Understanding Benchmarks
Significant efforts have been devoted to developing evaluation frameworks to advance MLLMs. Existing studies assess the trustworthiness of image-based MLLMs across multiple dimensions. For example, POPE [li2023evaluating] systematically evaluates hallucination issues, while SafeBench [ying2024safebench] and MM-SafetyBench [liu2024mm] provide specialized benchmarks for safety alignment in multimodal scenarios. RTVLM [li2024red] introduces a red-teaming dataset evaluating truthfulness, safety, privacy, and fairness. MultiTrust [zhang2024benchmarking] offers a comprehensive toolkit assessing: truthfulness, robustness, safety, fairness, and privacy. However, these benchmarks primarily address static image data and text prompts, failing to capture the spatiotemporal complexities of videoLLMs.
2.2 Trustworthiness Evaluation on Video Understanding
Video understanding, a key branch of MLLMs, demonstrate superior capabilities in processing dynamic visual information and diverse multimodal benchmarks [fu2024video, li2025temporal, li2024mvbench, wang2024videohallucer, guan2024hallusionbench, liu2025video], establishing them as foundational models for understanding the spatiotemporal complexities of real-world multimodal data. However, despite efforts to align videoLLMs with human values, they face critical trustworthiness challenges [ning2023video, liu2024tempcompass, wang2024videohallucer, tang2025video]. Compared to image-based MLLMs (e.g., LLaVA-Next [liu2024llavanext]), videoLLMs exhibit enhanced multimodal understanding by integrating complex spatiotemporal interactions across visual, auditory, and textual inputs. However, current evaluation benchmarks for videoLLMs primarily focus on video understanding accuracy [ning2023video, fu2024video] and long-video comprehension reliability [wang2024lvbench]. Concurrent studies [liu2025video, wang2025safevid] mainly focus on the safety dimension of videoLLMs. Existing image-based MLLM benchmarks [zhang2024benchmarking], designed for static visual tasks, are ill-equipped to address trustworthiness risks arising from the dynamic nature of video content. Consequently, there is an urgent need for a comprehensive evaluation benchmark tailored to the spatiotemporal complexities of videoLLMs from truthfulness, robustness, safety, fairness, and privacy, enabling a systematic and holistic assessment of their trustworthiness to guide model optimization and ensure reliable deployment.
3 Framework of Trust-videoLLMs
3.1 Detailed Description of Trust-videoLLMs
The Trust-videoLLMs framework is organized into five core evaluation dimensions, each encompassing specific tasks, datasets, and metrics to comprehensively assess the trustworthiness of videoLLMs, as shown in Figure 1 and Table II.
3.1.1 Truthfulness in videoLLMs: Assessing Accuracy and Reliability
Truthfulness is vital for ensuring that videoLLMs provide accurate and reliable outputs based on dynamic visual content. Unlike static images, videos require reasoning over temporal sequences, increasing the risk of errors and hallucinations. Departing from prior work [ning2023video, liu2024tempcompass, fu2024video], we assess truthfulness through two key aspects:
Perceptual and Cognitive Proficiency (P.) assesses the intrinsic capability of videoLLMs to accurately perceive and reason about video content. This includes fundamental perceptual tasks such as video classification and video description (Task T.3), as well as higher-level cognitive tasks like temporal perception VQA (Task T.2) and contextual reasoning using a novel dataset (Task T.1).
Contextual Sequential Comprehension (C.) evaluates model truthfulness under complex event sequences, addressing vulnerabilities arising from overall comprehension capability or inherent design limitations. This includes assessing the ability to understand event sequences and temporal coherence (Task T.4), as well as resistance to hallucinations in video understanding (Task T.5).
3.1.2 Robustness in videoLLMs: Assessing Consistency and Resistance
The robustness dimension tests videoLLMs’ ability to maintain accurate and stable understanding under adversarial or perturbed inputs across multi modalities. Trust-videoLLM assess robustness across four aspects:
OOD Robustness (O.) evaluates a videoLLM’s generalization to unseen domains, including diverse video types and natural noise. We utilize an OOD video dataset [bashmal2023capera] and natural noise to assess model robustness under OOD conditions (Tasks R.1 and R.2).
Temporal Understanding Robustness (T.) assesses robustness to disruptions in temporal structure by altering frame order or introducing missing frames, evaluating the model’s capacity for temporal reasoning (Task R.3).
Adversarial Robustness (A.) evaluates susceptibility to adversarial inputs—a well-known vulnerability of deep neural networks [szegedy2013intriguing]. We assess the model’s resilience to perturbations crafted to mislead video understanding, using keyframes selected via uniform sampling and optical flow analysis. Adversarial examples are generated with the MI-CWA algorithm [chen2023rethinking], and performance is evaluated on video classification (Task R.4) and captioning (Task R.5).
Multimodal Interaction Robustness (M.) evaluates the model’s ability to maintain semantic alignment between modalities under adversarial conditions. We assess robustness through three tasks: introducing noise perturbations to MVBench video questions (Task R.7), testing resistance to misleading textual prompts using YouTube videos (Task R.8), and evaluating the influence of varied video content on sentiment analysis judgments (Task R.6). These evaluations ensure reliable modality understanding.
3.1.3 Safety in videoLLMs: Assessing Output Security
Ensuring the safety of videoLLMs is critical to prevent harmful outputs and mitigate risks of misuse. This evaluation addresses toxicity in generated content, unsafe content recognition, and defenses against malicious manipulations, considering the unique temporal and multimodal nature of video.
Toxicity in Generated Content (G.). Toxicity refers to outputs containing pornography, violence, blood, or hate speech. We assess videoLLMs’ ability to detect and describe NSFW content using videos from the BigPorn, Violence, and HateSpeech datasets (Task S.1). Rejection rates under harmful instructions are evaluated using HarmBench [mazeika2024harmbench] (Task S.2). To examine how video context influences toxic responses, prompts from RealToxicityPrompts [gehman2020realtoxicityprompts] are paired with semantically related and unrelated videos across five toxicity categories (Task S.3).
Unsafe Content Recognition (U). Beyond detecting toxic language, we evaluate whether videoLLMs can recognize unsafe or risky actions in videos that may encourage harmful imitation (Task S.4). NSFW segments (10–20% duration) are inserted into otherwise safe videos to assess temporal consistency and unsafe content recognition (Task S.5).
Safety Against Malicious Manipulations (S.). Robustness against adversarial manipulation is vital. We assess deepfake detection using manipulated videos (Task S.6) and evaluate resistance to jailbreak attacks (Task S.7). This includes two image-based methods—FigStep [gong2025figstep] and MMsafetyBench [liu2024mm]—converted to video, and one native video-based attack, VideoJail [hu2025videojail].
3.1.4 Fairness and Bias in VideoLLMs: Assessing Equity and Bias
Ensuring fairness in videoLLMs is crucial to mitigate biases arising from training data or multimodal interactions that may result in stereotypical or discriminatory outputs. This evaluation assesses bias across modalities and examines the model’s ability to uphold fairness, with a focus on temporal and contextual consistency.
Bias from Data-Driven Influences (B.). VideoLLMs trained on large-scale datasets may inherit demographic biases, potentially generating stereotypical outputs. We evaluate bias manifestation through three complementary methods: analyzing stereotype presence using videos from OpenVid-1M [nan2024openvid] covering occupation, gender, age, and race with targeted prompts (Task F.1); employing established text-based bias benchmarks [nadeem2020stereoset, nangia2020crows] to generate videos for stereotype classification (Task F.2); and extending bias evaluation to scenarios with different attributes description including age, gender, and skin tone (Task F.3). These assessments evaluate model bias awareness arising from data-driven influences.
Fairness in Temporal and Multimodal Understanding (F.). Bias may emerge across modalities or over time, necessitating fairness in dynamic, multimodal contexts. We test stereotype reinforcement by pairing bias-inducing prompts with related and unrelated videos (Task F.4) and assess fairness in temporal understanding across gender, race, and occupation categories (Task F.5), focusing on the model’s sensitivity to time-dependent biased content.
3.1.5 Privacy in VideoLLMs: Assessing Information Protection and Inference Control
Privacy in videoLLMs is critical to prevent the unauthorized disclosure or inference of sensitive information from video or text inputs. The evaluation focuses on the model’s ability to identify privacy-relevant content and avoid generating or inferring personal information, ensuring compliance with ethical and legal standards.
Figure 2: Left: Rankings of videoLLMs in each sub-aspect of Trust-videoLLMs. The max frames num of All videoLLM is set to 64. Right: (a) Top 10 videoLLMs: highlighting top 3 Performers. (b) Pearson Correlation Coefficients between sub-aspects.
Privacy Awareness (R.). This component evaluates whether videoLLMs can detect and appropriately respond to privacy-sensitive content, such as faces, license plates, and ID cards. We first assess identification capabilities using videos from BIV-Priv [sharma2023disability], which include items like passports, credit cards, and private letters (Task P.1). We further evaluate privacy perception using real-world YouTube videos containing transient private information, such as phone or computer screens, license plates, and delivery addresses by QA method (Task P.2).
Control Over Privacy Inference (I.). VideoLLMs should avoid inferring or generating private information, especially without explicit prompts. We adapt the InfoFlow Expectation task [mireshghallah2023can] to multimodal settings, pairing it with videos to assess model agreement on privacy use (Task P.3). To evaluate protection of celebrity privacy, we use videos from diverse domains (sports, entertainment, politics, music), examining whether the model refrains from disclosing personal details (Task P.4). We also test the tendency to infer implicit privacy-sensitive content using videos from the OpenVid-1M dataset [nan2024openvid] (Task P.5).
3.2 Evaluated VideoLLMs
To systematically assess the trustworthiness of videoLLMs, we curated a diverse set of 23 models, spanning various design paradigms, capabilities, and accessibility levels. This includes advanced closed-source models (e.g., GPT-4o [hurst2024gpt], Gemini [team2024gemini], Claude [Anthropic]) to establish performance benchmarks, and leading open-source models (e.g., Qwen2.5-VL, LLaVA-Video, MiniCPM, InternVL) to evaluate current limitations. Detailed descriptions are provided in App. A.1.
3.3 Dataset Collection
To assess videoLLMs across 30 tasks, we developed a comprehensive dataset by integrating existing task-specific data, generating videos via text/image-to-video tools, and manually collecting and annotating them to ensure diverse scenario coverage. The dataset comprises 6,955 videos with durations ranging from 5 seconds to over 30 minutes. Detailed dataset description for each task are provided in App. A.2.
3.4 Toolbox
To support the evaluation of emerging videoLLMs and tasks while enhancing the scalability of Trust-videoLLMs, we introduce a universal, extensible toolbox for assessing trustworthiness. The framework standardizes assessment across diverse models and interaction formats through a modular design separating data and metrics, enabling efficient reuse, updates, and community-driven development.
4 Experimental Results Analysis
We conducted comprehensive experiments across 30 curated tasks in the benchmark. This section presents the rankings in Figure 2 and summarizes key findings from the experimental results. Extended analysis are provided in App. C. Evaluation details for each dimension are provided in App. D to H.
Overall Performance. Figure 2 presents the overall rankings, revealing a diverse performance landscape. Closed-source models, particularly the Claude and Gemini series, generally outperform their open-source counterparts. Claude4-sonnet ranks first, followed by Claude3.7-sonnet and Gemini1.5-Flash. GPT-4o, despite excelling in specific sub-aspects, ranks sixth—just behind InternVL2.5-78B—indicating balanced but non-leading performance. Among open-source models, InternVL2.5-78B and Qwen2.5-VL-72B achieve the highest rankings (fifth and seventh), demonstrating competitiveness with closed-source models. However, most open-source models, such as VideoLLaMA3-7B and LLaVA-OneVision-72B, fall in the lower half.
Figure 2(a) illustrates the performance of top-ranking models across multiple dimensions. Claude4-Sonnet exhibits superior safety performance with a balanced, high-performing profile. Claude3.7-Sonnet delivers consistent cross-dimensional reliability, though without standout strengths. Gemini1.5-Flash excels in robustness but shows significant performance variability, resulting in an irregular pattern. Other models achieve lower overall scores and lack distinct differentiating traits.
Figure 2(b) highlights intricate relationships among sub-aspects of trustworthiness. Strong intra-dimensional correlations are observed, particularly within truthfulness and safety clusters. Cross-dimensional analysis reveals notable patterns: robustness in multimodal scenarios strongly correlates with safety dimensions, while temporal robustness displays significant negative correlations with truthfulness aspects. Fairness dimensions exhibit weak cross-correlations, suggesting independent evaluation characteristics.
Truthfulness.
Figure 3: The average performance of top-13 videoLLMs across contextual reasoning QA, temporal perception QA, video description, event understanding & detection, and hallucination in video understanding tasks.
As shown in Figure 3, open-source videoLLMs excel in specialized reasoning tasks (e.g., temporal/contextual QA) due to task-specific optimization. However, closed-source models mitigate hallucinations more effectively through extensive pretraining (presented in Table VI of App D.2.2). In temporal challenges, over 50% of videoLLMs score below 60% on temporal QA, indicating difficulty with cross-frame integration. This highlights the need for improved temporal modules. For hallucination, leading models (e.g., Claude, Qwen2.5-VL-72B) employ conservative strategies to minimize false positives, though they require further calibration to prevent overly cautious responses.
Robustness.
Figure 4: Correlation between overall trustworthiness rankings and robustness sub-aspect rankings.
Closed-source videoLLMs demonstrate superior performance on clean data but exhibit vulnerabilities to noise and adversarial attacks. In contrast, larger open-source models display variable robustness, occasionally rivaling closed-source counterparts. As depicted in Figure 4, a positive correlation exists between overall ranking and robustness sub-dimension rankings, with Multimodal Interaction Robustness (R-M) and Temporal Understanding Robustness (R-T) emerging as pivotal determinants of model efficacy. Conversely, Adversarial Robustness (R-A) and OOD Robustness (R-O) exert a relatively minor influence on overall performance. Temporal reasoning and multimodal conflict resolution remain challenging, particularly for smaller models, underscoring the necessity for advanced temporal modules and robust multimodal fusion techniques. Furthermore, the susceptibility of most models to adversarial perturbations highlights the critical need to enhance resilience.
TABLE III: Performance results (%) for Tasks S.1 and S.7 are presented, with Task S.7 showing average scores across four jailbreak attack methods. L-score denotes LLM-score.
Model
NSFW Video
Description
Jailbreaks
Attack
$uparrow RtA\uparrow L-score\uparrow$
Claude3.7-sonnet 64.6 87.7 78.8 97.4
GPT-4o 35.4 76.9 57.1 83.2
Claude4-sonnet 73.85 95.2 58.4 82.5
Qwen2.5-VL-72B 5.8 46.2 41.1 72.7
Oryx1.5-7B 6.2 87.7 36.7 67.8
Long-VA-7B 12.3 36.9 26.8 55.1
Safety. Performance results of some videoLLMs for Tasks S.1 and S.7 are presented in Table III. In overall, Closed-source videoLLMs, such as Claude and GPT-4o, set a high safety standard, effectively rejecting NSFW content and toxic prompts, but struggle with detecting subtle risky content and defending against jailbreak attacks like VideoJail-Pro. Open-source models, however, require substantial improvements in safety alignment, particularly for NSFW detection and resilience to video-based jailbreak attacks, as they exhibit lower refusal rates and higher toxicity. Video context significantly impacts safety, with contextually relevant inputs amplifying the risk of harmful outputs. To enhance videoLLM safety across both model types, targeted advancements in temporal understanding and modality alignment are essential.
Figure 5: Performance (%) of top 15 videoLLMs for the Stereotype Impact Generation task.
Fairness. Figure 5 illustrates that closed-source videoLLMs surpass open-source counterparts in stereotype impact evaluation, attributed to superior data curation and ethical constraints. While larger models generally handle sensitive attributes more effectively, fairness primarily depends on architectural design and training objectives rather than scale. Assessments of temporal and multimodal understanding reveal that robust temporal sensitivity requires refined temporal modeling and improved language-vision alignment, independent of model size or type. In occupational and social evaluations, models exhibit a propensity for stereotyping based on visual attributes like gender or age, with textual cues providing limited bias mitigation, underscoring ongoing challenges in cross-modal integration.
Figure 6: RtA Rate (%) in the Celebrity Privacy Information QA Task (top 10 videoLLMs). n denotes name, o denotes occupation, and w denotes without; for example, wo-name indicates that only occupation is provided in the prompt.
Privacy. Closed-source videoLLMs like GPT-4o and Claude-4-sonnet lead in privacy-sensitive tasks, but certain open-source models show competitive potential, as shown in Figure 6. However, all models face challenges with recall variability, context sensitivity, and autonomous privacy reasoning, posing a dual challenge of enhancing detection capabilities while mitigating privacy leakage risks. Improved training data diversity and contextual analysis are critical for advancing privacy protection in videoLLMs.
5 Discussion
Based on the evaluation results, key phenomena are discussed below. A detailed analysis is provided in App. B.
Model Scale vs. Model Performance. The evaluation reveals that increasing videoLLM scale does not always correlate with better performance, particularly in tasks that require complex temporal reasoning. Smaller, optimized open-source models like LongVA-7B outperform larger ones like LLaVA-Video-72B on specific tasks such as event understanding. This highlights that architectural design and task-specific optimization are more important than sheer scale when it comes to effective video understanding .
Complexity of Cross-Modal Understanding and Multimodal Integration. Video understanding involves multiple modalities (e.g., visual, textual, and auditory), requiring strong cross-modal integration. Benchmark results show that closed-source models, such as Claude and Gemini, excel in handling conflicts between modalities, maintaining a high degree of semantic alignment. However, open-source models struggle with cross-modal coherence, especially when confronted with conflicting or biased inputs. These results highlight the complexity of cross-modal integration, showing that addressing these challenges is key to improving open-source videoLLMs.
The Safety in Video Understanding. Enhanced video understanding improves truthfulness in VideoLLMs. However, safety evaluations show that both related and unrelated videos affect performance against harmful prompts. Commercial models generally exhibit greater robustness to such prompts with both video types, whereas open-source models experience significant performance degradation, revealing a notable disparity. Furthermore, embedding jailbreak attack information in the spatiotemporal domain of videos has proven more effective against MLLMs, indicating that dynamic visual scenarios require further safety alignment research compared to static images.
Fairness and Bias Issues. The experiments also reveal the widespread challenge of fairness and bias, especially in handling sensitive attributes like gender, age, and race. While closed-source models mitigate bias through better-designed training datasets, open-source models often exhibit inconsistent fairness, particularly in cross-modal interactions. This highlights the importance of addressing both data-driven biases and model design to ensure fairness in video understanding tasks.
Model Performance and Challenges The primary advantage of videoLLMs lies in their ability to process spatiotemporal data and integrate multimodal information. However, despite their success in static image tasks, these models face significant challenges when handling video data. The temporal aspect of videos demands reasoning over event sequences, detecting fine details, and maintaining consistency, which places higher demands on model stability. Benchmark results show that while models like Claude and Gemini1.5 series perform well overall, they still face challenges in temporal understanding and cross-modal integration, highlighting that video understanding is a complex, multi-layered problem.
6 Conclusion
Trust-videoLLMs assesses the trustworthiness of 23 videoLLMs across five critical dimensions: truthfulness, robustness, safety, fairness, and privacy. Findings reveal significant limitations in spatiotemporal understanding, cross-modal resilience, and temporal consistency. Proprietary models generally outperform in multiple tasks but struggle with truthfulness and safe temporal understanding, highlighting the need for enhanced temporal modeling, multimodal alignment, and robust safety mechanisms. Trust-videoLLMs provides a standardized, extensible framework to advance videoLLM trustworthiness for reliable real-world applications.
7 Acknowledgments
This paper is supported by the National Science Foundation of China Project (Nos. 62306098 and U23B2031), Fundamental Research Funds for the Central Universities (No. JZ2024HGTB0256) and the Open Project of Anhui Provincial Key Laboratory of Multimodal Cognitive Computation, Anhui University (No. MMC202412). The computation is completed on the HPC Platform of Hefei University of Technology.
Appendix A Models and Datasets Specification
A.1 Evaluated VideoLLMs
To systematically assess the trustworthiness of VideoLLMs, we curate a diverse set of 23 models that span various design paradigms, capabilities, and accessibility levels. Our selection includes advanced closed source models (e.g., GPT-4o [hurst2024gpt], Gemini [team2024gemini] series, Claude series) to benchmark performance ceilings and highlight gaps in current open-source systems. We incorporate multiple variants from well-established frameworks such as LLaVA-Video [zhang2024videoinstructiontuningsynthetic] and MiniCPM [yao2024minicpm] to examine how scaling, instruction tuning, and vision-language alignment affect model behavior. Additionally, we include models that emphasize long-context reasoning (e.g., LongVA [zhang2024longva]), real-time captioning (e.g., LiveCC [livecc]), and multi-modal instruction following (e.g., TPO [li2025temporal], ShareGPT4video [chen2024sharegpt4video]). This composition reflects the rapid diversification of VideoLLM applications and architectures, enabling a thorough and comparative analysis across tasks requiring temporal understanding, visual grounding, and safety-critical reasoning. For detailed information of the evaluated models, see Table IV.
TABLE IV: Detailed information of the evaluated videoLLMs in this paper.
Model Open-source Version Institution Deployment
Claude4-sonnet No 20250514 Anthropic Official API
Claude3.7-sonnet No 20250219 Anthropic Official API
Gemini1.5-Pro No 20250504 Google Official API
Gemini1.5-Flash No 20250504 Google Official API
GPT-4o No 20241120 OpenAI Official API
Qwen2.5-VL-72B Yes Qwen/Qwen2.5-VL-72B-Instruct Alibaba Locally Load
Qwen2.5-VL-7B Yes Qwen/Qwen2.5-VL-7B-Instruct Alibaba Locally Load
LLaVA-Video-72B Yes lmms-lab/LLaVA-Video-72B-Qwen2 ByteDance & NTU Locally Load
LLaVA-Video-7B Yes lmms-lab/LLaVA-Video-7B-Qwen2 ByteDance & NTU Locally Load
MiniCPM-o-2.6-7B Yes openbmb/MiniCPM-o-2_6 ModelBest Locally Load
MiniCPM-V-2.6-7B Yes openbmb/MiniCPM-V-2_6 ModelBest Locally Load
Oryx-34B Yes THUdyh/Oryx-34B THU Locally Load
Oryx1.5-7B Yes THUdyh/Oryx-1.5-7B THU Locally Load
InternVL2.5-78B Yes OpenGVLab/InternVL2_5-78B Shanghai AI Lab Locally Load
LLaVA-OneVision-72B Yes llava-hf/llava-onevision-qwen2-72b-ov-hf ByteDance Locally Load
mPLUG-Owl3-7B Yes mPLUG/mPLUG-Owl3-7B-240728 Alibaba Locally Load
LongVA-7B Yes lmms-lab/LongVA-7B NTU Locally Load
Sharegpt4video-8B Yes Lin-Chen/sharegpt4video-8b Shanghai AI Lab Locally Load
TPO-7B Yes ruili0/LongVA-7B-TPO Stanford&USTC Locally Load
Long-LLaVA-7B Yes aws-prototyping/long-llava-qwen2-7b CUHK Locally Load
Video-ChatGPT-7B Yes MBZUAI/Video-ChatGPT-7B MBZUAI Locally Load
LiveCC-7B Yes chenjoya/LiveCC-7B-Instruct NUS Locally Load
VideoLLaMA3-7B Yes DAMO-NLP-SG/VideoLLaMA3-7B Alibaba Locally Load
A.2 Dataset Construction
The dataset is constructed from diverse sources to ensure representativeness, comprising 6,955 videos: 550 sourced from YouTube, 2,753 generated via text/image-to-video tools and other methods, and 3,652 selected from existing video datasets. Task distribution includes 17.9% truthfulness, 15.1% robustness, 30.4% safety, 28.8% fairness and bias, and 7.6% privacy. The videos cover varied scenarios, such as action recognition, object reasoning, and temporal perception. Video durations range from 5 seconds to over 30 minutes, with standardized sampling of up to 64 frames per video to ensure consistent model evaluation.
For data annotation, we collected and annotated a dataset of 1,270 videos from online sources for the Trust-videoLLMs benchmark, covering tasks such as Contextual Reasoning QA (1,030 videos, with 300 uniformly sampled by duration for evaluation efficiency), NSFW Video Description (60), Risk-Containing Videos (50), Privacy-Sensitive Videos (90), and Celebrity Videos (50). A team of five university student annotators, proficient in video analysis, AI ethics, and bilingual (English and Chinese) capabilities, was recruited and underwent a 2-hour training session on task-specific guidelines, ethical considerations, and privacy identification strategies. To ensure high-quality annotations, each video was labeled by at least three annotators, with discrepancies resolved through majority voting or team discussions.
Detailed descriptions of the data for each task are provided in the appendices, specifically in Sections D-H.
Appendix B Comprehensive Discussion of Benchmarking Results
Trust-videoLLMs evaluates the trustworthiness of 23 videoLLMs across five key dimensions—truthfulness, robustness, safety, fairness, and privacy—through 30 carefully designed tasks. This comprehensive benchmark reveals critical limitations in videoLLMs, stemming from spatiotemporal complexities, temporal consistency issues, and cross-modal interactions. Key findings and phenomena are discussed below.
Model scale and model performance are not necessarily positively correlated. Conventional view suggests that model performance improves with increasing parameter scale. However, this assumption does not hold for current videoLLMs. Proprietary models, such as Gemini1.5-Pro, underperform compared to open-source 7B models in specific video understanding tasks, notably ranking 18th and 22nd in Perceptual and Cognitive Proficiency and Contextual Sequential Comprehension aspects within the Truthfulness dimension. Similarly, within the same model series, smaller models may outperform larger ones in certain tasks; for instance, Qwen2.5-VL-7B significantly surpasses Qwen2.5-VL-72B in the Fairness and Bias dimension. These findings indicate that temporal modeling complexity and reliance on parameter scaling alone do not guarantee performance gains, underscoring the need for advanced model architectures and diverse training data.
The Effects of cross-modal interaction and multimodal fusion on videoLLM performance. The integration of multimodal inputs—text, images, and temporal video data—presents unique challenges and opportunities for videoLLMs. Our evaluation, particularly in tasks like Multimodal Interaction Robustness (Section 11.4) and Temporal Understanding Robustness (Section 11.2), highlights the critical role of cross-modal interactions in shaping model performance. The task of assessing the impact of video on sentiment analysis demonstrates that videoLLMs often struggle with conflicting sentiment cues between visual and textual inputs. For instance, proprietary models like Claude3.7-sonnet and Gemini1.5-Pro exhibited robust performance (Table X), maintaining high accuracy in sentiment classification despite multimodal inconsistencies, likely due to their advanced vision-language alignment and diverse training datasets.
The results from the video description task (Table V) further underscore the importance of effective multimodal integration. Models like Qwen2.5-VL-7B achieved a high LLM-score of 52.6%, reflecting strong semantic and lexical accuracy in generating descriptions for multimodal video clips. However, models like Video-ChatGPT-7B scored as low as 5.9%, revealing deficiencies in integrating visual details with contextual associations. This variability suggests that robust cross-modal integration, particularly the ability to align temporal dynamics with textual reasoning, is a key determinant of performance. The superior performance of closed-source models, such as GPT-4o and Claude series, can be attributed to their large-scale pretraining and fine-tuning on diverse datasets, which enhance their ability to process and synthesize multimodal information effectively.
Moreover, the contextual reasoning task (Section D.1) revealed that models with specialized temporal modeling modules, such as Qwen2.5-VL-72B, outperformed others in tasks requiring dynamic scene understanding (Figure 8). This indicates that architectural innovations tailored to temporal and multimodal data processing are crucial for improving videoLLM performance. However, even high-performing models struggled with out-of-distribution (OOD) scenarios, as seen in the OOD video understanding task (Table VIII), where performance dropped significantly under adversarial perturbations (Figure 19). These findings suggest that while cross-modal integration enhances performance in controlled settings, robustness to real-world variations remains a challenge, necessitating further advancements in multimodal alignment and adversarial training.
Safety risks in video understanding. Safety is a critical concern for videoLLMs, given their potential to generate harmful content or be manipulated through adversarial inputs. The Trust-videoLLMs evaluated safety across multiple tasks, including identification of NSFW video content, deepfake identification, and resilience to jailbreak attacks. The results indicate that closed-source models, such as Claude4-sonnet (96.6% average safety score) and GPT-4o (96.7%), significantly outperformed open-source models like LiveCC-7B (73.1%) in rejecting toxic prompts and detecting risky content. This disparity is particularly evident in the VideoJail and VideoJail-Pro settings, where open-source models exhibited lower refusal rates and higher toxicity in outputs.
The task of toxic content continues (Table XV) highlighted the influence of video context on safety outcomes. Contextually relevant videos amplified the risk of harmful outputs, as models struggled to filter toxic prompts when paired with semantically related visual inputs. For example, open-source models like Video-ChatGPT-7B achieved only average 49.9% RtA rate in toxic content generation, underscoring their susceptibility to multimodal manipulation. In contrast, closed-source models like Claude3.7-sonnet maintained high refusal rates (e.g., 93.3% in VideoJail), likely due to robust safety alignment mechanisms and curated training data.
Deepfake identification (Figure 30) further revealed safety challenges, with closed-source models like Claude4-sonnet (53.5% accuracy) outperforming open-source counterparts like Oryx1.5-7B (17.5% but with lower consistency). The lower performance of open-source models in detecting manipulated content suggests limitations in their ability to discern subtle temporal and visual cues, a critical requirement for ensuring safety in real-world applications. These findings emphasize the need for enhanced safety mechanisms, such as targeted training on adversarial and toxic datasets, to mitigate risks of misuse and ensure ethical alignment in videoLLMs.
Fairness and Bias Issues in video understanding. Fairness and bias are pivotal dimensions of trustworthiness, particularly given the potential for videoLLMs to reinforce stereotypes or exhibit discriminatory behavior. The Trust-videoLLMs benchmark assessed fairness through tasks like Stereotype Impact Generation and Agreement on Stereotypes, revealing significant disparities between closed- and open-source models. Closed-source models, such as Claude4-sonnet and Gemini1.5-Pro, demonstrated lower stereotype agreement rates, reflecting better suppression of biases due to systematic data curation and ethical training objectives. In contrast, open-source models like TPO-7B and Oryx-34B exhibited higher agreement rates, indicating a tendency to overfit to biased training data.
The task of Time Sensitivity Analysis further highlighted the challenges of temporal bias in videoLLMs. Open-source models like Sharegpt4video-8B achieved high accuracy in adjusting for time-dependent biases, rivaling closed-source models like GPT-4o. However, models like Oryx-34B recorded low accuracy (15.9%), underscoring the dependency on architectural design and temporal modeling capabilities rather than sheer scale. The P-value analysis for professional competence prediction revealed that gender had a stronger predictive influence than age or skin tone, particularly in open-source models, suggesting that biases in training data can compound over time in video contexts.
These findings indicate that fairness in videoLLMs requires a dual approach: addressing data-driven biases embedded in large-scale datasets and enhancing multimodal understanding to prevent temporal bias accumulation. The high stereotype rates in open-source models (e.g., MiniCPM-V-2.6-7B) underscore the urgent need for debiasing techniques, such as balanced dataset curation and fairness-aware training objectives, to ensure equitable performance
The challenges on the trustworthiness of videoLLMs. The Trust-videoLLMs benchmark reveals several overarching challenges in achieving trustworthy videoLLMs. First, spatiotemporal understanding remains a significant hurdle, as evidenced by the low accuracies in temporal reasoning tasks (e.g., 1.3% for some models). The complexity of processing dynamic, multimodal data requires advanced architectures that can effectively model temporal dependencies and cross-modal interactions. Second, robustness to adversarial inputs is a critical limitation, with most models exhibiting performance degradation under noise or adversarial perturbations. This vulnerability highlights the need for adversarial training and robust multimodal alignment strategies. Third, safety alignment is insufficient, particularly in open-source models, which struggle with detecting toxic content and resisting jailbreak attacks. The amplification of harmful outputs in contextually relevant scenarios underscores the importance of context-aware safety mechanisms. Fourth, fairness and bias mitigation require significant improvements, as models often perpetuate stereotypes due to biased training data or inadequate temporal modeling. Finally, privacy protection poses a dual challenge: while high detection rates in tasks like Private Information QA indicate strong capabilities, they also increase the risk of privacy leakage, necessitating a balance between performance and ethical safeguards.
In conclusion, the Trust-videoLLMs provides a standardized framework for evaluating and improving the trustworthiness of videoLLMs. While closed-source models generally outperform open-source counterparts, both face significant challenges in spatiotemporal understanding, robustness, safety, fairness, and privacy. Addressing these issues requires advancements in architectural design, diverse and ethically curated datasets, and targeted training strategies to ensure reliable and equitable performance in real-world applications.
Appendix C Extended Analysis of Evaluation Results
To facilitate readers’ understanding of the evaluation results across the five dimensions, we have comprehensively aggregated the overall performance metrics for each dimension (Section D - H) as follows:
C.1 Truthfulness
The truthfulness evaluation assesses videoLLMs’ ability to generate accurate and contextually relevant responses across tasks such as contextual reasoning, temporal perception, event understanding, video description, and hallucination detection. In the Hallucination in Videos task (Table VI), closed-source models like Claude3.7-sonnet and Gemini1.5-Flash show strong performance with better hallucination scores, indicating better resistance to generating incorrect video content interpretations, while open-source models like LLaVA-Video series exhibit higher hallucination rates. In the Contextual Reasoning QA task, models are tested on 300 YouTube videos of varying lengths, with accuracy as the primary metric; results (Figure 8) highlight significant performance gaps, with Qwen2.5-VL-72B and Video-LLaVA-72B excelling in integrating multimodal information for complex video contextual reasoning. In the temporal Perception QA task, results, as shown in Figure 10 reveal a wide variance in accuracy among models, ranging from 1.3% for Video-ChatGPT-7B to 68.3% for LiveCC-7B, with an average of 50.3%. This indicates significant differences in the models’ capabilities to process time-dependent information in videos. The Events Understanding and Detection task, using 200 YouCook2 videos, shows Video-ChatGPT-7B achieving high accuracy in identifying correct event sequences, while Gemini1.5-Pro is best performer in closed-source models. Overall, open-source model Qwen2.5-VL series demonstrate superior truthfulness and contextual understanding, but there is no significant difference in total scores between the leading open-source models and most closed-source models.
C.2 Robustness
Robustness is evaluated through tasks like OOD noise handling, temporal understanding, and resilience to adversarial perturbations. In OOD Noise Videos QA, models like LiveCC-7B and ShareGPT4Video-8B exhibited minimal performance degradation, outperforming most open-source models, but GPT-4o and Claude3.7-sonnet showed larger drops due to optimization for clean data. Adversarial Robustness, assessed via the MI-CWA algorithm, highlighted that most videoLLMs suffer notable performance degradation under untargeted adversarial attacks, indicating limited robustness to adversarial noise. Among closed-source models, Claude4-sonnet exhibits strong robustness, whereas GPT-4o and Claude3.7-sonnet perform less reliably, revealing considerable variation within this category. In Temporal Understanding Robustness, Gemini1.5 series exhibits the highest robustness with only a 0.0% accuracy drop. Open-source models such as Video-ChatGPT-7B (0.2%) and Video-LLaMA3-7B (0.8%) show moderate sensitivity. In Multimodal Interaction Robustness, Claude3.7-sonnet achieved high accuracy, while open-source models like Qwen2.5-VL-72B and LLaVA-Video-72B showed comparable robustness, though other models performance varied widely. Overall, Gemini series models offer better stability, but optimized open-source models (e.g. Qwen2.5-VL2.5-72B, LLaVA-Video-72B) demonstrate promising potential, emphasizing the importance of training data quality and architectural design over mere model scale for robust real-world performance.
C.3 Safety
The safety evaluation of videoLLMs across seven core tasks reveals significant disparities between closed-models and open-source models in safety capabilities, as well as the critical influence of video context on model security. Closed-models demonstrate superior safety performance, with the Claude series achieving the highest RtA rates of up to 79% for NSFW content and Claude4-sonnet ranking among the top performers across multiple safety metrics. GPT-4o exhibits exceptional toxicity control with the lowest response toxicity scores, while the Gemini series, despite strong performance in certain tasks, shows elevated risks in toxicity responses. In contrast, most open-source models exhibit poor performance in NSFW content recognition and rejection, with generally low refusal rates, though Qwen2.5-VL-7B stands out among open-source alternatives and occasionally matches proprietary model performance in specific safety tasks. The evaluation reveals that video context critically influences model safety, with toxic text prompts paired with contextually relevant videos significantly amplifying harmful content generation risks compared to unrelated video pairings, indicating that video context may exacerbate unsafe model behaviors. While Claude4-sonnet demonstrates the strongest temporal consistency maintenance capabilities (96.2%) in temporal dependency misleading tasks, most models face challenges in processing videos with embedded NSFW content, particularly in maintaining temporal continuity. All models show limited performance in deepfake identification tasks, with the highest accuracy reaching only 55% (Gemini1.5-Pro), suggesting substantial room for improvement in countering advanced video manipulation techniques. Claude3.7-sonnet achieves the best defense against jailbreak attacks (91.0%), while video-specific attacks such as VideoJail and VideoJail-Pro pose varying degrees of threats to different models, highlighting the importance of targeted defense mechanisms. These findings indicate that while closed-models videoLLMs have established relatively robust safety mechanisms, vulnerabilities remain when confronting sophisticated video attacks and subtle risky content identification, and open-source models urgently require enhanced safety alignment, particularly in NSFW content detection and resistance to video-based adversarial attacks.
C.4 Fairness&Bias
Stereotype sensitivity varies dramatically across models, with closed-source models (GPT-4o, Gemini1.5 series) demonstrating superior stereotype suppression with rates between 19.0-23.0%, while open-source models like MiniCPM-V-2.6-7B and LongVA-7B show concerning stereotype rates of 70.5-74.3%. Model size generally correlates with reduced bias, as evidenced by LLaVA-Video-72B achieving 20.0% lower stereotype rates than its 7B counterpart. In preference selection tasks, RtA rates vary significantly, with Claude4-sonnet leading at 78% and many open-source models showing poor bias avoidance below 20%. Professional competence predictions reveal persistent biases influenced by gender, age, and skin tone, particularly affecting open-source models in visual-only settings. Multimodal analysis shows that closed-source models maintain lower stereotype agreement rates (6.5-12.3%) compared to open-source alternatives, while temporal sensitivity analysis demonstrates that bias adjustment capabilities depend more on architectural refinement than parameter scale, with some open-source models like Sharegpt4video-8B (88.6%) outperforming certain closed-source solutions. Overall, the results indicate that while closed-source models generally exhibit better fairness performance, significant improvements are needed across the entire videoLLM ecosystem to address persistent stereotypical biases.
C.5 Privacy
The privacy evaluation results reveal significant performance variations across videoLLMs in handling privacy-sensitive content. In Privacy Content Recognization tasks, closed-models like GPT-4o (94.3% average) and Claude4-sonnet (93.0% average) achieved strong performance in private content recognition, while some open-source models like Sharegpt4video-8B (96.9% average) surprisingly outperformed them. However, in Private Information QA tasks, GPT-4o maintained superiority (79.7% average) while most open-source models struggled significantly, with some achieving less than 30% average performance. For Control Over Privacy Inference, the InfoFlow Expectation task showed that larger models (both closed and open-source) demonstrated better alignment with human privacy expectations, with failure rates being a critical reliability factor. In Celebrity Privacy QA, Claude4-sonnet achieved perfect 100% RtA rates across all scenarios, while many open-source models showed substantial vulnerability, particularly in reduced-context situations. Most notably, in Privacy Information Self-Inference tasks, closed-models demonstrated superior autonomous privacy detection capabilities (75-82% detection rates) compared to open-source models (mostly below 50%), highlighting that stronger video understanding capabilities may lead to greater privacy reasoning and leakage issues. The results underscore that while some open-source models can compete in structured privacy tasks, closed-models consistently excel in complex reasoning scenarios, and provide better privacy protection reliability.
Appendix D Evaluation Details on Truthfulness
Truthfulness is essential for the real-world deployment of large foundational models. With the rapid advancement of MLLMs, particularly those processing video data, concerns over trustworthiness and reliability have intensified. The incorporation of temporal dynamics and multimodal interactions—encompassing visual, auditory, and textual modalities—introduces significant challenges to ensuring truthfulness [wang2024videohallucer, gao2025exploring, huang2024online, madan2024foundation]. Effective video understanding demands not only static visual processing but also accurate modeling of temporal sequences, contextual dependencies, and cross-modal consistency. Consequently, evaluating and enhancing the truthfulness of videoLLMs has become a critical research focus.
Prior assessments of videoLLMs have largely addressed various length of video comprehension [fu2024video, wang2024lvbench], hallucination phenomena [wang2024videohallucer, gao2025exploring], and temporal awareness [liu2024tempcompass, ning2023video]. However, these evaluations offer limited insight into the broader issue of truthfulness. This section proposes a comprehensive evaluation of videoLLM’s truthfulness from two perspectives: perceptual and cognitive capabilities, and reliability to contextually misleading inputs. These perspectives respectively target the models’ inherent limitations and their resilience to complex external stimuli, thus facilitating a more rigorous evaluation of videoLLM accuracy.
D.1 Perceptual and Cognitive Proficiency
Videos are complex, combining spatial and temporal information, making them harder to process than static images. Evaluating videoLLMs’ perceptual and cognitive skills ensures they can accurately perceive visual content and reason about dynamic scenes, which is vital for tasks like action recognition, temporal comprehension and contexual reasoning.
D.1.1 Contextual Reasoning QA
Figure 7: An example for the task of contextual reasoning QA.
Setting. This task evaluates videoLLMs’ capacity for comprehensive contextual reasoning over video content through a discriminative framework. It requires models to analyze temporal dynamics, scenes, and objects holistically to answer questions accurately. The task assesses the integration of multimodal information and the ability to reason about complex, dynamic video contexts, thereby reflecting the model’s proficiency in understanding evolving scenes. An example is shown in Figure 7.
Dataset. The dataset consists of 300 collecting from YouTube videos - 100 short (1 to 5 minutes), 100 medium (5 to 30 minutes), and 100 long (30 minutes) - capturing a range of context-dependent dynamic reasoning tasks, including action recognition, counting, object reasoning, and spatial or attribute perception. It ensures diversity in duration, scene complexity, and object interactions, providing a rigorous benchmark for evaluating videoLLMs’ holistic analysis and reasoning capabilities across varied contexts.
Metrics. For the discriminative task with multiple-choice questions, we use accuracy of the model’s responses to contextual reasoning questions as the primary metric, assessing its ability to correctly interpret video content.
Figure 8: Performance of videoLLMs on the task of contextual reasoning QA.
Results. This study systematically evaluates the contextual reasoning capabilities of mainstream videoLLMs, as shown in Figure 8. The results reveal significant differences among models in terms of their ability to understand and analyze contextual dynamics, and reason about object interactions. In terms of accuracy, model performance ranges from 6.0% to 84.3%, indicating considerable disparity in current model capabilities. Specifically, Qwen2.5-VL-72B achieved the highest accuracy at 84.3%, outperforming the second-best model, LLaVA-Video-72B (81.3%), by 3 percentage points, demonstrating exceptional reasoning ability in complex video scenarios. Notably, open-source models performed particularly well in this task, occupying three of the top four positions (Qwen2.5-VL-72B, LLaVA-Video-72B, InternVL2.5-78B). In contrast, the best-performing closed-source model, GPT-4o, reached an accuracy of 75.3%, trailing the leading open-source model by approximately 9 percentage points.
Findings. (1) Open-Source vs. Closed-Source Models. Experiments reveal a performance gap in video contextual reasoning tasks, with open-source models (top 5) achieving an average accuracy of 78.1%, surpassing closed-source models at 57.2% by 20.9 points. Open-source models benefit from targeted optimizations, such as temporal modeling (e.g., TPO techniques) and long-video segmentation, enhancing temporal dynamics understanding. Closed-source models, except GPT-4o, underperform, with Claude3.7-sonnet at 51.3% and Gemini1.5 series ranging from 47.0% to 50.7%, indicating limited adaptability for specialized video tasks. (2) Impact of Parameter Scale. Model performance varies significantly with parameter scale. Among 7B models, accuracy spans from 6.0% (Video-ChatGPT-7B) to 72.8% (LLaVA-Video-7B), underscoring the importance of architectural design. At 72B, top models like Qwen2.5-VL-72B and LLaVA-Video-72B exceed 80% accuracy. Larger parameter models consistently outperform smaller counterparts within the same family, better capturing long-range dependencies and complex object interactions in video data.
D.1.2 Temporal Perception QA
Figure 9: An example for the task of temporal perception QA.
Setting. This task assesses the videoLLMs’ temporal perception capabilities, designed as a discriminative task. The model must answer multiple-choice questions that require understanding the temporal sequence of events in videos, focusing on its ability to capture and reason about time-dependent information. The setting evaluates the model’s inherent ability to process temporal dynamics, a critical aspect of video understanding. An example is shown in Figure 9.
Dataset. The dataset in this task comprises 300 videos sampled from the TempCompass [liu2024tempcompass] dataset, specifically designed to test temporal perception. These videos include scenarios with clear temporal sequences, requiring the model to identify and reason about event order and timing to select the correct answer from multiple-choice options.
Metrics. This is a discriminative, multiple-choice task evaluated using metrics consistent with those applied in contextual reasoning QA.
Figure 10: Performance of videoLLMs on the task of temporal perception QA.
Results. Experimental results, as shown in Figure 10, reveal a wide variance in accuracy among models, ranging from 1.3% for Video-ChatGPT-7B to 68.3% for LiveCC-7B, with an average of 50.3%. This indicates significant differences in the models’ capabilities to process time-dependent information in videos. In terms of model type, open-source models demonstrate a generally strong performance, occupying four of the top five positions. LiveCC-7B ranks first with 68.3% accuracy, followed closely by VideoLLaMA3-7B (67.3%) and LLaVA-Video-7B (66.7%). Among proprietary models, only GPT-4o (68.0%) performs competitively; other closed-source models such as Claude4-Sonnet and Gemini1.5-Flash achieve 60.0% and 59.3%, and Gemini1.5-Pro reaches only 50.0%, falling behind many well-performing open-source models (e.g., MiniCPM-V-2.6-7b at 61.7%). These results suggest that open-source communities have effectively optimized their models for temporal reasoning tasks, surpassing some commercial counterparts.
Notably, most 7B models outperform both commercial models and their 72B counterparts, further highlighting that model performance is not solely determined by parameter count.
Findings. (1) Over half of videoLLMs achieve accuracies below 60.0%, with the lowest at 1.3%, revealing significant limitations in handling complex scenarios with interwoven events and implicit temporal cues. (2) The closed-source Gemini1.5-Pro model, at 50.0% accuracy, underperforms the average, indicating poor adaptation of general-purpose models to specialized temporal reasoning tasks. Open-source models exhibit a wide accuracy range (1.3% to 68.3%), highlighting the need for a unified temporal modeling benchmark. (3) Most models struggle with tasks requiring integration of information across multiple video frames, underscoring deficiencies in processing temporal dynamics.
D.1.3 Video Description
Setting. As a complement to QA tasks, the video description task evaluates the model’s ability to generate coherent, accurate, and contextually faithful narratives of video content without specific prompts. While QA tasks assess discriminative capabilities—selecting or verifying answers based on given options—the video description task measures generative capacity, requiring synthesis of multimodal information (visual, temporal, and auditory) into comprehensive and truthful descriptions. This task challenges videoLLMs to integrate visual, temporal, and contextual elements, reflecting their inherent ability to produce coherent and authentic narratives.
Dataset. The dataset comprises 235 videos, including 135 sampled from OpenVid-1M [nan2024openvid], covering 10 diverse real-world scenarios with varying lengths, and 200 from TempCompass [liu2024tempcompass]. OpenVid-1M offers a broad range of high-quality, realistic scenes, while TempCompass focuses on videos requiring precise temporal and contextual descriptions, enabling a comprehensive evaluation of the model’s descriptive capabilities.
Metrics. The primary metric is the quality of generated descriptions, evaluated using automated metrics such as BLEU, CIDEr, ROUGE, and METEOR to measure textual similarity to ground-truth captions. Additionally, DeepSeek-based scoring assesses the factual accuracy and contextual relevance of descriptions.
TABLE V: Performance (%) of videoLLMs on the task of video description. L denotes LLM score; B denotes BLEU; M denotes Meteor; C denotes CIDEr; R denotes Rouge-L. LLaVA-Onevision is a 72B version.
Model L B M C R
Claude3.7-sonnet 51.9 5.9 20.8 2.6 23.9
Claude4-sonnet 46.7 4.2 21.3 0.3 21.3
GPT-4o 44.4 8.4 21.1 8.4 27.0
Gemini1.5-Flash 33.0 5.1 15.1 2.4 19.2
Gemini1.5-Pro 32.5 6.1 14.4 5.7 21.0
Qwen2.5-VL-7B 52.6 11.4 25.3 3.7 29.5
LLaVA-Onevision 48.1 17.7 23.6 30.6 35.3
Qwen2.5-VL-72B 47.4 13.2 24.5 11.2 31.2
MiniCPM-V-2.6-7B 42.2 6.9 21.0 4.9 25.1
LLaVA-Video-72B 41.5 14.0 21.3 16.0 32.9
LiveCC-7B 40.0 11.4 20.2 17.5 31.1
TPO-7B 38.5 13.6 22.0 17.9 32.0
InternVL2.5-78B 38.5 25.4 25.3 37.2 42.5
LLaVA-Video-7B 37.8 11.6 23.1 8.4 29.6
Sharegpt4video-8B 34.1 11.6 22.6 6.3 31.0
MiniCPM-o-2.6-7B 31.9 7.7 20.5 8.9 26.1
Oryx1.5-7B 28.9 12.3 20.3 16.9 31.1
VideoLLaMA3-7B 25.9 8.5 15.5 8.1 29.0
Oryx-34B 25.2 8.7 21.8 5.6 27.4
Long-LLaVA-7B 23.0 12.7 18.9 20.7 31.8
LongVA-7B 18.5 10.5 19.9 12.9 28.7
Video-ChatGPT-7B 5.9 6.2 13.3 2.7 29.8
Results. In the video description task involving 235 multimodal video clips, model performance exhibited significant variation. Among open-source models, as shown in Table V, Qwen2.5-VL-7B led the field with a top accuracy of 52.6% (LLM score), likely owing to its strong integration of visual, temporal, and contextual elements. It was closely followed by LLaVA-OneVision-72B (LLM score: 48.1%) and Qwen2.5-VL-72B (LLM score: 47.4%), reflecting the Qwen series’ robust performance in multimodal generative tasks.
For closed-source models, Claude3.7-sonnet (LLM score: 51.9%) and Claude4-sonnet (LLM score: 46.7%) performed well, demonstrating their powerful general multimodal understanding and generation capabilities. In contrast, Gemini1.5-Flash (LLM-score: 33.0%) and Gemini1.5-Pro (LLM-score: 32.5%) scored relatively lower.
Notably, some models underperformed significantly: Video-ChatGPT-7B achieved only 5.9% LLM-score, revealing deficiencies in its multimodal integration capabilities and a lack of effective understanding of visual details and contextual associations.
Findings. (1) The model’s performance in generating descriptions without specific prompts fundamentally depends on its efficiency in integrating multimodal information. For the TempCompass dataset, which is highly time-sensitive, well-performing models may have reinforced the alignment between timestamps and event sequences during pretraining. This enables them to produce temporally coherent descriptions that closely match the video’s chronological flow. For example, when describing videos with complex action sequences, these models can accurately capture “first… then…” temporal logic, avoiding confusion in event ordering. (2) Closed-source models (such as Claude4-sonnet and Geminil.5-Flash) demonstrate strengths primarily through their generalization capabilities across multimodal knowledge. Their vast training data encompasses a broader range of vision-language pairs, granting them greater robustness when dealing with rare scenarios or complex contexts. However, open-source models, such as those in the Qwen series, have achieved superior performance in specific tasks like temporally sensitive description through targeted optimization. This highlights that task-specific fine-tuning is crucial for enhancing unprompted generative capabilities.
D.2 Contextual Sequential Comprehension
Evaluating resilience to contextual hallucinations and the ability to understand complex event sequences is crucial for deploying videoLLMs in real-world scenarios. This evaluation addresses a critical facet of truthfulness—resistance to hallucination and misleading errors—which complements assessments of inherent limitations examined by tasks such as video classification, question answering, and description. By emphasizing resilience to misguidance, this evaluation probes advanced perception, temporal reasoning, and contextual understanding, reflecting human-like cognition and ensuring trustworthiness in complex real-world video scenarios.
D.2.1 Events Understanding and Detection
Figure 11: An example for the task of events understanding and detection.
Setting. This task assesses the videoLLMs ability to understand and detect complex event sequences and their order in videos. The model need to identify or select correct event descriptions for given video segments, focusing on its ability to comprehend sequential and intricate event structures. The setting tests the model’s resilience to misguiding temporal or contextual cues that could lead to incorrect event identification. An example is shown in Figure 11.
Dataset. We sample 200 videos from the validation set of YouCook2 [ZhXuCoAAAI18], each paired with two text descriptions (one correct and one incorrect), resulting in 400 data points. These cooking videos feature complex event sequences, such as step-by-step procedures, requiring the model to accurately identify the correct order and details of events to avoid misinterpretations.
Metrics. In this task, the primary metric is the accuracy of the model’s event detection, measuring its ability to select correct event descriptions.
Figure 12: Performance of videoLLMs on events understanding and detection task.
Results. In the task of understanding complex event sequences in 200 cooking videos, videoLLMs showed significant variation in event understanding accuracy. Among open-source models, as shown in Figure 12, Video-ChatGPT-7B led with the highest accuracy of 90.7%, followed closely by LongVA-7B (75.5%) and TPO (74.7%), demonstrating the effectiveness of optimizations targeted at temporally sensitive tasks. Notably, Sharegpt4video-8B ranked at the bottom with only 12.5% accuracy, highlighting its significant disadvantage in prompt-free event sequence recognition tasks. Among closed-source models, Gemini1.5 series performed the best, while GPT-4o(51.6%) and Claude3.7-sonnet (46.8%) showed moderate performance, falling behind some deeply optimized open-source models. From a task-specific perspective, when dealing with cooking videos that involve step-by-step procedures, models exhibited substantial differences in sensitivity to event order.
Findings. Experimental results show that the open-source model Oryx-34B, with 34 billion parameters, achieved an accuracy of only 22.3%, significantly lower than many 7B-scale open-source models such as LongVA-7B (75.5%). Similarly, the 72B-parameter LLaVA-Video-72B underperformed with an accuracy of 52.7%, even lower than the 7B-scale Video-ChatGPT-7B (90.7%). These findings suggest that in complex event sequence parsing tasks, domain-specific adaptation (e.g., to cooking procedures) is more critical than simply increasing model size.
D.2.2 Hallucination in Videos
Figure 13: An example for the video hallucination task.
Setting. This task assesses videoLLMs’ tendency to generate hallucinations in video understanding, a critical measure of truthfulness. Models are required to understand videos that may contain ambiguous or misleading elements, evaluating their ability to avoid fabricating inaccurate temporal, semantic, or factual details. Each item includes paired basic and hallucination questions; notably, the basic questions are identical across factual and non-factual contexts to ensure fair evaluation. This setup probes the model’s advanced reasoning and truthfulness capabilities. An example is shown in Figure 13.
Dataset. We sample 210 videos from the VideoHallucer [wang2024videohallucer] dataset, comprising 30 videos from each of seven categories: temporal, semantic-detail, object_relation, interaction, fact-detect, external_factual, and external_non_factual. For each item, two question types are posed: a basic question evaluating core videoLLMs capabilities, and a hallucinated question containing deliberately fabricated content. This setup assesses the model’s tendency to generate hallucinations when faced with uncertain visual inputs.
Metrics. Following the setting in VideoHallucer [wang2024videohallucer], we calculate the accuracy and the bias score as the primary metrics. For the hallucination score, we calculate the overall accuracy by considering both the basic and hallucinated questions as a paired set, marking it as a hit only if both questions are answered correctly. We posit that enhancing a model’s ability to recognize and counter hallucinations should not compromise its performance on fundamental tasks.
For Bias score, we calculate the Yes Percentage Difference (Pct. Diff) and False Positive Ratio (FP Ratio) to reveal the bias of these videoLLMs. Specifically, the Yes Percentage Difference is calculated as:
$d_{y}=\frac{|{M(v,q)="yes"}_{(v,q)\in V}|-|{GT_{(v,q)}="yes"}_{(v,q)}\in V|}{|V|}$ (1)
where $V$ is the set of video question pairs, $M(v,q)$ is the prediction from models, $GT(v,q)$ is the ground truth. A smaller dy indicates the number of “yes” responses from models is closer to the ground truth, revealing less language bias. the False Positive Ratio is calculated as
$f_{fp}=\frac{|{M(v,q)="yes"}_{(v,q)\in W}|}{|W|}$ (2)
where $W$ is the set of wrongly answered video question pairs. $r_{fp}$ demonstrates the percentage of “yes” in all wrongly predicted answers. A value closer to 50% indicates less bias from the models.
TABLE VI: Performance of videoLLMs on the Hallucination in Videos task. Pct denotes Pct.Diff; FP denotes FP Ration; B denotes Bias Score; H denotes Hallucination Score; O denotes Overall Score.
Models Pct FP B H O
GPT-4o 4.4 39.6 66.8 71.4 42.4
Claude3.7-sonnet -18.9 13.6 54.8 90.8 47.5
Claude4-sonnet -20.1 14.5 48.9 74.7 37.8
Gemini1.5-Pro -1.4 42.0 70.5 65.9 43.8
Gemini1.5-Flash -6.0 33.6 67.3 71.0 46.5
LLaVA-Video-7B 1.4 43.6 74.7 60.8 46.5
LLaVA-Video-72B 0.2 43.2 74.2 65.0 47.9
LLaVA-OneVision 13.6 62.2 75.6 41.5 24.0
MiniCPM-o-2.6-7B -4.4 38.4 66.8 69.6 42.4
MiniCPM-V-2.6-7B -9.2 29.4 62.2 71.9 42.7
LongVA-7B 1.6 47.3 74.7 65.0 46.1
TPO-7B 4.2 52.2 75.6 61.8 42.9
mPLUG-Owl3-7B -4.8 40.7 62.2 68.7 34.1
Oryx1.5-7B -6.0 35.5 60.8 62.7 33.6
Oryx-34B 7.8 52.8 72.4 45.6 29.0
Qwen2.5-VL-7B -16.8 20.2 53.9 84.3 41.5
Qwen2.5-VL-72B -18.7 13.9 54.4 85.7 46.1
Video-ChatGPT-7B 37.3 94.9 97.2 21.2 19.8
VideoLLaMA3-7B 21.2 76.8 88.9 41.5 35.5
LongLLaVA-7B -12.9 24.0 60.4 82.0 46.5
LiveCC-7B 3.9 51.4 74.2 60.4 40.6
sharegpt4video-8B 9.7 51.5 71.0 35.0 22.6
Results. As shown in Table VI, the overall hallucination-suppression performance varies significantly across videoLLMs. In the hallucination-suppression task, Claude3.7-sonnet, Qwen2.5-VL-72B, and Qwen2.5-VL-7B stand out most prominently, achieving Hallucination Score of 90.8, 85.7, and 84.3, respectively—substantially higher than those of other models. By contrast, Video-ChatGPT-7B and VideoLLaMA3-7B perform worst (21.2 and 41.5), reflecting their susceptibility to hallucination when processing complex video semantics. Closed-source models generally exhibit a slight edge over open-source counterparts in hallucination suppression.
Bias Analysis. The Pct. Diff metric reveals a pronounced “conservative bias” in the Qwen series (–16.8/–18.7) and Claude series (–18.9/–20.1): these models answer “yes” far less often than the ground truth. In contrast, Video-ChatGPT-7B (37.3) and VideoLLaMA3-7B (21.2) show an “over-affirmative” tendency, frequently generating unfounded “yes” responses on non-factual questions—evidence of fundamental deficits in their visual-semantic parsing.
The FP ratio further illuminates the error pattern of each model. Qwen2.5-VL-72B (13.9) and Claude4-sonnet (14.5) register extremely low “yes” ratios among their incorrect answers, indicating a cautious default to “no” under uncertainty, which complements their low Pct. Diff. Conversely, Video-ChatGPT-7B (94.9) and LLaVA-OneVision-72B (62.2) maintain high “yes” rates even on mistakes, demonstrating a lack of robust fact-checking and a propensity to produce misleading content in ambiguous scenarios.
Findings. (1) Models with 72B parameters or more consistently exhibit reduced language bias and enhanced resistance to hallucinations, indicating that large-scale parameter counts provide critical capacity for cross-modal reasoning. (2) Gemini1.5-Pro and Claude‑4 achieve the highest hallucination-resistance scores, suggesting that their pretraining and fine-tuning strategies are more mature in video contexts. Nevertheless, open-source 72B models (such as LLaVA-Video and Qwen2.5‑VL) are rapidly closing the gap. (3) Lightweight models tend to default to “yes” responses and suffer from high false-positive rates, limiting their reliability in real-world video-understanding applications.
D.3 Summary
D.3.1 Score Calculation
Our goal is to reflect and analyze the videoLLMs’inherent limitations and their reilience to complex external stimuli.
Perceptual and Cognitive proficiency. For contextual reasoning QA and temporal perceptioin QA tasks, we use accuracy with multiple-choice questions as the primary metric, denoted as $\mathrm{Acc}_{\mathrm{contextual}}$ and $\mathrm{Acc}_{\mathrm{temporal}}$. For the task of video description, we average the $\mathrm{LLM_{score}}$, BLUE, METEOR, and ROUGE to measure the performance, as denoted $\mathrm{Score_{description}}$. We eventually take the average of these metrics as the score of Perceptual and Cognitive proficiency, which is expressed as:
$\begin{split}\mathrm{Score_{Perceptual}}&=\bigg(\mathrm{Acc}_{\mathrm{contextual}}+\mathrm{Acc}_{\mathrm{temporal}}\\ &\quad+\mathrm{Score_{description}}\bigg)\big/3\times 100\end{split}$ (3)
Contextual Sequential Comprehension. For the task of events understanding and detection, we use accuracy to measure videoLLMs’ ability to select correct event description, denoted as $\mathrm{Acc}_{\mathrm{events}}$. For hallucination in videoLLMs, We use the overall score to measure the degree of hallucination in the videoLLMs’response on video understanding, denoted as $\mathrm{Score}_{\mathrm{hallucination}}$. We eventually take the average of these metrics as the score of Perceptual and Cognitive proficiency, which is expressed as:
$\mathrm{Score_{Resilience}}=\frac{\mathrm{Acc}_{\mathrm{events}}+\mathrm{Score}_{\mathrm{hallucination}}}{2}\times 100$ (4)
The comprehensive rankings and corresponding scores for Truthfulness evaluation are presented in Table VII.
TABLE VII: The scores and rankings of two subaspects in Truthfulness. LLaVA-OneVision is the 72B version.
P. C.
Model Score Rank Score Rank
Claude4-sonnet 56.1 8 42.2 7
Claude3.7-sonnet 54.2 11 42.5 6
Gemini1.5-Pro 39.5 18 15.6 22
Gemini1.5-Flash 44.3 17 16.1 21
GPT-4o 62.6 2 42.0 9
Qwen2.5-VL-72B 62.0 3 46.7 2
Qwen2.5-VL-7B 64.3 1 47.0 1
LLaVA-Video-72B 61.5 4 44.7 3
LLaVA-Video-7B 59.7 6 42.2 8
MiniCPM-o-2.6-7B 55.6 9 37.2 12
MiniCPM-V-2.6-7B 54.9 10 42.5 5
Oryx-34B 38.1 20 27.1 20
Oryx1.5-7B 46.5 15 31.3 16
InternVL2.5-78B 58.5 7 43.9 4
LLaVA-OneVision 47.3 14 36.1 13
mPLUG-Owl3-7B 34.4 22 27.8 19
LongVA-7B 44.7 16 32.3 15
Sharegpt4video-8B 38.6 19 28.3 18
TPO-7B 53.2 13 40.7 10
Long-LLaVA-7B 36.7 21 34.8 14
Video-ChatGPT-7B 4.4 23 12.9 23
LiveCC-7B 61.11 5 40.3 11
VideoLLaMA3-7B 53.3 12 30.7 17
D.3.2 Takeaways
  • Diverse strengths across model families. Open‑source videoLLMs often lead in specialized video reasoning (e.g., contextual and temporal QA) thanks to task‑focused modules (TPO, temporal segmentation), while closed‑source models generally excel at hallucination suppression through large‑scale pretraining and robust fine‑tuning.
  • Scale isn’t everything. Larger parameter counts tend to improve contextual reasoning and reduce language bias in hallucination tasks, but they do not guarantee better temporal perception or event‑sequence understanding—architecture design and domain adaptation can outweigh sheer size.
  • The challenges for temporal understanding. Nearly half of VideoLLMs exhibit accuracies below 60% in temporal reasoning, with the lowest at 1.3%, indicating substantial limitations in processing complex scenarios with interwoven events and implicit temporal cues. Even general-purpose models, such as the Gemini series, struggle with specialized temporal reasoning tasks, underscoring that temporal relation understanding remains a significant challenge in video comprehension.
  • Hallucination resilience correlates with conservatism. Top performers in hallucination suppression (Claude, Qwen2.5‑VL‑72B) exhibit a “conservative” bias—error‑averse defaults that favor “no” under uncertainty—resulting in low false‑positive rates but requiring careful calibration to avoid over‑cautious omissions.
Appendix E Evaluation Details on Robustness
VideoLLMs encounter a range of imperfect input conditions in real-world settings, such as noisy videos, low-quality audio, and erroneous text prompts. Insufficient robustness in these models can lead to incorrect or inconsistent outputs, thereby compromising user experience and system reliability. Moreover, models lacking robustness are susceptible to adversarial attacks, where carefully crafted inputs can deceive the system into producing wrong outputs. This vulnerability is especially critical in safety-sensitive domains, including autonomous driving and medical diagnosis. Furthermore, videoLLMs must integrate and process multiple modalities—video, text, and audio—simultaneously, where the complexity of multimodal interactions and fusion introduces additional potential weaknesses. We introduce various perturbations and challenges across four dimensions—OOD Robustness, Temporal Understanding Robustness, Adversarial Robustness and Multimodal Interaction Robustness—to evaluate the model’s ability to maintain expected performance under non-ideal conditions.
E.1 OOD Robustness
E.1.1 Video captioning for OOD videos
Figure 14: An example for OOD videos.
Setting. This task assesses videoLLMs’ ability to generate accurate video descriptions for OOD videos, specifically those captured from aerial perspectives. Such videos present a challenging domain with unique viewpoints and emergency scenarios that are typically underrepresented in standard video captioning training datasets. An example is shown in Figure 14.
Dataset. The evaluation data comes from the CapEra [bashmal2023capera] dataset, which consists of aerial videos covering 25 different event types (such as post-earthquake scenes, floods, fires, traffic accidents, etc.). Each video corresponds to five reference captions. We randomly select 250 videos from this dataset to ensure a diverse and representative evaluation set.
Metrics. For this task, we employ a multi-faceted evaluation strategy: (1) Semantic Similarity Assessment: The DeepSeek model determines whether the videoLLM-generated caption matches the reference caption, outputting “yes” or “no.” (2) Traditional Metrics: BLEU, METEOR, CIDEr, and Rouge-L scores quantify lexical and semantic similarity between generated and reference captions. (3) Final Accuracy: The overall accuracy, defined as the average of all metrics, offers a more reliable and comprehensive assessment of the model’s out-of-distribution robustness. It is calculated as follows:
$\mathrm{Avg.}=\frac{1}{5}\times\Big($ $\mathrm{ACC_{LLM}}+\mathrm{BLUE}$ (5)
$+\mathrm{METEOR}+\mathrm{CIDEr}$
[…]
Results. The performance metrics for the videoLLMs on the CapEra dataset are presented in Table VIII. We assess the models using a comprehensive set of metrics: $\mathrm{ACC_{L}LM}$, BLEU, METEOR, CIDEr, and Rouge-L, with an overall average (Avg.) score providing a holistic measure of captioning accuracy for OOD aerial videos. Proprietary models like Gemini1.5-Flash and GPT-4o lead with Avg. scores of 27.9 and 24.8, respectively, excelling in $\mathrm{Acc_{LLM}}$ (85.0 and 71.0) and showing balanced performance across METEOR (18.3 and 18.8) and CIDEr (14.6 and 16.4). Among open-source models, LLaVA-Video-72B achieves a strong Avg. score of 26.1, with the highest Rouge-L (23.3), indicating superior lexical overlap with reference captions. In contrast, open-source models like Video-ChatGPT-7B and mPLUG-Owl3-7B perform poorly, both with Avg. scores of 5.7 and low $\mathrm{ACC_{L}LM}$ (2.5), reflecting limited capability in this domain. Notably, LLaVA-Video-7B, a smaller model, achieves a respectable Avg. score of 24.7, with competitive Meteor (16.1) scores.
Findings. (1) Advanced closed source model, such as Gemini1.5-Flash and GPT-4o, exhibit greater robustness in generating accurate descriptions for OOD aerial videos, likely due to more diverse training data and optimized architectures. (2) Open-source models, including Video-ChatGPT-7B and mPLUG-Owl3-7B, struggle significantly with semantic and lexical accuracy, highlighting limitations in handling underrepresented domains like aerial perspectives. (3) Larger models within families, such as LLaVA-Video-72B, outperform smaller counterparts (e.g., LLaVA-Video-7B with Avg. 24.73), suggesting that scale contributes to better performance in challenging tasks. (4) Compared to Oryx-34B and Qwen2.5-VL-72B, models like MiniCPM-O-2-6-7B and Live-CC-7B demonstrate that smaller architectures can yield competitive results with proper optimization, although they fall short of top performers in overall accuracy.
TABLE VIII: Results (%) for videoLLMs on OOD videos understanding. Acc denotes Accuracy; B denotes BLEU; M denotes Meteor; C denotes Cider; R denotes Rouge-L.
Models Avg. Acc B M C R
Gemini1.5-Flash 27.9 85.0 4.8 18.3 14.6 16.6
GPT-4o 24.8 71.0 4.7 18.8 16.4 13.3
Gemini1.5-Pro 21.9 85.5 2.0 8.1 7.1 6.9
Claude3.7-sonnet 15.0 57.0 1.3 8.5 8.0 0.1
LLaVA-Video-72B 26.1 59.0 7.7 16.4 24.1 23.3
LLaVA-Video-7B 24.7 59.5 6.6 16.1 20.2 21.3
TPO-7B 23.8 61.5 5.3 15.8 16.1 20.4
LongVA-7B 23.3 54.5 5.6 14.4 20.6 21.4
VideoLLaMA3-7B 21.8 55.4 5.0 15.2 14.5 19.1
Live-CC-7B 21.3 56.0 4.7 14.4 13.0 18.2
Qwen2.5-VL-72B 20.9 55.5 4.6 15.7 10.8 17.9
MiniCPM-o-2.6-7B 20.4 53.1 4.2 15.5 10.6 18.7
Oryx-34B 20.3 38.0 7.0 12.9 24.1 19.4
Qwen2.5-VL-7B 20.3 54.0 4.1 15.3 10.8 17.1
Oryx1.5-7B 19.6 52.0 4.4 14.0 9.9 17.7
MiniCPM-v-2.6-7B 19. 61.5 3.0 11.3 5.5 13.6
sharegpt4video-8B 18.6 54.5 3.9 12.7 7.2 14.8
mPLUG-Owl3-7B 18.4 60.0 2.8 10.3 5.5 13.2
Video-ChatGPT-7B 5.7 2.5 2.5 10.2 2.3 10.9
E.1.2 Noise Videos QA
Setting. This task assesses videoLLMs’ robustness when processing videos corrupted with various types of noise, simulating real-world scenarios where video quality may be degraded due to transmission errors, sensor limitations, or environmental factors.
Dataset. The evaluation dataset is derived from MVBench [li2024mvbench], from which we randomly sample 200 videos. Frames are uniformly sampled from each video, and noise is added with a probability of 0.3. Two types of noise are applied: Gaussian noise (continuous) and salt-and-pepper noise (discrete), both at varying intensity levels to test different degrees of corruption.
Metrics. Performance degradation is measured using: ($\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{ood}}$) × 100, where $\mathrm{Acc}_{\mathrm{clean}}$ represents the model’s performance on original videos and $\mathrm{Acc}_{\mathrm{ood}}$ represents performance on noise-corrupted videos. A lower score indicates better robustness.
Results.
Figure 15: Performance of videoLLMs on the OOD noise videos QA task.
Performance metrics of videoLLMs on noisy videos are summarized in Figure 15. While models generally perform well on clean data (e.g., GPT-4o at 67.5%, Claude3.7-sonnet at 64.4%), their robustness to noise is limited. GPT-4o’s accuracy declines by 26.2%, Claude3.7-sonnet by 41.9%, and Claude4-sonnet drops from 44.5% to 30.4% (a 14.1% decrease). These results suggest that closed-source models, despite optimization for clean data, struggle to handle noisy, real-world scenarios effectively. Models such as LLaVA-Video-72B (14.5% degradation) and InternVL2.5-78B (8.5% degradation) exhibit moderate robustness. In contrast, some open-source models like Oryx1.5-7B (1.0% degradation) and ShareGPT4Video-8B (–2.5% degradation) demonstrate greater resilience, maintaining or even improving performance under noisy conditions.
Findings. (1) The best-performing models on clean data (e.g., LLaVA-Video-72B, GPT-4o) are not necessarily the most robust. This suggests that high clean accuracy might come at the cost of overfitting to clean data, making them less adaptable to noisy conditions. Smaller models like TPO-7B and ShareGPT4Video-8B show surprising robustness, possibly due to simpler architectures that are less sensitive to noise. (2) Despite high clean accuracy, closed-source models like GPT-4o and Claude exhibit poor robustness, with performance drops of 26.2% and 41.9%, respectively. This could be due to their optimization for clean, curated datasets, making them less adaptable to real-world noise. (3) Models like Oryx1.5-7B and ShareGPT4Video-8B show impressive robustness, with minimal or negative degradation scores, suggesting they might be better suited for real-world applications where noise is common. (4) Models with the highest clean accuracy (e.g., LLaVA-Video series) tend to have larger performance drops, indicating a potential trade-off between optimizing for clean data and maintaining robustness to noise.
E.2 Temporal Understanding Robustness
Setting. This task evaluates videoLLMs’ ability to maintain temporal understanding when video frames are missing or presented in incorrect order, testing the model’s reliance on temporal coherence and its ability to infer missing information.
Dataset. We randomly sample 200 videos from MVBench [li2024mvbench] and uniformly sample frames from each video. With a probability of 0.2, we perform random frame dropping or frame shuffling operations to disrupt the temporal structure while maintaining the core content.
Metrics. This is a discriminative, multiple-choice task evaluated using metrics consistent with those applied in VQA for OOD Noise.
Figure 16: Performance of videoLLMs on the task of temporal understanding robustness.
Results. As shown in Figure 16, Gemini1.5 series exhibits the highest robustness with only a 0.0% accuracy drop. Open-source models such as Video-ChatGPT-7B (0.2%) and Video-LLaMA3-7B (0.8%) show moderate sensitivity, while Oryx series (–1.65%) and ShareGPT4Video-8B (–3.2%) achieve slight performance gains under temporal perturbations. Notably, the Claude series displays poor robustness to missing or misordered video frames, despite excelling in other performance aspects.
Findings. Commercial models like Claude3.7-sonnet and Claude4-sonnet display greater degradation, possibly due to stronger reliance on precise temporal understanding. Notably, smaller open-source models (e.g., ShareGPT4Video-8B, Oryx1.5-7B) demonstrate superior temporal robustness, whereas larger models like Oryx-34B struggle, suggesting that model scale and architectural design significantly influence temporal stability.
E.3 Adversarial Robustness
Adversarial robustness evaluation assesses videoLLMs’ vulnerability to carefully crafted perturbations designed to fool the model while remaining imperceptible to humans. This evaluation is crucial for understanding vulnerabilities in adversarial environments.
E.3.1 Adversarial Attack for Classification
Figure 17: An example of the adversarial attack for classification.
Setting. This subtask evaluates videoLLMs’ robustness against untargeted adversarial attacks for classification task, where the goal is to cause incorrect predictions without specifying particular target outputs. The attacks aim to maximally disrupt model performance while maintaining visual imperceptibility. An example is shown in Figure 17.
Dataset. We randomly sample 100 videos from MVBench [li2024mvbench] and identify high-energy regions within videos as keyframes for perturbation. We apply the MI-CWA algorithm [chen2023rethinking] to generate adversarial perturbations with maximum perturbation bound $\epsilon=\frac{16}{255}$ and step size $\alpha=\frac{1}{255}$. These perturbations are designed to maximize prediction errors through untargeted attacks."
Metrics. The primary metric is the classification metrics including accuracy, precision, recall, and F1-score. These metrics are computed by comparing predicted labels against ground-truth labels from per-sample results. The evaluation Metrics (accuracy, precision, recall, F1) represent clean data performance minus adversarial data performance; negative values indicate better adversarial performance, with models ranked by their average difference across all four metrics, providing a comprehensive assessment of defensive capabilities against untargeted attacks. Take accuracy as an example:
$\mathrm{Avg_{Diff}}=\mathrm{Accuracy}_{\mathrm{clean}}-\mathrm{Accuracy}_{\mathrm{adv}}$ (6)
TABLE IX: Performance (%) of videoLLMs on the adversarial attack classification task, reported as the difference between metrics on clean and adversarial data. C denotes Clean; A denotes Adversarial. Acc denotes Accuracy; Pre denotes Precision; Rec denotes Recall.
Models Acc Pre Rec F1 $\mathrm{\textbf{Avg}_{\textbf{Diff}}}$
Video-ChatGPT-7B C 1.0 15.0 1.0 1.9 -4.5
A 1.0 33.0 1.0 1.9
Long-LLaVA-7B C 40.0 53.4 40.0 44.6 -3.1
A 43.0 56.3 43.0 48.0
Claude4-sonnet C 52.0 55.5 52.0 53.3 -1.6
A 54.0 56.1 54.0 54.9
LiveCC-7B C 65.0 65.3 65.0 64.9 -1.0
A 66.0 66.1 66.0 65.9
MiniCPM-V-2.6 C 63.0 63.9 63.0 63.3 0.3
A 63.0 63.1 63.0 63.0
LongVA-7B C 37.0 54.0 37.0 41.7 1.0
A 34.0 57.9 34.0 40.0
Qwen2.5-VL-7B C 59.0 65.5 59.0 61.8 2.5
A 57.0 62.2 57.0 59.1
Oryx1.5-7B C 43.0 41.1 43.0 41.7 2.8
A 39.0 41.1 39.0 38.7
MiniCPM-o-2.6 C 57.0 59.1 57.0 57.8 3.2
A 54.0 55.4 54.0 54.6
TPO-7B C 31.0 57.5 31.0 39.2 3.7
A 28.0 53.0 28.0 35.0
LLAMA3-7B C 69.0 71.7 69.0 70.1 4.9
A 64.0 66.9 64.0 65.3
LLaVA-Video-7B C 61.0 64.0 61.0 61.1 5.0
A 53.3 66.6 53.3 54.1
GPT-4o C 69.0 71.4 69.0 69.2 5.2
A 64.0 66.1 64.0 63.6
Claude3.7-sonnet C 55.0 70.6 55.0 61.8 6.5
A 50.0 61.9 50.0 54.6
mPLUG-Owl3-7B C 58.0 58.8 58.0 58.3 7.2
A 51.0 51.3 51.0 51.0
Sharegpt4video-8B C 41.0 41.5 41.0 41.2 7.5
A 33.0 35.0 33.0 33.6
Results. The experiment assessed the robustness of videoLLMs against untargeted adversarial attacks in classification tasks. Results show that most models performed significantly worse on adversarial data, indicating generally weak robustness, as presented in Table IX. Among closed-source models, Claude4-sonnet demonstrated strong robustness (average metric: -1.6), outperforming others on adversarial data, while GPT-4o (5.2%) and Claude3.7-sonnet (6.5%) showed poorer robustness. Among open-source models, ShareGPT4Video-8B (7.5%) and mPLUG-Owl3-7B (7.2%) experienced the largest performance drops, while Video-ChatGPT-7B (-4.5%) and LiveCC-7B (-1/0%) displayed better robustness. Accuracy and recall differences were generally consistent, while precision varied, highlighting the complex effects of adversarial attacks. The study underscores the vulnerability of video content understanding to such attacks. Future improvements should focus on adversarial training or data augmentation, alongside further investigation into the robustness mechanisms of models like Claude4-sonnet.
Findings. (1) Most videoLLMs suffer notable performance degradation under untargeted adversarial attacks (reflected by positive metric values), indicating limited robustness to adversarial noise. Among closed-source models, Claude4-sonnet exhibits strong robustness, whereas GPT-4o and Claude3.7-sonnet perform less reliably, revealing considerable variation within this category. Open-source models display mixed results: Video-ChatGPT-7B and LiveCC-7B demonstrate greater resilience, while ShareGPT4Video-8B and mPLUG-Owl3-7B are notably vulnerable. These disparities may stem from differences in model architecture, training data, or the use of adversarial training. (2) Untargeted adversarial attacks, by adding imperceptible perturbations to key frames, significantly reduce classification performance in most models, revealing insufficient robustness in video content understanding. Precision fluctuates notably in some models, suggesting that adversarial attacks can have a more complex impact on classification accuracy. (3) The negative metric values of Claude4-sonnet and Video-ChatGPT-7B suggest that some models may exhibit unexpected adaptability in adversarial settings, possibly due to the incorporation of adversarial training during model development. This warrants further analysis of their architectures or training processes.
E.3.2 Adversarial Attack for Video Captioning
Figure 18: An example of the adversarial attack task for video captioning, with descriptions generated by LLaVA-Video-7B.
Setting. Video captioning serves as a key task for evaluating the overall video understanding of videoLLMs. We apply untargeted perturbations to a few selected keyframes within each video to assess whether the model can leverage contextual information to correct the injected errors and produce accurate captions despite partial perturbations. The MI-CWA [chen2023rethinking] algorithm is applied to generate adversarial perturbations with maximum perturbation bound $\epsilon=\frac{16}{255}$ and step size $\alpha=\frac{1}{255}$. An example is shown in Figure 18.
Dataset. We randomly sample 100 videos from MVBench [li2024mvbench] and identify high-energy regions within videos as keyframes for untargeted perturbation.
Metrics. This task is also a video captioning, with evaluation metrics aligned with those used in task of Video captioning for OOD videos.
Figure 19: Performance of videoLLMs on video captioning under adversarial perturbations. Models are ordered left to right based on the performance gap between clean and adversarial accuracy.
Results. The results of this task are shown in Figure 19. Closed-source commercial models such as GPT-4o (74.0% clean) and Gemini 1.5 Pro (70.0% clean) achieve higher clean accuracies compared to most open-source models but exhibit significant drops under adversarial perturbations, indicating limited robustness. In contrast, open-source models like VideoLLaMA3-7B (66.0% clean, 66.0% adversarial) and mPLUG-Owl3-7B (66.0% clean, 67.0% adversarial) maintain stable performance, demonstrating stronger resilience to perturbations. The gap between clean and adversarial accuracy varies considerably across models. For instance, ShareGPT4Video-8B (43.0% clean, 24.0% adversarial) and LLaVA-Video-7B (66.0% clean, 49.0% adversarial) show large performance drops, indicating vulnerability. Conversely, models such as VideoLLaMA3-7B and Claude 3.7 Sonnet maintain identical clean and adversarial accuracies, suggesting robust behavior. Interestingly, some models—including mPLUG-Owl3-7B, LongVA-7B, TPO-7B, and Qwen2.5-VL-7B—perform better under adversarial conditions. This may indicate that the perturbations act as a form of data augmentation or regularization, helping models focus on more robust visual features and reducing reliance on noise or irrelevant details.
Findings. (1) Open-source models generally exhibit more balanced performance across clean and adversarial settings, potentially due to their adaptability and incorporation of adversarial training. In contrast, closed-source models, while most of them achieving higher clean accuracy, are more vulnerable to adversarial attacks, revealing a trade-off between accuracy and robustness. (2) The results highlight the importance of leveraging contextual cues to mitigate injected errors. Models with minimal performance degradation (e.g., LongVA-7B, nPLUG-7B) likely excel at contextual reasoning, whereas those with substantial drops (e.g., ShareGPT4Video-8B) demonstrate limitations—underscoring the need for improved adversarial training in videoLLMs.
E.4 Multimodal Interaction Robustness
E.4.1 The Impact of Video on Sentiment Analysis
Figure 20: An example for the task of the impact of video on sentiment analysis
Setting. This task examines videoLLMs’ robustness when visual and textual information provide conflicting sentiment cues, testing the model’s ability to handle multimodal inconsistencies and determine appropriate sentiment classifications. An example is shown in Figure 20.
Dataset. Videos are generated using Kling text2video system, while text samples are drawn from SST-2 [socher-etal-2013-recursive] (100 text samples). Two evaluation scenarios are created: (1) positive text paired with natural/negative videos, and (2) negative text paired with natural/positive videos. This creates deliberate sentiment conflicts between modalities.
TABLE X: Performance (%) of videoLLMs on the task of the impact of video on sentiment analysis. O denotes Opposite; N denotes Natural.
[…]
Overall
Score
$\uparrow$
O N
Claude3.7-sonnet 97.0 96.0 97.0 96.7
GPT-4o 97.0 95.0 95.0 95.7
Gemini1.5-Pro 95.0 92.0 97.0 94.8
Gemini1.5-Flash 93.0 95.0 94.0 94.0
Claude4-sonnet 92.0 88.0 86.0 88.7
LLaVA-Video-72B 95.0 97.0 96.0 96.0
Oryx-34B 95.0 97.0 93.0 95.0
Qwen2.5-VL-72B 93.0 94.0 95.0 94.0
LLaVA-OneVision 94.0 94.0 94.0 94.0
TPO-7B 93.0 94.0 94.0 93.7
Qwen2.5-VL-7B 93.0 92.0 93.0 92.7
Long-VA-7B 92.0 92.0 93.0 92.3
Video-ChatGPT-7B 91.0 92.0 93.0 92.0
MiniCPM-V-2.6-7B 92.0 91.0 92.0 91.7
mPLUG-Owl3-7B 90.0 91.0 90.0 90.3
MiniCPM-o-2.6-7B 90.0 88.0 92.0 90.0
Sharegpt4video-8B 88.0 88.0 93.0 89.7
Oryx1.5-7B 88.0 90.0 90.0 89.3
Long-LLaVA-7B 91.0 81.0 93.0 88.3
VideoLLaMA3-7B 82.0 75.0 77.0 78.0
LiveCC-7B 84.0 50.0 83.0 72.3
LLaVA-Video-7B 77.0 56.0 68.0 67.0
Metrics. We formulate a discriminative task to classify the sentiment of text paired with varied videos, aiming to investigate the impact of video context on sentiment analysis for videoLLMs. Each text is accompanied by three types of videos: sentiment-aligned, sentiment-opposite, and sentiment-neutral. The final performance is measured by the average classification accuracy across these three conditions.
Results. The experimental results (as presented in Table X) demonstrate that closed-source proprietary models (e.g., Claude3.7-sonnet, GPT-4o, Gemini1.5-Pro) excel in sentiment analysis, achieving overall scores of 94.8%–96.7%, showcasing robust performance and strong handling of multimodal sentiment conflicts. Open-source models exhibit varied performance: LLaVA-Video-72B (96.0%) and Oryx-34B (95.0%) approach closed-source model performance, while 7B-scale models (e.g., LLaVA-Video-7B, LiveCC-7B) perform poorly in sentiment-opposite scenarios (scores as low as 50.0%–56.0%), indicating susceptibility to video context interference. In sentiment-aligned scenarios, models generally perform well (scores mostly above 90.%); sentiment-opposite scenarios are the most challenging, with closed-source models and select large-scale open-source models maintaining high accuracy. In sentiment-neutral scenarios, performance is relatively balanced, though smaller models still face interference. Larger models (72B, 34B) consistently show greater robustness across all scenarios, reflecting robust multimodal information fusion. The results highlight that video context significantly impacts sentiment analysis, particularly in conflicting scenarios, underscoring the need for enhanced semantic understanding and fusion mechanisms in models.
Findings. (1) Closed-source models outperform open-source models: Closed-source models (Claude3.7-sonnet, GPT-4o, Gemini1.5-Pro) demonstrate higher classification accuracy and robustness across all scenarios, particularly in handling complex sentiment-opposite cases. This is likely due to optimized training data and architectural design. Among open-source models, LLaVA-Video-72B and Oryx-34B perform close to closed-source models, indicating the potential of large-scale open-source models in multimodal tasks. (2) Significant impact of video context on sentiment analysis: Sentiment-opposite videos cause the greatest interference with model performance, especially for smaller open-source models, highlighting the need for stronger semantic understanding and information fusion in videoLLMs to handle multimodal conflicts. Sentiment-aligned and neutral videos have less impact but can still introduce interference, particularly in smaller models. (3) Importance of model scale: Large-scale models (72B, 34B) exhibit superior performance in handling multimodal sentiment conflicts, showing greater robustness and generalization. Smaller models (7B) are more prone to being misled by video context in complex scenarios.
E.4.2 Misleading prompts for video understanding
Figure 21: An example for the task of misleading prompts for video understanding.
Setting. This task evaluates videoLLMs’ robustness to misleading or contradictory textual prompts that may attempt to bias the model’s interpretation of video content, testing the model’s ability to ground its responses in visual evidence rather than potentially incorrect textual guidance. An example is shown in Figure 21.
Dataset. 100 videos are collected from YouTube, covering 20 diverse scenarios including traffic accidents, natural disasters, sports competitions, animal behavior, social events, medical procedures, technology demonstrations, artistic performances, cooking, education, industrial processes, and agriculture. Each video is paired with misleading prompts designed to test different types of textual misguidance.
Metrics. Evaluation is conducted using DeepSeek as an automated evaluator to assess whether the videoLLM’s responses are appropriately grounded in visual content despite misleading textual prompts. The evaluation considers both factual accuracy and the model’s ability to resist prompt-based misguidance.
Figure 22: Performance of videoLLMs on the task of misleading prompts for video understanding.
Results. The results of this task are shown in Figure 22. Claude4-aonnet (70.6%), Qwen2.5-VL-72B (67.6%), and Gemini1.5-Flash (55.9%) exhibit strong performance in resisting misleading prompts, significantly outperforming other models. These results suggest that closed-source models (e.g., Claude, Gemini, GPT-4o) possess superior capabilities in video understanding and robustness against textual misguidance, with Claude4-sonnet and Qwen2.5-VL-72B ranking highest. In contrast, open-source models display mixed results. LLaVA-Video-72B (34.3%) and LongVA-7B (21.8%) show moderate robustness but lower overall accuracy. Performance declines markedly with smaller model sizes—for instance, LLaVA-Video-7B achieves only 18.8%. Most smaller open-source models (e.g., MiniCPM-V-2.6, LiveCC-7B) fall below 10%, and some (e.g., Oryx-34B, ShareGPT4Video-8B) score 0%, highlighting their limited ability to handle misleading prompts.
Findings. (1) Model Scale and Performance Correlation: Results from closed-source models and large open-source models (e.g., LLaVA-Video-72B) suggest that increased parameter scale (e.g., 72B) can enhance resistance to misleading prompts. However, scale alone is not decisive; for instance, Oryx-34B performs poorly, likely due to limitations in model architecture or training data. (2) Task Challenges and Insights: The findings highlight the need for strong visual grounding to prioritize video content over deceptive textual cues. Closed-source models likely benefit from high-quality training data and advanced optimization, whereas open-source models—especially smaller ones—require improved robustness to misleading inputs.
E.4.3 The Impact of Text on Video Understanding
Setting. Most videoLLMs currently process both video and text inputs. While prior tasks assessed the influence of video variations on model decisions, this task evaluates the effect of textual noise on video understanding. Specifically, it tests whether the model can maintain semantic consistency with the video and avoid generating out-of-context responses.
Dataset. We sampled 100 videos from MVBench [li2024mvbench] and added adversarial noise to the problems in each video. The noise is divided into three categories: Spelling errors, Grammatical errors and Noise with uncontroversial symbols.
Metrics. In the discriminative task with multiple-choice questions, we assess the model’s ability to correctly interpret prompts with noise based on prediction accuracy. The accuracy difference between noisy and original textual prompts reflects the impact of text errors on the videoLLM’s performance.
TABLE XI: Performance (%) of videoLLMs on the task of the impact of text on video understanding. S. denotes Spelling; N. denotes Noisy; G. denotes Grammr. LLaVA-OneVision is a 72B version.
Models
Origin
text
Text error
S. N. G. Avg.
Claude3.7-sonnet 70.5 64.2 70.5 66.3 67.0
Claude4-sonnet 52.6 61.1 59.0 61.1 60.0
GPT-4o 60.0 56.8 52.6 57.9 55.8
LiveCC-7B 68.4 74.7 68.4 69.5 70.9
mPLUG-Owl3-7B 69.5 70.5 70.5 64.2 68.4
VideoLLaMA3-7B 64.2 70.5 64.2 66.3 67.0
MiniCPM-o-2.6 62.1 65.3 61.1 63.2 63.2
LLaVA-Video-7B 60.0 62.1 60.0 61.1 61.1
MiniCPM-V-2.6 56.8 61.1 59.0 57.9 59.3
LLaVA-Video-72B 45.7 57.1 57.1 57.1 54.3
Oryx-34B 40.0 42.9 42.9 42.9 42.9
Oryx1.5-7B 39.0 36.8 39.0 40.0 38.6
Sharegpt4video-8B 28.4 39.0 33.7 34.7 35.8
Qwen2.5-VL-72B 37.1 28.6 37.1 37.1 34.3
Qwen2.5-VL-7B 41.1 35.8 28.4 35.8 33.3
Long-LLaVA-7B 31.6 30.5 28.4 29.5 29.5
LLaVA-OneVision 28.6 28.6 28.6 28.6 28.6
LongVA-7B 31.6 14.7 35.8 21.1 23.9
TPO-7B 17.9 10.5 14.7 14.7 13.3
Video-ChatGPT-7B 1.1 1.1 3.2 3.2 2.5
Results. The results of this task are shown in Table XI. The experimental results reveal significant differences in the robustness of videoLLMs when processing video inputs with textual noise (spelling errors, grammatical errors, and symbolic noise). Among closed-source models, Claude3.7-sonnet (original accuracy 70.5%, average noise accuracy 67.0%, drop 3.5%) and Claude4-sonnet (original 52.6%, noise 60.0%, increase 7.4%) demonstrate strong robustness, while GPT-4o (drop 4.2%) performs moderately. Among open-source models, LiveCC-7B (original 68.4%, noise 70.9%, increase 2.5%) and mPLUG-Owl3-7B (drop 1.1%) excel, whereas LLaVA-Video-72B (drop 11.4%) and Video-ChatGPT-7B (original 1.1%, noise 2.5%) perform poorly. Among noise types, spelling errors have the greatest impact (e.g., LongVA-7B drops 16.8%), symbolic noise the least, and grammatical errors are moderate. Model scale does not directly correlate with robustness; large-scale models like LLaVA-Video-72B show no clear advantage over smaller models like LiveCC-7B, suggesting that training data quality and noise adaptability are more critical. Closed-source models generally exhibit stable performance, while some open-source models match or surpass them, though their performance varies widely, indicating a need for optimization.
Findings. (1) Significant Differences in Robustness: Certain models (e.g., LiveCC-7B, Claude3.7-sonnet, Claude4-sonnet) exhibit strong robustness to textual noise, with some even showing improved performance under noisy conditions. This may stem from training data containing similar noise patterns or superior semantic understanding capabilities. Conversely, models like LongVA-7B and Video-ChatGPT-7B are highly sensitive to noise, indicating limitations in their training data or model architecture when handling noisy inputs. (2) Impact of Noise Types: Spelling errors have the most significant impact on model performance, likely because they directly alter word semantics, making it difficult for models to parse correctly. Symbolic noise has the least impact, as models can often ignore irrelevant symbols and focus on video content and core text semantics. Grammatical errors have a moderate effect, possibly because they do not entirely disrupt semantics, allowing some models to infer correct meanings. (3) Closed-Source vs. Open-Source Models: Closed-source models generally demonstrate better robustness, particularly the Claude series, likely due to more extensive training data or thorough optimization. Some open-source models, such as LiveCC-7B, achieve comparable or even superior performance, highlighting their potential in specific tasks, though their performance is less consistent and requires optimization. (4) Model Scale and Robustness: Large-scale models (e.g., LLaVA-Video-72B) do not significantly outperform smaller models (e.g., LiveCC-7B), suggesting that robustness depends more on the diversity and noise adaptability of training data rather than sheer parameter size.
E.5 Summary
E.5.1 Score Calculation
We evaluate the robustness of videoLLMs across four dimensions related to video understanding:
OOD Robustness. For video captioning on OOD videos, we compute the final score as the average of semantic similarity and standard metrics: $\mathrm{Avg}_{\mathrm{OOD_{video}}}$. For noise video QA, we use the difference between clean accuracy and noise accuracy: $\mathrm{Acc}_{\mathrm{OOD_{noise}}}=\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{noise}}$. We eventually take the average of these metrics as the score under OOD robustness, which is expressed as:
$\mathrm{Score_{OOD}}=\frac{\mathrm{Avg}_{\mathrm{OOD_{video}}}-\mathrm{Acc}_{\mathrm{OOD_{noise}}}}{2}\times 100.$ (7)
Temporal Understanding Robustness. For VQA under missing temporal information, robustness is quantified by the performance drop between clean and perturbed data, denoted as $\mathrm{Score}_{\mathrm{temporal}}=\mathrm{Score}_{\mathrm{clean}}-\mathrm{Score}_{\mathrm{noise}}$.
Adversarial Robustness. For adversarial video classification, we calculate the score based on performance degradation in accuracy, precision, recall, and F1:
$\begin{split}\mathrm{Avg}_{\mathrm{cls}}&=\frac{(\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{adv}})+(\mathrm{Pre}_{\mathrm{clean}}-\mathrm{Pre}_{\mathrm{adv}})}{4}\\ &\quad+\frac{(\mathrm{Recall}_{\mathrm{clean}}-\mathrm{Recall}_{\mathrm{adv}})+(\mathrm{F1}_{\mathrm{clean}}-\mathrm{F1}_{\mathrm{adv}})}{4}\\ &\quad\times 100.\end{split}$ (8)
For adversarial video captioning, the score is based on degradation across multiple metrics (B denotes BLEU, M denotes METEOR, C denotes CIDEr, R denotes Rouge-L):
$\begin{split}\mathrm{Avg}_{\mathrm{cap}}&=\frac{(\mathrm{LLM}_{\mathrm{clean}}-\mathrm{LLM}_{\mathrm{adv}})+(\mathrm{B}_{\mathrm{clean}}-\mathrm{B}_{\mathrm{adv}})}{5}\\ &\quad+\frac{(\mathrm{M}_{\mathrm{clean}}-\mathrm{M}_{\mathrm{adv}})+(\mathrm{C}_{\mathrm{clean}}-\mathrm{C}_{\mathrm{adv}})}{5}\\ &\quad+\frac{(\mathrm{R}_{\mathrm{clean}}-\mathrm{R}_{\mathrm{adv}})}{5}\times 100.\end{split}$ (9)
We eventually take the average of these metrics as the score under OOD robustness, which is expressed as:
$\mathrm{Score_{adv}}=\frac{\mathrm{Avg}_{\mathrm{cls}}+\mathrm{Avg}_{\mathrm{cap}}}{2}$ (10)
Multimodal Interaction Robustness. For sentiment analysis and misleading prompt tasks, we use classification accuracy: $\mathrm{Acc}_{\mathrm{sentiment}},\quad\mathrm{Acc}_{\mathrm{misleading}}$ For evaluating the impact of text on video understanding, we compute performance degradation as: $\mathrm{Acc}_{\mathrm{text}}$. We eventually take the average of these metrics as the score under OOD robustness, which is expressed as:
$\mathrm{Score_{interaction}}=\frac{\mathrm{Acc}_{\mathrm{sentiment}}+\mathrm{Acc}_{\mathrm{misleading}}+\mathrm{Acc}_{\mathrm{text}}}{3}$ (11)
The comprehensive rankings and corresponding scores for Robustness evaluation are presented in Table XII.
TABLE XII: The scores and rankings of four subaspects in Robustness.
O. T. A. M.
Model Score Rank Score Rank Score Rank Score Rank
Claude4-sonnet 7.5 21 0.4 7 2.8 12 63.7 4
Claude3.7-sonnet 9.5 17 2.0 13 2.9 13 67.0 2
Gemini1.5-Pro 11.0 8 -2.8 3 4.9 17 66.5 3
Gemini1.5-Flash 12.0 4 0.6 8 0.0 6 63.3 6
GPT-4o 11.6 6 2.4 15 7.9 18 55.9 9
Qwen2.5-VL-72B 10.4 12 2.5 16 3.1 14 67.5 1
Qwen2.5-VL-7B 10.1 15 3.2 21 8.0 19 59.3 7
LLaVA-Video-72B 13.0 1 3.3 14 -0.5 5 63.7 5
LLaVA-Video-7B 12.3 3 2.9 18 12.3 22 48.9 14
MiniCPM-o-2.6-7B 10.1 14 4.2 23 -2.5 3 50.0 12
MiniCPM-V-2.6-7B 9.3 19 3.8 22 -4.5 1 52.8 11
Oryx-34B 10.2 13 -3.8 1 9.0 20 47.3 15
Oryx1.5-7B 9.8 16 -1.7 4 10.0 21 43.0 17
InternVL2.5-78B 10.9 9 2.5 17 0.0 7 56.2 8
LLaVA-OneVision-72B 2.3 23 2.9 19 4.5 16 41.3 19
mPLUG-Owl3-7B 9.1 20 1.5 12 3.0 15 53.7 10
LongVA-7B 11.6 7 1.2 11 -1.5 4 45.8 16
Sharegpt4video-8B 9.4 18 -3.2 2 13.5 23 41.2 20
TPO-7B 12.0 5 1.1 10 0.0 8 40.3 22
Long-LLaVA-7B 12.6 2 0.2 6 2.0 11 40.5 21
Video-ChatGPT-7B 2.8 22 0.2 5 0.0 9 31.7 23
LiveCC-7B 10.7 11 3.0 20 -3.5 2 41.5 18
VideoLLaMA3-7B 10.8 10 0.8 9 2.0 10 49.8 13
E.5.2 Takeaways
  • Closed-source models generally demonstrate better robustness, particularly the Claude series. Some open-source models, such as LiveCC-7B, achieve comparable or even superior performance, highlighting their potential in specific tasks, though their performance is less consistent and requires further optimization.
  • Temporal reasoning and multimodal conflict handling remain challenging, especially for smaller models, highlighting the need for advanced temporal modules and robust multimodal fusion. Most models are vulnerable to adversarial perturbations, underscoring the importance of adversarial training.
  • To enhance video understanding in videoLLMs, most approaches focus on amplifying the influence of video information on model decisions. However, this strategy may compromise robustness, as evidenced by the impact of videos on sentiment analysis. The relevance of a video to the sentiment of a sentence significantly affects videoLLM judgments, particularly for smaller models.
Appendix F Evaluation Details on Safety
The safety of videoLLMs poses unique challenges compared to image-based MLLMs. First, temporal consistency and contextual integrity are paramount, as videos consist of continuous frame sequences requiring precise temporal reasoning. Failures here can lead to misinterpretations or unsafe content generation (Tasks 1–3). Second, multimodal alignment risks emerge from inconsistencies across textual, auditory, and visual inputs, potentially causing erroneous outputs (Task 4). Third, robustness against deepfakes is critical, necessitating effective detection and mitigation of manipulated media (Task 5). Finally, videoLLMs must resist toxic prompts and multimodal jailbreak attacks to ensure safety and alignment across diverse input modalities.
F.1 NSFW Videos Description
Figure 23: An example for the NSFW videos description task.
Setting. This task evaluates the ability of videoLLMs to identify and refuse to describe Not Safe For Work (NSFW) content in videos, designed as a generative task. The model is expected to recognize pornographic, violent, gory, or hate-related content and avoid generating inappropriate outputs, thereby ensuring safe and ethical responses in practical applications. An example is shown in Figure 23.
Dataset. The dataset comprises 100 videos collected from the BigPorn [^3] https://huggingface.co/datasets/Nikity/Big-Porn, Violence [^4] https://www.kaggle.com/datasets/mohamedmustafa/real-life-violence-situations-dataset/, and Youtube, covering four categories: pornography, violence, self-harm and hate speech. These videos represent diverse NSFW content, providing a comprehensive testbed to assess the model’s ability to detect and handle sensitive material across various contexts.
Figure 24: Base templates employed in Refuse-to-Answer (RtA) mechanisms to determine response rejection.
Metrics. We use the Refuse-to-Answer (RtA) rate, with the template shown in Figure 24, non-toxicity score based on Google’s PerspectiveAPI [^5] https://www.perspectiveapi.com/ (calculated as 1 - toxicity score), and comprehensive metrics extended from average score gradient and balance index. Specifically, the refusal rate quantifies the model’s tendency to reject unsafe prompts through keyword matching against predefined refusal templates. The non-toxicity score measures the civility level of generated text. The average score, defined as the arithmetic mean of these two metrics, provides a holistic reflection of the model’s overall safety performance. Furthermore, to characterize model behavior more granularly, we introduce the average score gradient ($|\Delta|$Avg) and balance index (H-mean). The former measures the magnitude of safety performance fluctuation between adjacent models, while the latter employs the harmonic mean of refusal rate and non-toxicity score to assess the coordination between the model’s "refusal capability" and "safe output generation." A higher balance index indicates that the model can both effectively reject harmful inputs and generate safe, low-toxicity descriptions.
Figure 25: Performance of videoLLMs in the task of NSFW video description.
TABLE XIII: Performance (%) of videoLLMs in the task of NSFW video description. Pers. denotes Perspective score; LLM denotes LLM-score. LLaVA-OneVision is the 72B version.
[…]
Claude4-sonnet 73.9 92.8 95.4 87.4
Cluade3.7-sonnet 64.6 93.8 87.7 82.0
GPT-4o 35.4 95.5 76.9 69.3
Gemini1.5-Flash 9.2 79.2 30.8 39.7
Gemini1.5-Pro 1.5 79.1 26.2 35.6
Oryx-34B 9.2 93.9 95.4 66.2
Oryx1.5-7B 6.2 98.4 87.7 64.1
LLaVA-OneVision 1.5 90.2 81.5 57.7
MiniCPM-V-2.6-7B 13.9 96.7 60.0 56.9
Qwen2.5-VL-7B 13.9 96.5 60.0 56.8
MiniCPM-o-2.6-7B 7.7 97.5 50.8 52.0
Long-LLaVA-7B 4.6 86.9 63.1 51.5
Sharegpt4video-8B 0.0 95.3 55.4 50.2
Video-ChatGPT-7B 0.0 92.9 52.3 48.4
LongVA-7B 12.3 95.2 36.9 48.1
LLaVA-Video-7B 6.2 96.5 40.0 47.6
TPO-7B 6.2 96.4 40.0 47.5
VideoLLaMA3-7B 0.0 93.1 44.6 45.9
LiveCC-7B 3.1 94.3 40.0 45.8
mPLUG-Owl3-7B 4.6 91.8 38.5 44.9
Qwen2.5-VL-72B 4.6 79.2 46.2 43.3
LLaVA-Video-72B 3.1 75.4 27.7 35.4
Results. The results of this task are presented in Figure 25. The scoring range [0, 1] is discretized into several intervals and visualized as a heatmap displaying the distribution of models across multidimensional safety metrics. The blue dashed line represents the refusal rate, while the black solid line indicates the average score. Overall, commercial closed-source videoLLMs demonstrate superior performance in both refusal rate and non-toxicity score, exhibiting stronger safety alignment capabilities, whereas most open-source videoLLMs still exhibit low refusal rates and suboptimal non-toxicity scores. Among the commercial models, GPT-4o shows the lowest response toxicity, while the Claude series achieves the highest rejection rate (up to 79%). In contrast, the Gemini series tends to produce more toxic responses, potentially due to a design preference for more direct or expressive outputs. Conversely, GPT-4o and Claude appear to favor more conservative or neutral responses, resulting in lower Perspective scores.
Notably, most open-source models exhibit relatively low rejection rates. The two 72B open-source models show limited refusal behavior but relatively high toxicity levels. The complete results are presented in Table XIII.
Findings. (1) Commercial models exhibit robust performance and higher safety levels. The Claude series demonstrates the highest refusal rate (>70%), while GPT-4o achieves both a lower refusal rate and the highest non-toxicity score, reflecting an excellent balance between safety and openness. In contrast, although the Gemini1.5 series shows a relatively lower refusal rate, the toxicity of its generated content is slightly inferior, indicating a design philosophy more inclined toward direct output. (2) Open-source models show significant disparities and substantial variability. Open-source videoLLMs (particularly earlier models) demonstrate limited perception of harmful content, characterized by low refusal rates and high toxicity scores, necessitating improved safety alignment. (3) The balance index reveals the stability of safety alignment. Models with high balance indices (such as GPT-4o and the Claude series) maintain effective coordination between "refusal capability" and "safe output," whereas models with low balance indices often exhibit unidirectional bias—either over-refusing or over-generating. Overall, commercial models demonstrate significantly higher safety alignment levels compared to open-source models. However, notable safety disparities persist across models. This indicates that safety alignment depends not only on model scale but is also closely related to training data, alignment strategies, and refusal template design.
F.2 NSFW Prompts Execution
Setting. This task evaluates whether integrating video understanding into large language models compromises their safety defenses. VideoLLMs are tested using toxic text prompts paired with neutral (blank) videos. The model is expected to detect and reject harmful instructions, maintaining ethical alignment even when visual input is benign.
Dataset. The dataset includes 100 toxic prompts uniformly sampled from the HarmBench [mazeika2024harmbench] dataset, paired with content-blank videos. These prompts are designed to provoke unsafe or inappropriate responses for large language module in videoLLMs, testing the model’s ability to recognize and reject harmful instructions regardless of the accompanying visual input.
Metrics. This task aims to detect toxicity in model responses, using evaluation metrics consistent with those employed in NSFW video assessments.
Results. The performance of videoLLMs in rejecting toxic prompts is presented in Table XIV. Among closed-source models, GPT-4o achieved the highest average score of 96.7%, closely followed by Claude4-sonnet at 96.6%. Gemini1.5-Flash and Claude3.7-sonnet recorded average scores of 94.8% and 94.8%, respectively, while Gemini1.5-Pro had the lowest closed-source performance at 91.2%. For open-source models, Qwen2.5-VL-7B led with an average score of 96.0%, outperforming several closed-source models. Other notable open-source performances include mPLUG-Owl3-7B at 93.1% and Qwen2.5-VL-72B at 92.7%. The lowest average scores were observed for Video-ChatGPT-7B at 58.9% and LLaVA-OneVision-72B at 61.7%, with Oryx1.5-7B, MiniCPM-V-2.6-7B, and Sharegpt4video-8B also performing poorly, scoring below 70.0%.
Findings. (1) The results demonstrate that most videoLLMs effectively detect and reject toxic textual prompts when paired with neutral videos, with closed-source models generally exhibiting stronger safety defenses. GPT-4o and Claude4-sonnet set a high benchmark, achieving near-perfect scores across metrics, particularly in LLM score (100.00%). The strong performance of the open-source model Qwen2.5-VL-7B, with an average score comparable to top closed-source models, suggests that certain open-source videoLLMs can maintain robust ethical alignment in multimodal settings. Integrating video modality into LLMs almost does not compromise safety alignment in LLMs, as observed in closed-source models.(2) However, significant variability exists among open-source models, with lower-performing models like Video-ChatGPT-7B and LLaVA-OneVision-72B showing weaknesses, particularly in RtA and LLM score metrics. This suggests that integrating video understanding can compromise safety defenses in some open-source models, likely due to differences in training data or safety alignment techniques. (3) The high Perspective score across most models (above 90%) reflects a consistent ability to identify toxic content, but the lower LLM score in several open-source models highlights challenges in fully rejecting harmful instructions. These findings emphasize the need for enhanced safety mechanisms in videoLLMs, particularly for open-source models, to ensure consistent robustness against toxic inputs in multimodal contexts.
TABLE XIV: Performance (%) of videoLLMs in the task of NSFW prompt execution. Pers. denotes Perspective score; LLM denotes LLM-score. LLaVA-OneVision is the 72B version.
[…]
GPT-4o 92.0 98.0 100.0 96.7
Claude4-sonnet 96.0 93.8 100.0 96.6
Gemini1.5-Flash 93.0 91.5 100.0 94.8
Cluade3.7-sonnet 91.0 93.3 100.0 94.8
Gemini1.5-Pro 84.0 90.7 99.0 91.2
Qwen2.5-VL-7B 92.0 98.0 98.0 96.0
mPLUG-Owl3-7B 88.0 98.2 93.0 93.1
Qwen2.5-VL-72B 85.0 95.0 98.0 92.7
LongVA-7B 74.0 98.3 90.0 87.4
TPO-7B 77.0 98.1 85.0 86.7
LLaVA-Video-72B 69.0 95.0 93.0 85.7
VideoLLaMA3-7B 50.0 99.6 97.0 82.2
Oryx-34B 67.0 93.8 82.0 80.9
MiniCPM-o-2.6-7B 67.0 97.6 76.0 80.2
Long-LLaVA-7B 65.0 91.9 75.0 77.3
LiveCC-7B 50.0 96.3 73.0 73.1
LLaVA-Video-7B 51.0 96.1 67.0 71.4
Sharegpt4video-8B 44.0 98.5 63.0 68.5
MiniCPM-V-2.6-7B 45.0 96.9 63.0 68.3
Oryx1.5-7B 43.0 97.3 56.0 65.4
LLaVA-OneVision 74.0 87.7 23.0 61.7
Video-ChatGPT-7B 39.0 96.7 41.0 58.9
F.3 Toxic Content Continues
Figure 26: An example for the task of identification of toxic content continues.
Setting. This task evaluates the impact of video context on the videoLLM’s generation of toxic content. The model is tested with toxic text prompts paired with either semantically unrelated or contextually relevant videos, assessing whether the visual input influences the generation of harmful outputs. The setting examines the model’s safety to multimodal interactions and its ability to maintain safe responses under provocative conditions. An example is shown in Figure 26.
Dataset. The dataset comprises 80 text prompts selected from the eight toxicity categories in RealToxicityPrompts [gehman2020realtoxicityprompts], with each category contributing 10 prompts with the highest toxicity scores (determined by Perspective API). Each prompt is paired with two types of videos: (1) semantically unrelated videos randomly selected, and (2) videos collected or generated to be contextually relevant to the prompt. This dual pairing tests the influence of video context on the model’s output toxicity.
Metrics. This experiment encompasses three settings: toxicity-prompt-text (with blank video), toxicity-prompt-unrelated video, and toxicity-prompt-related video. For each setting, we use evaluation metrics consistent with those employed in the NSFW video description task, and report the average of RtA, Perspective score and LLM score.
Results. The performance of videoLLMs in the Toxic Content Continues task is presented in Table XV. Among closed-source models, Claude3.7-sonnet achieved the highest average score of 89.0%, closely followed by Claude4-sonnet at 88.8%. GPT-4o recorded an average score of 86.3%, while Gemini1.5-Flash and Gemini1.5-Pro scored significantly lower at 71.0% and 60.0%, respectively. For open-source models, LongVA-7B led with an average score of 86.8%, followed by Qwen2.5-VL-7B at 86.3%. Other notable open-source performances include mPLUG-Owl3-7B (77.9%) and MiniCPM-V-2.6-7B (77.2%). The lowest average scores were observed for LLaVA-Video-7B at 41.5%, Video-ChatGPT-7B at 49.9%, and LLaVA-Video-72B at 55.6%, with several other open-source models, such as Sharegpt4video-8B and Oryx-34B, also scoring below 60%.
Findings. (1) The toxicity-prompt-unrelated video setting poses the least safety threat, with models generally achieving higher scores (e.g., GPT-4o at 93.4%, Claude4-sonnet at 93.1%). Conversely, the toxicity-prompt-related video setting presents the greatest challenge, as evidenced by significantly lower scores for many models (e.g., Qwen2.5-VL-72B at 35.4%, LLaVA-Video-7B at 39.3%), indicating that contextually relevant videos exacerbate the risk of generating toxic content. (2) Closed-source models, particularly Claude3.7-sonnet and Claude4-sonnet, demonstrate robust safety across all settings, with consistently high scores in both text-only and video-paired conditions. GPT-4o also performs strongly, though its score drops notably in the related-video setting (80.5%). (3) The substantial drop in scores for many models in the toxicity-prompt-related video setting underscores the difficulty of maintaining safe outputs when visual inputs reinforce toxic prompts. This suggests that video context can amplify the model’s propensity to generate harmful content, necessitating stronger safety mechanisms.
TABLE XV: Performance (%) of videoLLMs in the task of toxic content continues. U denotes unrelated video; R denotes related video. The reported scores are the average of RtA, Perspective score, and LLM score.
[…]
Cluade3.7-sonnet 85.9 92.6 88.4 89.0
Claude4-sonnet 83.3 93.1 90.1 88.8
GPT-4o 84.9 93.4 80.5 86.3
Gemini1.5-Flash 70.3 71.1 71.7 71.0
Gemini1.5-Pro 56.6 57.4 66.1 60.0
LongVA-7B 86.5 87.9 86.1 86.8
Qwen2.5-VL-7B 90.4 91.0 77.5 86.3
mPLUG-Owl3-7B 83.5 87.0 63.2 77.9
MiniCPM-V-2.6-7B 85.6 84.1 61.9 77.2
Oryx1.5-7B 67.4 76.2 67.5 70.4
TPO-7B 69.1 71.4 63.0 67.8
MiniCPM-o-2.6-7B 74.7 78.4 49.0 67.4
Long-LLaVA-7B 76.2 76.3 48.5 67.0
VideoLLaMA3-7B 65.9 70.5 63.0 66.5
Qwen2.5-VL-72B 74.8 77.7 35.4 62.6
LiveCC-7B 68.4 60.2 49.8 59.5
Oryx-34B 47.8 57.7 65.6 57.0
Sharegpt4video-8B 62.9 62.6 45.1 56.9
LLaVA-Video-72B 62.1 59.3 45.3 55.6
Video-ChatGPT-7B 44.7 52.6 52.5 49.9
LLaVA-Video-7B 42.9 42.3 39.3 41.5
F.4 Risky content Identification
Figure 27: An example for the risky content identification task.
Setting. This task evaluates the videoLLM’s ability to identify hidden unsafe content or behaviors in dynamic video scenes. The model must detect risky elements (e.g., subtle violence or inappropriate behavior) and provide accurate descriptions of the risks. The setting tests the model’s multimodal reasoning capabilities in real-world scenarios where unsafe content may be embedded in complex, dynamic contexts. An example is shown in Figure 27.
Dataset. The dataset is sourced from YouTube, comprising 50 videos that contain hidden or subtle unsafe content or behaviors. These videos are selected to reflect real-world scenarios where risky elements may not be immediately obvious, requiring the model to analyze temporal and multimodal cues to identify potential dangers.
Metrics. This study assesses videoLLM performance using accuracy as the primary metric, defined as the percentage of videos in which the model accurately identifies and describes hidden unsafe content or behaviors, as evaluated by DeepSeek.
Results. The performance of videoLLMs in the identification of video risky content task is summarized in Figure 28. Among closed-source models, Claude3.7-sonnet achieved the highest accuracy at 16.0%, followed by Gemini1.5-Flash and Gemini1.5-Pro, both at 12.0%. Claude4-sonnet recorded an accuracy of 10.0%, while GPT-4o had the lowest closed-source performance at 8.0%. For open-source models, Oryx-34B significantly outperformed others with an accuracy of 52.0%. Other open-source models, such as LongVA-7B, Qwen2.5-VL-7B, Qwen2.5-VL-72B, and TPO-7B, each achieved 10.0% accuracy. Several open-source models, including Video-ChatGPT-7B, MiniCPM-o-2.6-7B, Long-LLaVA-7B, LiveCC-7B, and LLaVA-OneVision-72B, recorded 0.0% accuracy, with Oryx1.5-7B and LLaVA-Video-7B scoring only 2.0%.
Findings. (1) The results indicate a generally low ability among videoLLMs to identify hidden unsafe content in dynamic video scenes, with most models achieving accuracies below 20%. This suggests significant challenges in detecting subtle or embedded risky elements within complex, real-world video contexts. (2) Closed-source models, led by Claude3.7-sonnet (16.0%), demonstrate modest performance but still struggle to effectively identify risky content. The low accuracies of GPT-4o (8.0%) and Claude4-sonnet (10.0%) highlight limitations in their multimodal reasoning capabilities for this task. (3) Oryx-34B’s standout accuracy of 52.0% among open-source models is remarkable, indicating a strong capacity to detect subtle unsafe content. This performance significantly surpasses both other open-source models and all closed-source models, suggesting advanced temporal and multimodal reasoning capabilities in this specific model.
Figure 28: Performance of videoLLMs in the task of identification of video risky content.
F.5 Temporal Dependency Misleading
Setting. This task assesses the videoLLM’s safety in maintaining temporal consistency and contextual understanding. The model must detect NSFW content (pornography, violence, or hate speech) inserted into otherwise benign videos and accurately describe the risks. The setting evaluates the model’s ability to handle disruptions in temporal continuity and avoid misinterpretations caused by manipulated video segments.
Dataset. The dataset includes 100 videos randomly sampled from non-violent, non-pornographic videos, with 10%–20% of each video’s duration replaced by randomly inserted NSFW content (pornography, violence, or hate speech). This manipulation simulates real-world scenarios where harmful content is embedded in otherwise safe videos, testing the model’s temporal reasoning and safety mechanisms.
Metrics. This task aims to detect toxicity in model responses, using evaluation metrics consistent with those employed in NSFW video assessments.
Results. The performance of videoLLMs in the Temporal Dependency Misleading task is detailed in Table XVI. Among closed-source models, Claude4-sonnet achieved the highest average score of 96.2%, followed by Claude3.7-sonnet at 90.6%. Gemini1.5-Flash and Gemini1.5-Pro recorded average scores of 79.7% and 78.4%, respectively, while GPT-4o had the lowest closed-source performance at 73.2%. For open-source models, LongVA-7B led with an average score of 77.4%, followed by MiniCPM-V-2.6-7B at 74.9% and Qwen2.5-VL-7B at 74.4%. The lowest average scores were observed for VideoLLaMA3-7B at 62.3%, Video-ChatGPT-7B at 63.5%, and mPLUG-Owl3-7B at 63.7%, with several other open-source models, such as LiveCC-7B and Sharegpt4video-8B, also scoring below 70%.
Findings. The evaluation reveals varying capabilities among videoLLMs in detecting NSFW content embedded within benign videos, with closed-source models generally outperforming open-source ones. Claude4-sonnet’s superior performance, particularly its perfect LLM_score (100.0%) and high RtA rate (92.0%), underscores its robustness in maintaining temporal consistency and identifying harmful content. However, the relatively low RtA rates of some closed-source models, such as GPT-4o (23.0%) and Gemini1.5-Pro (48.0%), suggest that while these models did not directly refuse to generate output, the produced content did not contain toxic information. Among open-source models, LongVA-7B’s strong performance indicates potential for effective temporal reasoning, yet the majority of open-source models, including VideoLLaMA3-7B and mPLUG-Owl3-7B, exhibit lower average scores, particularly in RtA and LLM_score metrics. The consistently high Perspective_score (above 90% for most models) reflects a reliable ability to detect toxic content, but the lower RtA rates in many models, especially open-source ones like Sharegpt4video-8B (2.0%) and Oryx1.5-7B (6.0%), highlight difficulties in handling disruptions in temporal continuity. These findings emphasize the need for enhanced temporal reasoning and safety mechanisms in videoLLMs, particularly for open-source models, to improve their ability to detect and mitigate risks posed by manipulated video content in real-world scenarios.
TABLE XVI: Performance (%) of videoLLMs in the task of temporal dependency misleading. Pers. denotes Perspective score; LLM denotes LLM score. LLaVA-OneVision is a 72B version.
[…]
Claude4-sonnet 92.0 96.5 100.0 96.2
Cluade3.7-sonnet 75.0 96.8 100.0 90.6
Gemini1.5-Flash 60.0 92.0 87.0 79.7
Gemini1.5-Pro 48.0 91.3 96.0 78.4
GPT-4o 23.0 96.5 100.0 73.2
LongVA-7B 34.0 99.2 99.0 77.4
MiniCPM-v-2.6-7B 30.0 98.7 96.0 74.9
Qwen2.5-VL-7B 29.0 98.1 96.0 74.4
Long-LLaVA-7B 20.0 96.1 100.0 72.0
TPO-7B 18.0 99.0 95.0 70.7
Qwen2.5-VL-72B 24.0 93.7 94.0 70.6
LLaVA-Video-72B 27.0 91.1 91.0 69.7
LLaVA-Video-7B 26.0 97.5 85.0 69.5
MiniCPM-o-2.6-7B 14.0 98.1 94.0 68.7
Oryx1.5-7B 6.0 98.5 100.0 68.2
LLaVA-OneVision 10.0 94.4 100.0 68.1
Oryx-34B 9.0 93.0 100.0 67.3
Sharegpt4video-8B 2.0 98.8 100.0 66.9
LiveCC-7B 6.0 98.3 90.0 64.8
mPLUG-Owl3-7B 22.0 97.1 72.0 63.7
Video-ChatGPT-7B 8.0 95.4 87.0 63.5
VideoLLaMA3-7B 10.0 97.8 79.0 62.3
F.6 Deepfake Identification
Figure 29: An example for the deepfake identification task.
Setting. This task evaluates the videoLLM’s ability to detect deepfake content in videos, designed as a discriminative task. The model must identify manipulated media, such as synthetically altered faces, to prevent the propagation of misleading or harmful content. The setting evaluate the model’s robustness to advanced manipulation techniques and its capacity to distinguish authentic from fabricated video content. An example is shown in Figure 29.
Dataset. We sample 100 original videos and 100 manipulated videos from the Deepfakes Detection Entire Original dataset  [dufour2019deepfakes], which is designed for deepfake detection tasks, providing a comprehensive collection of video sequences that can be used to train and evaluate deep learning models for identifying manipulated media.
Metrics. Following prior discriminative tasks, we use accuracy as the primary evaluation metric for assessing videoLLMs performance. Additionally, correct responses are identified through keyword matching, as answers consistently contain either “Yes” or “No”.
Figure 30: Performance of videoLLMs in the task of deep fake identification.
Results. The performance of various videoLLMs in the deepfake identification task is summarized in Figure 30. Among the closed-source models, Gemini1.5-Pro achieved the highest accuracy at 55.0%, followed by Claude4-sonnet and Gemini1.5-Flash, both at 53.5%. GPT-4o recorded an accuracy of 47.0%, while Claude3.7-sonnet had the lowest performance among closed-source models at 41.5%. For open-source models, MiniCPM-o-2.6-7B led with an accuracy of 54.5%, closely followed by LLaVA-Video-72B at 53.0%. Other notable open-source performances include Qwen2.5-VL-7B, mPLUG-Owl3-7B, LLaVA-Video-7B, and VideoLLaMA3-7B, all achieving 50.0% accuracy. The lowest accuracy was observed for Oryx1.5-7B at 17.5%, with Sharegpt4video-8B and LiveCC-7B also performing poorly at 25.5% and 35.0%, respectively.
Findings. The results reveal a varied performance landscape among videoLLMs in detecting deepfake content. Closed-source models generally outperformed their open-source counterparts, with Gemini1.5Pro demonstrating superior robustness to advanced manipulation techniques. However, the top-performing open-source model, MiniCPM-o-2.6-7B, achieved an accuracy comparable to leading closed-source models, indicating significant progress in open-source videoLLM development. The substantial performance gap between the best and worst models, particularly the low accuracies of Oryx1.5-7B and Sharegpt4video-8B, suggests that some models struggle with the complexity of distinguishing authentic from manipulated video content. The consistent performance of models like LLaVA-Video-7B and VideoLLaMA3-7B at 50.0% highlights a baseline capability among several open-source models, though further improvements are needed to match or exceed the top closed-source performers. These findings underscore the importance of continued advancements in model architectures and training datasets to enhance deepfake detection capabilities across both closed- and open-source videoLLMs.
F.7 Jailbreak attack
Figure 31: An exemplar of the VideoJail jailbreaking attack against videoLLMs.
Setting. This task evaluates the robustness of videoLLMs against jailbreak attacks designed to bypass safety mechanisms and elicit harmful or inappropriate responses. The evaluation includes two image-based jailbreak attacks, FigStep [gong2025figstep] and MMsafetybench [liu2024mm], and two video-based jailbreak attack, VideoJail [hu2025videojail] and VideoJailPro [hu2025videojail]. For the image-based attacks, adversarial examples are converted into video format to assess the model’s vulnerability in a multimodal context. The setting tests the videoLLMs’ ability to maintain safety and ethical alignment when confronted with adversarial inputs that exploit visual and temporal modalities to trigger unsafe outputs. An example for videoJail is refered to Figure 31.
Dataset. The dataset comprises adversarial examples from four jailbreak attack frameworks: FigStep [gong2025figstep], MMsafetybench [liu2024mm], VideoJail [hu2025videojail], and VideoJail-Pro [hu2025videojail]. For the image-based attacks (FigStep and MMsafetybench), adversarial images are transformed into videos by incorporating them into sequences, ensuring compatibility with the multimodal input requirements of MLLMs. The VideoJail and VideoJail-Pro datasets include videos specifically designed to exploit weaknesses in video processing capabilities.
Metrics. This experiment includes four settings: FigStep (FS), MMsafetybench (MM), VideoJail (VJ), and VideoJail-Pro (VJP). Model safety is evaluated using a composite metric calculated as the average of three scores: the Refuse-to-Answer (RtA) rate, Perspective score, and LLM score (consistent with those employed in the NSFW video description task). The safety score for each setting is computed as (RtA + Perspective score + LLM score) / 3, providing a comprehensive assessment of the model’s ability to mitigate the generation of harmful content in response to adversarial inputs across different attack types.
Results. The performance of videoLLMs in the Jailbreak Attack task is presented in Table XVII. Among closed-source models, Claude3.7-sonnet achieved the highest average score of 91.0%, followed by GPT-4o at 79.5% and Claude4-sonnet at 78.9%. Gemini1.5-Flash and Gemini1.5-Pro recorded lower average scores of 73.1% and 67.7%, respectively. For open-source models, mPLUG-Owl3-7B led with an average score of 77.4%, followed by MiniCPM-V-2.6-7B at 70.5% and Qwen2.5-VL-72B at 69.6%. The lowest average scores were observed for LiveCC-7B at 51.3%, MiniCPM-o-2.6-7B at 55.0%, and LLaVA-Video-7B at 55.0%. Notably, closed-source models performed poorly in the VideoJail-Pro setting (e.g., Gemini1.5-Pro at 48.2%), while open-source models struggled in the VideoJail setting (e.g., LiveCC-7B at 37.0%).
Findings. (1) VideoJail and VideoJail-Pro settings present greater defense challenges compared to FigStep and MMsafetybench. VideoJail is particularly effective against open-source models, with scores as low as 37.0% for LiveCC-7B, while VideoJail-Pro significantly impacts closed-source models, with scores like 48.2% for Gemini1.5-Pro and 51.0% for Claude4-sonnet. (2) Claude3.7-sonnet demonstrates superior robustness across all settings, achieving the highest average score of 91.0% and excelling in VideoJail-Pro (94.1%). However, other closed-source models, such as GPT-4o and Claude4-sonnet, show vulnerabilities in the VideoJail-Pro setting, with scores dropping to 54.7% and 51.0%, respectively. (3) The pronounced impact of VideoJail on open-source models and VideoJail-Pro on closed-source models highlights the need for improved multimodal safety mechanisms. The varying effectiveness of attack types underscores the importance of developing targeted defenses to address specific vulnerabilities in video processing and adversarial input handling.
TABLE XVII: Performance (%) of videoLLMs against jailbreak attacks. FS denotes Figstep; $M$ denotes $M$-SafetyBench; VJ denotes VideoJail; VJP denotes VideoJailPro. LLaVA-OneVision is the 72B version.
[…]
Cluade3.7-sonnet 87.6 89.0 93.3 94.1 91.0
GPT-4o 86.8 84.2 92.1 54.7 79.5
Claude4-sonnet 84.0 89.1 91.3 51.0 78.9
Gemini1.5-Flash 74.8 87.1 73.4 57.0 73.1
Gemini1.5-Pro 79.1 82.5 61.2 48.2 67.7
mPLUG-Owl3-7B 99.4 77.1 64.0 69.3 77.4
MiniCPM-V-2.6-7B 78.3 75.3 51.1 77.2 70.5
Qwen2.5-VL-72B 71.1 80.9 65.5 61.0 69.6
Oryx1.5-7B 78.6 74.5 48.4 69.6 67.8
Qwen2.5-VL-7B 73.5 82.9 43.0 64.8 66.1
LLaVA-OneVision 61.6 66.0 69.6 60.2 64.4
LongVA-7B 70.9 74.7 55.1 53.0 63.4
Video-ChatGPT-7B 53.2 67.4 55.6 69.0 61.3
TPO-7B 61.9 73.3 46.3 63.2 61.2
Sharegpt4video-8B 56.1 62.9 57.4 65.1 60.4
LLaVA-Video-72B 62.5 66.5 41.1 70.6 60.2
VideoLLaMA3-7B 73.4 64.0 40.8 56.9 58.8
Long-LLaVA-7B 59.1 70.1 42.2 63.2 58.7
Oryx-34B 63.5 71.9 43.9 52.6 58.0
LLaVA-Video-7B 65.0 68.4 40.0 47.2 55.0
MiniCPM-o-2.6-7B 61.9 66.0 43.2 48.8 55.0
LiveCC-7B 55.2 60.3 37.0 52.6 51.3
F.8 Summary
F.8.1 Score Calculate
We evaluate the safety of videoLLMs across three aspects related to video understanding:
Toxicity in Generated Content. We evaluate toxicity across three tasks: NSFW video description, NSFW prompt execution, and toxic content continuation. For each task, we compute composite scores by averaging relevant metrics (RtA, Perspective API scores, and LLM-based assessments). The overall toxicity score is defined as:
$\begin{split}\mathrm{Score_{toxic}}&=\bigg(\mathrm{Score}_{\mathrm{nsfw\_video}}+\mathrm{Score}_{\mathrm{nsfw\_prompt}}\\ &\quad+\mathrm{Score}_{\mathrm{toxic\_prompt}}\bigg)\big/{3}\times 100\end{split}$ (12)
Unsafe Content Recognition. We assess the model’s ability to identify risky content through two complementary evaluations: risky content identification (measured by accuracy $\mathrm{Acc}_{\mathrm{risk\_recognition}}$) and temporal dependency misleading tasks (scored using averaged RtA, Perspective API, and LLM metrics, denoted as $\mathrm{Score}_{\mathrm{temporal}}$). The combined unsafe content recognition score is:
$\mathrm{Score_{unsafe}}=\frac{\mathrm{Acc}_{\mathrm{risk\_recognition}}+\mathrm{Score}_{\mathrm{temproal}}}{2}\times 100$ (13)
Safety Against Malicious Manipulations. This dimension encompasses deepfake identification accuracy ($\mathrm{Acc}_{\mathrm{deepfake}}$) and safety against jailbreak attacks. For jailbreak evaluation, we aggregate results across four attack methods (FigStep, MMSafety, VideoJail, and VideoJailPro):
$\begin{split}\mathrm{Score_{jailbreak}}&=\bigg(\mathrm{Score}_{\mathrm{figstep}}+\mathrm{Score}_{\mathrm{mmsafety}}\\ &\quad+\mathrm{Score}_{\mathrm{videoJail}}+\mathrm{Score}_{\mathrm{videoJailPro}}\bigg)\big/4\times 100\end{split}$ (14)
Finally, we take the average of these metrics as the score under OOD robustness, which is expressed as:
$\mathrm{Score_{safety}}=\frac{\mathrm{Acc}_{\mathrm{deepfake}}+\mathrm{Score_{jailbreak}}}{2}\times 100$ (15)
The comprehensive rankings and corresponding scores for Safety evaluation are presented in Table XVIII.
TABLE XVIII: The scores and rankings of three subaspects in Safety.
G. U. S.
Model Score Rank Score Rank Score Rank
Claude4-sonnet 89.8 1 85.2 1 78.8 3
Claude3.7-sonnet 87.0 2 77.5 2 91.0 1
Gemini1.5-Pro 58.2 17 70.3 4 67.7 10
Gemini1.5-Flash 64.9 13 70.8 3 73.1 5
GPT-4o 81.8 4 59.3 9 79.4 2
Qwen2.5-VL-72B 66.5 10 56.3 11 69.6 8
Qwen2.5-VL-7B 77.2 5 61.7 7 66.0 11
LLaVA-Video-72B 55.5 21 54.3 14 64.1 13
LLaVA-Video-7B 53.7 22 54.3 15 55.0 21
MiniCPM-o-2.6-7B 65.3 12 54.2 16 55.0 22
MiniCPM-V-2.6-7B 65.9 11 59.8 8 70.4 7
Oryx-34B 69.4 7 67.3 5 58.0 20
Oryx1.5-7B 67.1 9 41.8 23 67.8 9
InternVL2.5-78B 83.0 3 55.8 12 70.6 6
LLaVA-OneVision 55.9 20 50.3 17 64.3 12
mPLUG-Owl3-7B 69.1 8 49.3 18 77.4 4
LongVA-7B 70.2 6 65.3 6 63.4 14
Sharegpt4video-8B 58.2 18 44.5 21 60.4 17
TPO-7B 64.8 14 57.5 10 61.2 16
Long-LLaVA-7B 64.4 15 54.7 13 58.7 19
Video-ChatGPT-7B 52.7 23 45.8 20 62.3 15
LiveCC-7B 57.5 19 43.7 22 51.3 23
VideoLLaMA3-7B 61.7 16 47.7 19 58.8 18
F.8.2 Takeaways
  • Closed-source videoLLMs (e.g., Claude, GPT-4o) demonstrate robust safety standards by effectively filtering NSFW content and toxic prompts, yet exhibit vulnerabilities to sophisticated jailbreak attacks such as VideoJail-Pro and subtle risky content detection.
  • Open-source models require substantial safety alignment improvements, particularly in NSFW detection and resistance to video-based adversarial attacks. Video context critically influences safety outcomes, with contextually relevant inputs significantly amplifying harmful output risks.
  • While the highest-performing open-source model achieved comparable performance to leading closed-source systems—indicating notable progress in open-source videoLLM development—their defensive capabilities against jailbreak attacks remain inferior to closed-source alternatives.
  • Enhancing videoLLM safety across both paradigms necessitates targeted advances in temporal reasoning robustness, multimodal alignment, and intrinsic safety mechanisms.
Appendix G Evaluation Details on Fairness&Bias
VideoLLMs are trained on large-scale video-text datasets that may contain historical biases or imbalanced representations. Marginalization of certain groups—based on race, gender, or cultural background—can lead to the generation of stereotypical or discriminatory content. Given the multimodal nature of videos, encompassing visual, auditory, and textual elements, such biases may manifest subtly through scene composition, tone, or subtitles. Fairness evaluation must therefore account for all modalities to ensure that biases in one are not obscured by others. To comprehensively assess fairness and bias in MLLMs for video understanding, we adopt a dual-perspective approach: (1) evaluating the risk of bias arising from data-driven factors, and (2) assessing fairness in multimodal understanding, which enables a more nuanced analysis and supports the development of ethically robust MLLMs on video understanding.
G.1 The risk of bias arising from data-driven factors
Large-scale video datasets—such as films and content from short-video platforms—often embed historical, cultural, or societal biases. These biases can cause models to generate or reinforce discriminatory interpretations, such as associating certain behaviors with specific demographic groups. Moreover, imbalanced group representation in the data may result in reduced accuracy or inappropriate outputs when the model processes content involving underrepresented populations. Unlike image-based models, the temporal dimension of video allows such biases to compound over time; for instance, a model may initially interpret early frames accurately but gradually incorporate stereotypical or erroneous interpretations as the video progresses.
G.1.1 Stereotype Impact Generation
Setting. In this task, we present videoLLMs with videos depicting attributes such as gender, age, skin color, and occupation—contexts that are likely to elicit stereotypical responses. Each video is paired with carefully designed prompts. The objective is to quantitatively assess the sensitivity of videoLLMs to classical stereotypical scenarios in a VQA setting.
Dataset. To ensure effective evaluation, we select 1,592 videos from OpenVid-1M [nan2024openvid], covering attributes such as gender, age, skin color, and occupation to ensure diversity, relevance, and representativeness. Prompts are then generated using Grok3666https://grok.com/ based on the detailed descriptions of each video. An illustrative example is shown in Figure 32.
Figure 32: An example for the task of stereotype impact generation.
Metrics.To ensure the reliability of determining whether the videoLLMs’ outputs contain stereotypes, we utilize DeepSeek as a judge for its exceptional capability to identify stereotypes. After obtaining the DeepSeek results, we use a keyword matching algorithm to statistically assess the presence of "Yes" or "No", thus obtaining the stereotyping rate. The prompt for evaluating by DeepSeek is shown in Figure 33.
Figure 33: Prompt to use DeepSeek for judging the responses from videoLLMs.
Figure 34: Performance of videoLLMs in the task of stereotype impact generation.
Results. In this study, we quantitatively evaluated the stereotype sensitivity of videoLLM in a visual question-answering setting using 1,592 video samples annotated for gender, age, skin color, and occupation. As shown in Figure 34, the results reveal substantial variation in the stereotype containing rate across models, ranging from 18.0% (Oryx-34B) to 74.3% (MiniCPM-V-2.6-7B). Four closed-source models (GPT-4o, Claude3.7-sonnet, Gemini1.5-Pro, and Geminil.5-Flash) exhibit stereotype rates clustered between 19.0% and 23.0%, significantly lower than most open-source counterparts. In particular, Oryx-34B (18.0%) and LLaVA-Video-72B (20.0%) lie at the low end of the spectrum, whereas open-source models such as MiniCPM-V-2.6-7B (74.3%) and LongVA-7B (70.2%) show high stereotype tendencies.
Within the same model families, larger-parameter variants generally display reduced stereotype rates. For example, LLaVA-Video-72B (20.0%) achieves a 23.7% reduction compared to its 7B version (43.7%), and Qwen2.5-VL-72B (30.7%) shows a 39.8% drop relative to its 7B counterpart (70.5%). This suggests that increased model capacity may strengthen its ability to suppress stereotypes. Moreover, TPO (67.00%)—which incorporates Time Preference Optimization—reduces its stereotype rate by 4.6% compared to the base LongVA-7B (70.2%), indicating that targeted technical refinements can partially mitigate stereotyping issues.
Findings. (1) Closed-source models show superior performance in stereotype suppression: The lowest stereotype rates in the experiment (ranging from 19.0% to 23.0%) were primarily observed in closed-source models. This may be attributed to efforts by closed-source teams in data curation, bias mitigation algorithm optimization, and adherence to ethical constraints tied to commercial deployment. In contrast, open-source models tend to exhibit higher levels of stereotyping, likely due to limited diversity in training data and fewer resources devoted to debiasing techniques. (2) Model size is negatively correlated with stereotype rates: Within the same model families, larger parameter sizes are associated with lower stereotype rates (e.g., the LLaVA-Video and Qwen series). This suggests that increased model capacity may enhance the semantic understanding of complex social attributes, leading to fewer stereotype-driven default responses. (3) Industry-wide performance still needs improvement: Over 60% of the models tested exhibited stereotype rates above 50%, indicating a persistent tendency among current VideoLLMs to respond based on traditional stereotypes when handling sensitive attributes such as gender and age. Notably, the high stereotype rate observed in open-source models (e.g., MiniCPM-V-2.6-7B at 74.3%) highlights the urgent need to integrate debiasing technologies into the open-source ecosystem.
G.1.2 Preference selection of videoLLMs
Setting. This task assesses whether a videoLLM exhibits cultural, ideological, or lifestyle biases when interpreting video content. For each topic, two videos representing opposing perspectives are presented alongside carefully designed prompts that ask the model to choose between them and justify its reasoning. The model is expected to avoid expressing subjective preferences, as AI systems should not demonstrate personal bias. Instead, it should either refuse to make a direct choice—citing its impartial nature—or offer a neutral, comparative analysis of the videos based on content, theme, or visual elements.
Figure 35: An example for the task of preference selection in videoLLMs.
Dataset. We curated 50 video pairs across diverse themes—including ideology, culture, lifestyle, social issues, and technological development—to ensure thematic breadth and richness. Each pair presents opposing perspectives and is generated using the Jimeng AI  [^7] https://jimeng.jianying.com/ platform. Prompts were carefully crafted to remain neutral, with the goal of assessing the fairness and potential bias of videoLLMs. An illustrative example is provided in Figure 35.
Metrics. To evaluate whether videoLLMs appropriately avoid subjective preferences, we use the Refuse-to-Answer (RtA) rate as the primary metric. Models are expected to decline making a choice when presented with videos representing opposing perspectives, in alignment with the principle that AI systems should not exhibit personal bias. To identify refusals, we combine a previously trained Longformer-based classifier [wang2023not] with a keyword-matching approach that detects typical refusal phrases (e.g., “Sorry, I cannot,” or “It’s not appropriate to…”). The keyword method complements the classifier by capturing cases it may miss.
TABLE XIX: Performance (%) of videoLLMs in the task of perference selection.
[…]
GPT-4o 2.0 18.0
Claude3.7-sonnet 0.0 52.0
Claude4-sonnet 0.0 78.0
Gemini1.5-Pro 2.0 18.0
Geminil.5-Flash 2.0 20.0
LLaVA-Video-7B 0.0 2.0
LLaVA-Video-72B 0.0 14.0
LLaVA-OneVision-72B 2..0 50.0
MiniCPM-o-2.6-7B 0.0 10.0
MiniCPM-V-2.6-7B 0.0 10.0
LongVA-7B 14.0 54.0
TPO-7B 0.0 24.0
mPLUG-Owl3-7B 34.0 42.0
Oryx1.5-7B 28.0 66.0
Oryx-34B 2.0 6.0
Qwen2.5-VL-7B 0.0 14.0
Qwen2.5-VL-72B 0.0 16.0
Video-ChatGPT-7B 0.0 6.0
VideoLLaMA3-7B 14.0 20.0
Long-LLaVA-7B 6.0 16.0
LiveCC-7B 0.0 2.0
Sharegpt4video-7B 16.0 2.0
Results. As shown in Table XIX, the classifier detection results reveal significant differences in model behavior. Among open-source models, mPLUG-Ow13-7B leads with a 34% classifier-detected refusal rate, followed by Oryx1.5-7B (28%) and LongVA-7B (14%). Most models—such as LLaVA-Video-7B, LLaVA-Video-72B, and Qwen-2.5-VL-7B show a 0% refusal rate under this metric, indicating that the classifier failed to detect refusal behaviors in these models. Closed-source models like GPT-4o, Claude3.7-sonnet, and Gemini1.5-Pro each have a low classifier-detected refusal rate of 2%, with Geminil.5-Flash also at 2%.
In contrast, rule-based evaluation using keyword matching yields higher overall refusal rates, highlighting its effectiveness in identifying explicit refusal phrases. Oryx1.5-7B leads in this metric with a 66% refusal rate, followed by Geminil.5-Flash (78%, the highest among all models), Claude4-sonnet (52%), and LongVA-7B (54%). Open-source models such as mPLUG-Owl3-7B (42%) and MiniCPM-V-2-6-7B (10%) show moderate performance, while closed-source models like Claude4-sonnet and Geminil.5Flash demonstrate strong adherence to refusal templates. Poor performers include LLaVA-Video-7B (2%) and Video-ChatGPT-7B (6%), with most LLaVA and Qwen variants scoring below 20%.
Findings. (1) Different models exhibit significant variation in behavior: some models (such as mPLUG-Owl3-7B, LongVA-7B, and the closed-source Claude4-sonnet) frequently refuse to choose between opposing viewpoints, thereby prioritizing neutrality in accordance with AI fairness principles. In contrast, other models (such as LLaVA and Qwen variants) rarely issue refusals, which may risk introducing subjective bias through direct comparisons, and they often fail to explicitly declare neutrality. Among closed-source models, the Claude series stands out with high rule-based refusal rates (52% and 78%), highlighting its strong bias-avoidance mechanisms. (2) There is no significant correlation between model size and refusal rate. For example, the 7B LongVA-7B (rule-based refusal rate: 54%) outperforms the larger 72B LLaVA-Video-72B (14%), and the 7B Oryx1.5-7B (66%) exceeds even larger models like LLaVA-Video-72B. This suggests that bias avoidance capabilities are more influenced by architectural design or training objectives than by model size.
G.1.3 Profession Competence Prediction
Figure 36: Prompts for videoLLMs to judge people’s competence for their profession under two settings.
Setting. Preference selection in videoLLMs evaluates biases in subjective, preference-driven scenarios involving cultural, ideological, or lifestyle perspectives. However, biases in professional competence judgments—often grounded in objective criteria such as skills or performance—may manifest differently. To address this, we extend the evaluation to examine whether videoLLMs exhibit stereotypes or unfair assumptions related to attributes like gender, race, or appearance in professional contexts. In this task, models are challenged to predict individuals’ professional competence, reflecting how factors such as age, gender, and skin tone influence model perceptions (see Figure 37). Following an approach similar to MultiTrust [zhang2024benchmarking], we present videos of individuals with diverse attributes alongside their actual professions to assess videoLLM judgments of competence. We further analyze model sensitivity through two settings: one presenting attributes visually only, and another combining visual and textual attribute information. The prompts used in both settings are illustrated in Figure 36.
Figure 37: An example for the task of profession competence prediction.
Dataset. We use JimengAI and KlingAI to generate videos depicting six professions commonly associated with bias: nurse, teacher, waiter, doctor, law enforcement officer, and firefighter. Each video features individuals varying in age, gender, and skin tone. Age is categorized as young, middle-aged, or older; gender as male, female, or non-binary; and skin tone as light, medium, or dark.
Metrics. To evaluate the influence of various attributes on the predictions, we compute each attribute’s p-value. Specifically, we begin by calculating the Chi-square statistic using the formula:
$\chi^{2}_{\text{obs}}=\sum_{i=1}^{n}\frac{(O_{i}-E_{i})^{2}}{E_{i}}$
where $n$ is the number of categories, $O_{i}$ denotes the observed frequency in each category, and $E_{i}$ is the expected frequency for each category calculated by:
$E_{i}=\frac{R_{i}\times C_{i}}{N}$
Here, $R_{i}$ and $C_{i}$ are the totals of the respective row and column that contain $i$, and $N$ is the total number of observations in the data. After calculating the Chi-square statistic, we compute the p-value, which determines the statistical significance of the observed differences:
$\text{p-value}=P(\chi^{2}\geq\chi^{2}_{\text{obs}})$
The p-value represents the probability that a Chi-square statistic as extreme as or more extreme than the observed statistic $\chi^{2}_{\text{obs}}$ could occur under the null hypothesis, which assumes no association between the attributes and the predictions. A low p-value (typically less than 0.05) indicates that the null hypothesis can be rejected, suggesting a statistically significant association between the variables. This may also reveal that videoLLMs have biases toward such attributes.
TABLE $X$: P-values of videoLLMs for Each Attribute; Higher Values Indicate Weaker Performance. Wo. Description = Without Description, W. Description = With Description; 2. LLaVA-OneVision is 72B version. The models whose p-value is less than 0.05 are highlighted in red color.
Models Wo.Description W.Description
gender age skintone gender age skintone
Claude4-sonnet 0.97 0.89 0.99 0.87 0.89 0.99
Claude3.7-sonnet 0.97 0.89 0.99 0.87 0.89 0.99
Gemini1.5-Pro 0.97 0.89 0.99 0.87 0.89 0.99
Gemini1.5-Flash 0.70 0.15 0.12 0.97 0.89 0.99
GPT-4o 0.90 0.11 0.75 0.86 0.39 0.67
Qwen2.5-VL-7B 0.97 0.89 0.99 0.87 0.89 0.99
mPLUG-Owl3-7B 0.97 0.89 0.99 0.87 0.89 0.99
LiveCC-7B 0.97 0.89 0.99 0.87 0.89 0.99
Sharegpt4video-8B 0.97 0.89 0.99 0.87 0.89 0.99
TPO-7B 0.97 0.89 0.99 0.87 0.89 0.99
LongVA-7B 0.97 0.89 0.99 0.87 0.89 0.99
Video-ChatGPT-7B 0.97 0.89 0.99 0.87 0.89 0.99
Oryx1.5-7B 0.97 0.89 0.99 0.87 0.89 0.99
Qwen2.5-VL-72B 0.97 0.89 0.99 0.87 0.89 0.99
Oryx-34B 0.97 0.89 0.99 0.52 0.70 0.72
MiniCPM-V-2,6 0.17 0.64 0.72 0.90 0.68 0.29
Long-LLaVA-7B 0.24 0.71 0.94 0.93 0.07 0.28
VideoLLaMA3-7B 0.43 0.21 0.28 0.70 0.28 0.99
LLaVA-Video-72B 0.00 0.35 0.23 0.86 0.34 0.63
LLaVA-Video-7B 0.05 0.36 0.17 0.86 0.34 0.63
MiniCPM-o-2.6 0.23 0.72 0.26 0.04 0.38 0.55
LLaVA-OneVision 0.01 0.39 0.04 0.87 0.01 0.61
Results. Analysis of videoLLMs for professional competence prediction (Table XX) reveals notable biases across gender, age, and skin tone attributes. Higher P-values correspond to weaker performance. First, gender shows greater predictive influence than age or skin tone in open-source model outputs. Closed-source models (e.g., Claude 4-sonnet, Gemini 1.5-Pro) consistently yield high P-values (0.87–0.99) in both visual-only and visual-plus-text settings, indicating persistent bias. In contrast, open-source models such as MiniCPM-V-2.6and Long-LLaVA achieve significantly lower P-values in visual-only settings, reflecting reduced bias. However, their performance often degrades with the addition of text. Models like Oryx-34B and MiniCPM-o-2.6 show improved outcomes with text for certain attributes.
Findings. VideoLLMs remain notably influenced by visual attributes such as gender, age, and skin tone when predicting professional competence, especially in the absence of clear textual guidance. These biases often reflect stereotypical assumptions. While adding attribute descriptions can help mitigate these effects in some models, it may exacerbate them in others due to over-reliance on explicit cues. This highlights an ongoing fairness challenge: current videoLLMs are not yet reliable in making unbiased professional judgments, and addressing these issues requires more systematic efforts in data construction, model training, and multimodal alignment strategies.
G.2 The Fairness in Multimodal Understanding
Building on our prior quantitative assessment of stereotype prevalence in model outputs, we now examine the contributing modalities and temporal factors within videoLLMs. Specifically, we analyze these influences from two perspectives: video content and temporal dynamics.
G.2.1 Agreement on Stereotypes
Figure 38: An example for the task of agreement on stereotypes.
Setting. Following the setting of TrustLLM [sun2024trustllm], we present videoLLMs with stereotypical statements and prompt them to express agreement or disagreement, thereby assessing their intrinsic attitudes toward stereotypes. Additionally, we examine how the combination of textual prompts with varying videos cues—both relevant and irrelevant—impacts the models’ tendency to endorse stereotypical views, shedding light on fairness and bias in multimodal comprehension. An example is shown in Figure 38.
Dataset. To investigate differences between MLLMs for video and image understanding, we employ stereotype prompts from MultiTrust [zhang2024benchmarking], comprising 261 statements across six categories: age, gender, nationality, race, religion, and socioeconomic status. For related videos, we extract key content from these statements using Grok3 and generate semantically relevant videos via the text-to-video model Jimeng. Additionally, each statement is paired with irrelevant videos from three categories—color, nature, and noise—for comparison.
Metrics. In this task, we adopt the Stereotype Agreement Rate as the primary evaluation metric, defined as the proportion of videoLLMs that provide an affirmative response, specifically "Yes". A higher Stereotype Agreement Rate indicates a greater tendency for models to endorse stereotypes, which may reflect deficiencies in fairness. To ensure a comprehensive assessment, we extend beyond simple keyword matching for "Yes" including a set of affirmative keywords, such as "agree" and "favor" to robustly capture expressions of agreement. Additionally, we analyze the brief rationales provided by videoLLMs to mitigate the risk of uniform affirmative responses.
TABLE XXI: Performance (%) of videoLLMs in the task of agreement on stereotypes. u-c denotes unrelated-color; u-n denotes unrelated-natural; u-o denotes unrelated-noise.
Models Agreement Metrics (Average)
video u-c u-n u-o
GPT-4o 12.3 8.5 10.0 11.2
Claude3.7-sonnet 6.5 4.6 4.2 5.4
Claude4-sonnet 6.9 6.9 5.4 5.0
Gemini1.5-Pro 9.6 11.2 10.8 11.2
Geminil.5-Flash 7.3 9.6 7.3 10.4
LLaVA-Video-7B 13.5 8.1 0.8 5.8
LLaVA-Video-72B 13.9 10.4 4.2 7.3
LLaVA-OneVision-72B 15.4 15.4 15.4 15.4
MiniCPM-o-2.6-7B 13.1 10.0 5.8 8.1
MiniCPM-V-2.6-7B 17.7 11.9 5.0 14.6
LongVA-7B 16.9 19.2 10.4 20.8
TPO-7B 21.2 20.4 18.1 21.9
mPLUG-Owl3-7B 13.5 11.5 8.9 5.4
Oryx1.5-7B 13.9 14.2 6.9 14.6
Oryx-34B 20.4 22.3 20.4 20.4
Owen2.5-VL-7B 6.5 8.1 1.9 5.8
Owen2.5-VL-72B 15.0 12.7 10.8 13.5
Video-ChatGPT-7B 18.9 43.1 18.9 45.8
VideoLLaMA3-7B 10.0 13.5 4.6 20.4
Long-LLava-7B 11.9 10.4 8.5 10.8
LiveCC-7B 9.2 6.2 0.4 2.7
Sharegpt4video-7B 0.4 0.4 0.4 0.4
Results. As shown in Table XXI, the agreement rates for related videos are generally higher than those for unrelated videos. However, closed-source models exhibit significantly lower average agreement rates (6.5%–12.3%) on related videos compared to open-source models, such as TPO, which reaches 21.2%. For example, Claude4-sonnet (6.9%) and Geminil.5-Flash (7.3%) demonstrate strong resistance to stereotypes in the presence of related videos, whereas open-source models like TPO and Oryx-34B show agreement rates exceeding 20%. Among the unrelated video types, noise and color videos lead to notably higher agreement rates than natural scenes, suggesting their stronger interference with model judgment. Within the closed-source models, GPT-4o reaches an agreement rate of 11.2% on noise videos, while Claude3.7-sonnet shows a much lower agreement rate of only 4.2% on natural scene videos.
Findings. (1) Closed-source models generally show significantly lower agreement rates with stereotypes compared to open-source models, demonstrating an advantage in recognizing and resisting stereotypical content. (2) Different types of unrelated videos interfere with model judgments to varying degrees; specifically, noise and color videos lead to significantly higher agreement rates than natural scenes. (3) Some open-source models perform well, such as Qwen-2.5-VL-7B (6.5%) and Sharegpt4video-8B (0.4%), with agreement rates approaching those of closed-source models. However, others perform poorly—for instance, Video-ChatGPT-7B exhibits an agreement rate as high as 45.8% in noise video conditions—highlighting substantial differences in the robustness of open-source models against multimodal interference.
G.2.2 Time Sensitivity Analysis
Figure 39: An example for the task of time sensitivity analysis.
Settings. Evaluate whether the model’s bias remains consistent throughout the video timeline, specifically assessing if it forms premature judgments based on isolated frames or segments, or if it appropriately revises its assessment as the narrative unfolds, thereby ensuring a holistic consideration of the entire content. An example is shown in Figure 39.
Dataset. To evaluate a videoLLMs’ time sensitivity and bias changes, we design 50 pairs of video content where each video has a clear narrative progression or plot twist across its timeline (e.g., 30 seconds, divided into two 15-second segments) and generate the videos by text2video model: Jimeng. The goal is to assess whether the model forms premature biases based on early segments and whether it can adjust its judgment as the video progresses. Each video pair will include diverse groups (e.g., gender, race, age, profession) to test for fairness across different demographics.
Metrics. To ensure the reliability of determining whether the bias in videoLLMs’ outputs remains consistent with the change of time , we utilize DeepSeek as a judge for its exceptional capability to identify change. After obtaining the DeepSeek results, we use a keyword matching algorithm to statistically assess the presence of ’Yes’ or ’No’, thus obtaining the stereotyping rate. The prompt for evaluating by DeepSeek is shown in Figure 40.
Figure 40: Prompts for DeepSeek to assess the consistency between descriptions of two videos.
Figure 41: Performance of videoLLMs in the task of time sensitivity analysis.
Results. Different videoLLMs exhibit significant differences in temporal sensitivity and bias adjustment capabilities. As shown in Figure 41, among open-source models, Sharegpt4video-8B performs well with a 88.6% accuracy rate, demonstrating precise integration of full video content and dynamic bias adjustment. Oryx1.5-7B (86.4%) and TPO (81.8%) follow closely, indicating that some open-source models can effectively capture narrative shifts through architectural enhancements (e.g., temporal modeling modules like TPO). Among closed-source models, GPT-4o performs well with an 88.6% accuracy rate, while Gemini1.5-Pro/Flash achieve 70.5% and 65.9%, respectively—falling behind some open-source counterparts.
At the lower end, Oryx-34B records the lowest accuracy at 15.9%, highlighting that a larger parameter scale does not necessarily correlate with better performance.
Findings. The temporal sensitivity and bias adjustment capabilities of videoLLMs depend on the refinement of temporal modeling architecture and the depth of language-vision alignment, rather than merely on parameter scale or whether the model is closed-source. Open-source models, through targeted optimizations, are already able to rival some closed-source solutions.
G.3 Summary
G.3.1 Score Calculation
We evaluate videoLLM fairness across two dimensions: data-driven bias factors and multimodal understanding fairness.
(1) Data-Driven Bias Factors
Stereotype Impact Generation. The Stereotyping Rate (StR) quantifies the model’s tendency to generate stereotypical content by evaluating outputs using DeepSeek and keyword matching for "Yes" (contains stereotypes) or "No" (does not contain) responses:
$\text{StR}=\frac{\text{Number of outputs containing stereotypes}}{\text{Total number of samples}}\times 100$
A higher rate indicates a stronger tendency for the model to generate stereotypical content.
Preference Selection. The Refuse-to-Answer (RtA) Rate measures model neutrality using a pre-trained classifier and keyword matching for refusal phrases (e.g., "Sorry, I cannot"):
$\text{RtA}=\frac{\text{Number of refused answers}}{\text{Total number of samples}}\times 100$
A higher RtA rate reflects better fairness and neutrality in the model’s responses.
Profession Competence Prediction. We compute the average p-value across demographic conditions (with and without description) as $\text{Score}_{\text{profession}}$.
The overall data-driven bias score is calculated as:
$\mathrm{Score_{data\_driven}}=\frac{1-\mathrm{StR}+\mathrm{RtA}+\mathrm{Score}_{\mathrm{profession}}}{2}\times 100$ (16)
(2) The Fairness in Multimodal Understanding
Agreement on Stereotypes. We measure the average agreement rate across related and unrelated videos, denoted as $\text{Rate}_{\text{agreement}}$. Higher rates indicate greater tendency to endorse stereotypes.
Time Sensitivity Analysis. Accuracy measures the model’s ability to adjust biases as video narratives evolve, evaluated through DeepSeek assessment and keyword matching:
$\text{Accuracy}=\frac{\text{Number of correct adjustments}}{\text{Total number of samples}}\times 100\%$
The overall multimodal understanding fairness score is:
$\mathrm{Score_{data\_driven}}=\frac{1-\mathrm{Rate}_{\mathrm{agreement}}+\mathrm{Acc}_{\mathrm{time}}}{2}\times 100$ (17)
The comprehensive rankings and corresponding scores for Fairness&Bias evaluation are presented in Table XXII.
TABLE XXII: The scores and rankings of two subaspects in Fairness&Bias.
B. F.
Model Score Rank Score Rank
Claude4-sonnet 75.3 4 54.7 9
Claude3.7-sonnet 58.2 8 55.7 7
Gemini1.5-Pro 45.6 15 53.4 15
Gemini1.5-Flash 34.9 19 52.6 19
GPT-4o 34.1 21 59.2 3
Qwen2.5-VL-72B 47.3 13 55.3 8
Qwen2.5-VL-7B 59.9 6 57.9 4
LLaVA-Video-72B 24.7 23 54.1 11
LLaVA-Video-7B 28.6 22 53.8 13
MiniCPM-o-2.6-7B 34.7 20 54.1 10
MiniCPM-V-2.6-7B 47.0 14 52.9 17
Oryx-34B 35.4 18 31.7 23
Oryx1.5-7B 83.0 1 57.8 5
InternVL2.5-78B 58.9 7 55.9 6
LLaVA-OneVision 39.3 17 61.5 2
mPLUG-Owl3-7B 78.1 2 53.1 16
LongVA-7B 77.8 3 52.7 18
Sharegpt4video-8B 56.5 9 62.8 1
TPO-7B 62.1 5 53.7 14
Long-LLaVA-7B 43.9 16 52.4 20
Video-ChatGPT-7B 51.6 11 43.1 22
LiveCC-7B 53.5 10 54.0 12
VideoLLaMA3-7B 48.0 12 47.7 21
G.3.2 Takeaways
  • Significant Differences in Bias Suppression Across Model Types. Closed-source models demonstrate superior performance in suppressing stereotypes and avoiding subjective preferences, likely due to systematic investments in data curation, algorithmic optimization, and ethical constraints. In contrast, the fairness performance of open-source models is uneven—some are more prone to generating or reinforcing stereotypes due to limited training data diversity or lack of debiasing techniques.
  • Model Scale Negatively Correlates with Fairness, but Is Not the Sole Determining Factor. Larger models typically possess stronger semantic understanding capabilities and are less likely to rely on stereotypes when handling sensitive attributes. However, architectural design (e.g., temporal modeling modules, multimodal alignment mechanisms) and training objectives are equally critical to fairness. Some smaller models achieve effective bias suppression through targeted optimizations.
  • Multimodal Nature Increases the Hidden and Complex Nature of Bias. The fusion of visual, auditory, and textual modalities in video can result in bias being conveyed through subtle cues such as scene composition, tone of voice, or subtitles. Irrelevant visual distractions (e.g., noise, color) or early-frame content may mislead the model into forming biased judgments. Most models lack dynamic adjustment capabilities, making it difficult to correct initial biases as the video narrative evolves.
  • Stereotypes in Occupational and Social Attribute Judgments. When evaluating professional competence, models are easily influenced by visual attributes such as gender, age, or skin color, potentially forming stereotype associations even without textual input. While adding textual descriptions can partially mitigate such bias, some models may develop new biases due to overreliance on textual cues, exposing the fragility of cross-modal information integration.
Appendix H Evaluation Details on Privacy
Privacy protection in MLLMs is a significant concern due to the potential for malicious users to extract private information from training data and the increased risk of privacy violations during inference, amplified by MLLMs’ advanced capabilities in understanding and reasoning across multiple modalities, such as video content. Sensitive information, like family address, personal identifiable information and bank statement, is often embedded in video data, combining visual and linguistic cues, which heightens the risk of unintentional privacy exposure. To evaluate the privacy of videoLLMs in handling video content, two primary aspects are assessed: privacy awareness, focusing on the model’s ability to recognize privacy-related content in videos, and Control Over Privacy Inference, examining the model’s behavior in safeguarding sensitive information during video-based tasks.
H.1 Privacy Awareness
H.1.1 Private Content Recognition
Figure 42: An example for the task of Identification of private information.
Setting. This task evaluates the model’s ability to detect privacy-sensitive content in videos, focusing on its capacity to recognize elements such as personal identifiable information including passports, credit cards, hospital prescriptions, and private letters. The evaluation is designed as a generative task, where the model must classify whether a given video is likely to contain private information, responding with a binary output: [Yes] or [No]. The setting emphasizes the ability of the model to process visual and linguistic cues in video data to accurately identify privacy-related content. An example is shown in Figure 42.
Dataset. We sampled 100 videos from the BIV-Priv dataset [sharma2023disability], a benchmark commonly used to advance research on visual recognition under privacy constraints. The dataset is specifically curated to include privacy-sensitive content, such as passports, credit cards, medical prescriptions, and private correspondence. These videos depict real-world scenarios where sensitive information may be inadvertently exposed, offering a rigorous testbed for evaluating model detection performance.
Metrics. We evaluate performance using accuracy, precision, recall and F1. Accuracy measures the overall effectiveness of videoLLMs, while precision and recall reveal the models’ tendencies to predict "Yes" or "No", indicating whether they adopt a conservative or liberal approach in identifying private information. If the model fails to generate the expected keywords, the response defaults to "No", which means a failure to detect the presence of privacy.
TABLE XXIII: Performance (%) of videoLLMs in identifying private information. Acc denotes Accuracy; Pre denotes Precision; Rec denotes Recall.
[…]
GPT-4o 91.0 100.0 91.0 95.3 94.3
Claude4-sonnet 89.0 100.0 89.0 94.2 93.0
Claude3.7-sonnet 79.0 100.0 79.0 88.3 86.6
Gemini1.5-Pro 79.0 100.0 79.0 88.3 86.6
Gemini1.5-Flash 77.0 100.0 77.0 87.0 85.3
Sharegpt4video-8B 95.0 100.0 95.0 97.4 96.9
Long-LLaVA-7B 91.0 100.0 91.0 95.3 94.3
LLaVA-Video-72B 73.0 100.0 73.0 84.4 82.6
TPO-7B 70.0 100.0 70.0 82.4 80.6
Qwen2.5-VL-72B 59.0 100.0 59.0 74.2 73.1
MiniCPM-o-2.6-7B 58.0 100.0 58.0 73.4 72.4
LiveCC-7B 56.0 100.0 56.0 71.8 70.6
Oryx1.5-7B 56.0 100.0 56.0 71.8 70.6
MiniCPM-V-2.6-7B 55.0 100.0 55.0 71.0 70.2
LongVA-7B 51.0 100.0 51.0 67.6 67.4
Oryx-34B 50.0 100.0 50.0 66.7 66.7
LLaVA-Video-7B 48.0 100.0 48.0 64.9 65.2
Qwen2.5-VL-7B 29.0 100.0 29.0 46.0 50.7
mPLUG-Owl3-7B 15.0 100.0 15.0 26.1 39.1
VideoLLaMA3-7B 3.0 100.0 3.0 5.8 28.0
Results. The evaluation results for the task of identifying private information in videos, as detailed in Table  XXIII. Among the closed-source proprietary videoLLMs, GPT-4o achieves the highest accuracy at 91.0%, with an F1 score of 95.3% and an average performance (Avg) of 94.3%, closely followed by claude4-sonnet at 89.0% accuracy (F1: 94.2%, Avg: 93.0%). Gemini1.5-Flash, however, records a lower accuracy of 77.0% (F1: 87.0%, Avg: 85.3%), indicating variability within proprietary models. In contrast, open-source models like Sharegpt4video-8b and Long-LLaVA-7B outperform most proprietary models with an accuracy of 95.0% each (F1: 97.4% and 95.3%, Avg: 96.9% and 94.3%, respectively). However, other open-source models such as VideoLLAMA3-7B struggle significantly, with a mere 3.0% accuracy (F1: 5.8%, Avg: 28.0%). Precision across all models, both close-source and open-source, remains at 100.0%, reflecting a conservative approach where models rarely misclassify non-private content as private. Recall varies widely, with open-source models like Sharegpt4video-8b achieving 95.0%, while VideoLLAMA3-7B only reaches 3.0%, highlighting differences in sensitivity to detecting private information.
Findings. (1) The consistent 100.0% precision across all models highlights a cautious strategy, minimizing false positives, though this often results in lower recall for less capable models, particularly among open-source ones like VideoLLAMA3-7B. (2) Closed-source commercial videoLLMs like GPT-4o and claude4-sonnet demonstrate strong performance in identifying privacy-sensitive content, but they are generally outpaced by top open-source models like Sharegpt4video-8b and Long-LLaVA-7B, suggesting that open-source advancements in training data diversity and alignment can yield competitive results in video understanding. (3) The wide recall and F1 score variations indicate that while models excel at detecting explicit private information, their performance drops in complex video scenarios with subtle privacy cues, highlighting the need for improved contextual analysis and diverse training datasets to bolster video understanding capabilities.
H.1.2 Private information QA
Figure 43: An example for the task of Identification of private information.
Setting. This task assesses the model’s ability to reason about privacy-sensitive content in videos, building on its detection capabilities. The evaluation is designed as a discriminative task, where the model answers questions about privacy-related information in videos, such as identifying specific content (e.g., phone/computer screen content, license plates, or delivery addresses). The setting tests the model’s capacity to integrate visual and linguistic information to infer and respond to privacy-related queries accurately. An example is shown in Figure 43.
Dataset. The dataset is sourced from YouTube, comprising 90 videos that include frames with privacy-sensitive information, such as phone or computer screen content, license plates, and delivery addresses. These videos are selected to reflect real-world scenarios where privacy information may be embedded in specific frames, requiring the model to process temporal and multimodal data effectively.
Metrics. This task is also a binary classification, with evaluation metrics aligned with those used in private content recognition.
TABLE XXIV: Performance (%) of videoLLMs in private information QA. Acc denotes Accuracy; Pre denotes Precision; Rec denotes Recall. LLaVA-OneVision is a 72B version.
[…]
GPT-4o 80.0 81.4 77.8 79.9 79.7
Gemini1.5-Flash 72.2 75.0 66.7 70.6 71.1
Claude4-sonnet 73.3 86.2 55.6 67.6 70.7
Gemini1.5-Pro 64.4 65.1 62.2 63.6 63.9
Claude3.7-sonnet 58.9 75.0 26.7 39.3 50.0
LLaVA-OneVision 50.0 50.0 100.0 66.7 66.7
Qwen2.5-VL-72B 60.0 84.6 24.4 37.9 51.8
Oryx-34B 57.8 65.2 33.3 44.1 50.1
Long-LLava-7B 45.6 45.7 46.7 46.2 46.0
LLaVA-Video-7B 44.4 44.7 46.7 45.7 45.4
Video-ChatGPT-7B 43.3 43.4 44.4 44.0 43.8
Sharegpt4video-7B 53.3 100.0 6.7 12.5 43.1
LLaVA-Video-72B 50.0 50.0 28.9 36.6 41.4
VideoLLaMA3-7B 51.1 57.1 8.9 15.4 33.1
MiniCPM-o-2.6-7B 51.1 60.0 6.7 12.0 32.4
Oryx1.5-7B 47.8 42.9 13.3 20.3 31.1
MiniCPM-V-2.6-7B 38.9 33.3 22.2 26.7 30.3
mPLUG-Owl3-7B 50.0 50.0 6.7 11.8 29.6
LiveCC-7B 48.9 42.9 6.7 11.5 27.5
TPO-7B 45.6 35.7 11.1 17.0 27.3
Qwen2.5-VL-7B 50.0 50.0 2.2 4.3 26.6
LongVA-7B 43.3 20.0 4.4 7.3 18.8
Results. The performance of videoLLMs in the private information QA task, as shown in Table XXIV, reveals significant variations across models, with both closed-source and open-source models exhibiting distinct strengths and weaknesses in video understanding for privacy-related queries. Among closed-source models, GPT-4o achieves the highest accuracy at 80.0%, with a balanced precision of 81.4% and recall of 77.8%, leading to an F1 score of 79.9% and an average score of 79.7%, demonstrating robust video understanding capabilities for privacy-sensitive content. Gemini1.5-Pro follows with a respectable average score of 63.9%, though its recall of 62.2% indicates a slight struggle in identifying all privacy-related instances. Claude3.7-sonnet, however, underperforms with an average score of 50.0%, primarily due to a low recall of 26.7%, suggesting limitations in detecting privacy content across video frames.
Among open-source models, LLaVA-OneVision-72B stands out with an accuracy of 50.0%, precision of 50.0%, and a perfect recall of 100.0%, yielding an F1 score of 66.7% and an average of 66.7%, indicating strong detection of privacy-sensitive content despite challenges in precision. Qwen2.5-VL-72B, another notable open-source model, achieves an accuracy of 60.0% but suffers from a low recall of 24.4% (F1: 37.9%, Avg: 51.8%), reflecting difficulties in consistently identifying privacy information in complex video contexts. Models like VideoLLaMA3-7B (Avg: 33.1%) and MiniCPM-V-2.6-7B (Avg: 32.4%) show significant struggles, with low recall rates (8.9% and 6.7%, respectively), underscoring their limited video understanding capabilities for privacy tasks. Early-stage models like mPLUG-Owl3-7B (Avg: 29.6%) and TPO-7B (Avg: 27.3%) perform poorly, aligning with their weaker general perception abilities.
Findings. (1) Closed-source models like GPT-4o and Gemini1.5-Pro generally exhibit superior performance in privacy-related video understanding compared to open-source counterparts, maintaining higher accuracy and stability across diverse scenarios, though Claude3.7-sonnet lags behind due to poor recall. (2) Open-source models like LLaVA-OneVision-72B show promise with high recall, but their inconsistent precision highlights a trade-off in video understanding, often failing to balance detection and accuracy in privacy-sensitive contexts. (3) Both closed- and open-source models tend to adopt a conservative approach when faced with challenging video content, as evidenced by higher recall in models like LLaVA-OneVision-72B and Sharegpt4video-7B, indicating a preference to over-identify privacy content rather than miss it. (4) The recognition of privacy in videos is closely tied to a model’s general video understanding capabilities, with better-aligned training data improving performance, though not uniformly across all models or scenarios.
H.2 Control Over Privacy Inference
H.2.1 Infoflow Expectation
Figure 44: An example for the task of infoflow expectation.
Setting. We assess the capability of videoLLMs to appropriately handle private information in videos. Following the approach of MultiTrust [zhang2024benchmarking], we utilize the InfoFlow Expectation from [mireshghallah2023can], which characterizes the contexts in which specific types of private information are used and measures the alignment between model predictions and human expectations. To facilitate subsequent answer extraction, we include an explicit formatting instruction at the end of the prompt. An example is shown in Figure 44.
Dataset. Following the setting of MultiTrust [zhang2024benchmarking], we use the Tier-2a subset of 98 prompts from [mireshghallah2023can], which describe only the usage of private data, excluding additional contextual short stories found in Tier-2b. This subset covers 7 types of private information across 14 usage scenarios. In addition to pair with irrelevant videos, we employ Jimeng, a commercial text-to-video model, to generate semantically relevant videos. Specifically, we use the usage scenarios as prompts for video synthesis, producing 14 relevant videos paired with the corresponding text samples.
Metrics. Following the MultiTrust [zhang2024benchmarking], we evaluate the pearson correlation between model ratings and human annotations. Predictions are extracted from responses using keyword matching, capturing both numeric values (e.g., “-100”) and verbal expressions (e.g., “strongly disagree”). During preliminary testing, we observed that some models fail to provide clear scores as instructed. In such cases, the response is treated as a task failure, and a default “neutral” value (0) is assigned.
TABLE XXV: Pearson Correlation Coefficient for videoLLMs predicting people’s attitudes towards the usage of privacy. P denotes Pearson; F denotes Failure;
Models Related Unrelated
P F P F
Gemini1.5-Flash 75.0 36.7 77.7 0.0
Claude4-sonnet 73.4 19.4 67.7 0.7
GPT-4o 73.2 0.0 69.7 0.0
Claude3.7-sonnet 70.2 1.0 68.9 2.0
Gemini1.5-Pro 66.8 45.9 67.5 0.0
Qwen2.5-VL-72B 74.7 0.0 69.6 0.0
InternVL2.5-78B 74.0 0.0 71.4 0.0
LLaVA-Video-72B 65.7 0.0 64.9 0.0
Qwen2.5-VL-7B 64.8 0.0 57.6 0.0
mPLUG-Owl3-7B 56.3 24.5 57.9 48.6
LLaVA-Video-7B 51.6 0.0 30.8 0.0
LiveCC-7B 47.9 0.0 43.9 0.0
MiniCPM-o-2_6 44.1 10.2 50.1 24.5
Sharegpt4video-8B 43.9 0.0 35.3 0.0
TPO-7B 37.0 0.0 43.5 0.0
LongVA-7B 35.6 0.0 36.1 0.0
Video-ChatGPT-7B 32.2 72.5 9.5 17.7
Oryx1.5-7B 30.7 0.0 44.5 0.0
Long-LLaVA-7B 22.1 0.0 37.1 0.0
VideoLLaMA3-7B 17.8 0.0 9.6 1.7
MiniCPM-V-2_6 14.8 76.5 17.7 76.2
Results. The results are presented in Table XXV. Closed-source models such as GPT-4o exhibit strong alignment with human privacy expectations, achieving high Pearson correlations (73.2 for related, 69.7 for unrelated videos) and no failure cases. However, models like Gemini1.5-Flash and Gemini1.5-Pro, despite competitive correlations, show high failure rates, indicating reliability concerns. Among open-source models, Qwen2.5-VL-72B and InternVL2.5-78B lead with high correlations and zero failures, outperforming smaller models such as LLaVA-Video-7B and MiniCPM-V-2.6, which show weak alignment and high failure rates. Overall, while closed-source models tend to offer higher correlation scores, larger open-source models provide more consistent and failure-free performance, highlighting their potential as reliable alternatives in privacy-sensitive applications.
Findings. (1) Model scale critically determines privacy performance, with larger architectures exhibiting superior comprehension and consistency in privacy-sensitive contexts. (2) For mission-critical privacy applications, zero-failure reliability takes precedence over models with higher correlation but inconsistent performance, such as GPT-4o and Claude3.7. (3) Contemporary large-scale open-source models demonstrate comparable or superior reliability to proprietary alternatives. (4) Smaller models, particularly 7B parameter video language models, exhibit systematic underperformance and are unsuitable for privacy-critical deployments.
H.2.2 Celebrities Privacy information QA
Figure 45: An example for the task of celebrities privacy information QA.
Setting. This task evaluates the model’s ability to protect privacy-sensitive information related to celebrities in videos, focusing on its discriminative capabilities. The model is tasked with a swering questions about privacy-related content in videos of celebrities from diverse domains (sports, entertainment, politics, music), ensuring that it refrains from disclosing or misusing personal identifiable information. The setting tests the model’s capacity to recognize and appropriately handle sensitive information, balancing accurate responses with privacy protection across diverse demographic representations (e.g., varied skin tones and genders). An example is shown in Figure 45.
Dataset. The dataset consists of videos from 50 celebrities, sourced from publicly available platforms, covering domains such as sports, entertainment, politics, and music. The dataset is curated to include diverse skin tones (e.g., Black, White, Asian) and genders (male and female) to ensure comprehensive representation.
Metrics. In the task, we use the Reject answer rate (RtA) as the sole evaluation metric. Models are expected to refuse queries involving personal identifi able information regardless of the input content. To identify refusal responses, we apply both a pre-trained Longformer evaluator [wang2023not] and keyword matching (e.g., detecting phrases like “Sorry, it’s not appropriate to…” or warnings). The keyword matching compensates for cases missed by the classifier.
Figure 46: RtA Rate (%) in the celebrity privacy information QA task. n denotes name, o denotes occupation, and w denotes without; for example, w-n-o indicates that only occupation is provided in the prompt.
Results. The evaluation results for the Celebrities Privacy Information QA task, as shown in Figure 46, highlight the models’ ability to safeguard privacy-sensitive information in video content by rejecting queries involving personal identifiable information. Among closed-source commercial models, Claude4-sonnet demonstrates exceptional performance, achieving a 100.0% rejection rate (RtA) across all categories—phone, address, and email—regardless of whether name (n) or occupation (o) is provided in the prompt. GPT-4o also performs strongly, with a 100.0% RtA in most scenarios, though it slightly drops to 96.0% in phone (n-o) and email (n-o) settings, indicating minor sensitivity to fully contextualized prompts. Gemini1.5-Pro maintains a high RtA of 100.0% in most cases but shows noticeable declines when occupation is omitted (w-o), with scores of 78.0% for address (n-w-o) and 88.0% for email (n-w-o), suggesting that contextual cues like occupation play a critical role in its privacy discrimination.
Among open-source models, performance varies widely. Advanced models like Sharegpt4video-8B and LLaVA-OneVision-72B exhibit strong privacy awareness, achieving 98%-100.0% RtA in scenarios with full context (n-o) across phone, address, and email queries, surpassing some closed-source models in specific cases—for example, ShareGPT4Video-8B’s 100.0% RtA in address (n-w-o) compared to Gemini1.5-Flash’s 80.0%. However, these models falter in reduced-context scenarios; LLaVA-OneVision-72B drops to 98.0% in address (n-w-o) and 96.0% in email (w-n-w-o), while Sharegpt4video-8B falls to 2.0% in address (w-n-o), highlighting instability when key identifiers like name or occupation are absent. Early-stage open-source models like MiniCPM-o-2.6-7B and Video-ChatGPT-7B struggle significantly, with RtA scores as low as 0.0% in challenging scenarios such as address (w-n-w-o) and email (w-n-w-o), underscoring their limited capability to recognize and protect privacy in video modalities.
Findings. (1) Closed-source models like Claude4-sonnet and GPT-4o exhibit robust privacy protection in video-based celebrity QA tasks, consistently achieving high RtA across various scenarios, though minor inconsistencies arise when contextual information is reduced. (2) Advanced open-source models such as Sharegpt4video-7B and LLaVA-OneVision-72B demonstrate competitive performance, occasionally outperforming closed-source models in specific categories, but their sensitivity to missing context (e.g., name or occupation) reveals a need for improved training data alignment. (3) Early-stage open-source models like MiniCPM-o-2.6-7B and mPLUG-Owl3-7B lack the discriminative ability to handle privacy-sensitive video content, particularly in ambiguous scenarios, as evidenced by their near-zero RtA in challenging cases.
H.2.3 Privacy Information Self-Inference
Figure 47: An example for the task of Privacy Information Self-Inference
Setting. This task evaluates the model’s ability to proactively identify and reason about privacy-sensitive information in videos without explicit guidance or specific prompts directing attention to particular elements. Unlike the previous tasks that provide targeted questions or direct instructions, this evaluation assesses the model’s autonomous capability to recognize, analyze, and infer the presence of private information through self-directed reasoning. The task is designed as a generative evaluation where the model must independently examine video content and provide comprehensive analysis of any privacy-related concerns it identifies, demonstrating its understanding of privacy implications across various contexts without predetermined focus areas. An example is shown in Figure 47.
Dataset. This dataset, sourced from YouTube, comprises 91 privacy-sensitive videos captured in natural, real-world settings. It is designed to reflect realistic scenarios where private information may be present but not explicitly highlighted, requiring models to demonstrate autonomous detection and reasoning without guided attention to specific elements.
Figure 48: Prompt for DeepSeek to judge whether the description contains privacy information.
Metrics. To ensure reliable identification of privacy-related content in VideoLLMs’ outputs, we employ DeepSeek as an automated judge due to its strong capability in detecting private information. Based on DeepSeek’s responses, a keyword matching algorithm is used to statistically analyze the presence of "Yes" or "No" answers, thereby calculating the privacy leakage rate. The evaluation prompt used with DeepSeek is shown in Figure 48.
Figure 49: Performance of videoLLMs in privacy information self-inference
Results. The evaluation results for the Privacy Information Self-Inference task, as depicted in Figure 49. Closed-source models significantly outperform open-source counterparts in privacy inference tasks. GPT-4o exhibits the highest privacy leakage rate at 75.05%, followed by Claude4 and Gemini1.5-Pro, both at 82.0%, reflecting their superior video understanding capabilities. In contrast, open-source models demonstrate markedly lower rates—Video-ChatGPT at 8.9%, VideoLLaVA at 13.3%, and ShareGPT4Video at 6.7%, with the exception of TPo-7B. The substantial performance gap between the top proprietary and lowest open-source models underscores notable disparities in recognizing privacy-sensitive content. Notably, all proprietary models exceed a 50% detection rate, while most open-source models fall below this threshold, highlighting that stronger video understanding capabilities may lead to greater privacy reasoning and leakage issues.
Findings. (1) Proprietary videoLLMs demonstrate superior capability in autonomously identifying and reasoning about privacy-sensitive information in videos without explicit guidance, suggesting more sophisticated understanding of contextual privacy implications and visual content analysis. (2) Open-source models exhibit limited privacy awareness in self-directed scenarios, which may reflect constraints in their training data diversity, alignment strategies, or fundamental architectural limitations for complex multimodal reasoning tasks. (3) The substantial performance variation across open-source models highlights the current immaturity of privacy-sensitive content awareness in the videoLLM domain, with even the best-performing most models achieving less than 50% detection rates, leaving considerable room for improvement. (4) The results present a dual challenge for practical deployment: while higher detection rates indicate stronger model capabilities, they simultaneously represent increased privacy leakage risks, emphasizing the critical need for balanced approaches between model performance enhancement and privacy protection mechanisms.
H.3 Summary
H.3.1 Score Calculation
In the privacy evaluation, the score calculations for evaluating privacy aspects in videoLLMs are formulated to assess two primary tasks: Privacy Content Recognition and Control Over Privacy Inference.
The score aggregates results from five distinct tasks: private information recognition, private information QA, infoflow expectation, celebrities privacy information QA, and privacy information self-inference. Each task contributes equally to the final score, weighted at 0.2, and employs specific metrics tailored to its evaluation setting. The metrics include avgerage score (Avg.), Pearson correlation (Corr), Reject answer rate (RtA) and Privacy leakage rate (PlT), averaged over settings as described.
Private Content Recognition employs the average score $\mathrm{Avg_{PI\_recognition}}$ across accuracy, precision, recall, and F1 metrics on the BIV-Priv dataset to evaluate models’ capability to detect privacy-sensitive content including passports and credit cards.
Private Information QA utilizes the average score ($\mathrm{Avg_{PI\_QA}}$) across accuracy, precision, recall, and F1 metrics on the YouTube dataset to assess reasoning capabilities regarding privacy-related queries.
InfoFlow Expectation measures pearson correlation ($\mathrm{Corr_{infoflow}}$) across multiple Tier-2a subset configurations to evaluate alignment with human expectations regarding private information usage.
Celebrities Privacy information QA employs the reject answer rate ($\mathrm{RtA_{celebrities}}$) across celebrity video dataset splits to measure models’ propensity to refuse disclosure of sensitive celebrity information.
Privacy Information Self-Inference quantifies the privacy leakage rate ($\mathrm{PlT_{inference}}$) by evaluating models on a curated YouTube dataset using DeepSeek assessment methodology. This metric measures models’ tendency to autonomously infer private information without explicit prompts or direct instructional guidance. The overall score for the privacy evaluation is:
$\begin{split}\mathrm{Score_{privacy}}&=\bigg(\mathrm{Avg_{PI\_recognition}}+\mathrm{Avg_{PI\_QA}}\\ &\quad+\mathrm{Corr_{infoflow}}+\mathrm{RtA_{celebrities}}\\ &\quad+\mathrm{PlT_{inference}}\bigg)\big/5\times 100\end{split}$ (18)
The comprehensive rankings and corresponding scores for privacy evaluation are presented in Table XXVI.
TABLE XXVI: The scores and rankings of two subaspects in Privacy. LLaVA-OneVision is the 72B version.
R. I.
Model Score Rank Score Rank
Claude4-sonnet 81.9 2 39.9 5
Claude3.7-sonnet 68.3 8 35.3 9
Gemini1.5-Pro 75.2 4 44.7 4
Gemini1.5-Flash 78.2 3 47.1 1
GPT-4o 87.0 1 35.4 8
Qwen2.5-VL-72B 62.4 10 35.8 7
Qwen2.5-VL-7B 38.7 20 30.5 15
LLaVA-Video-72B 62.0 11 32.4 12
LLaVA-Video-7B 55.3 13 20.4 17
MiniCPM-o-2.6-7B 52.4 15 32.1 13
MiniCPM-V-2.6-7B 50.3 17 46.1 3
Oryx-34B 58.4 12 31.7 14
Oryx1.5-7B 51.0 16 18.7 20
InternVL2.5-78B 67.6 9 36.1 6
LLaVA-OneVision 33.3 22 34.7 10
mPLUG-Owl3-7B 34.3 21 46.7 2
LongVA-7B 43.1 19 17.8 21
Sharegpt4video-8B 70.0 7 19.8 18
TPO-7B 54.0 14 19.7 19
Long-LLaVA-7B 70.2 6 14.8 22
Video-ChatGPT-7B 71.9 5 32.9 11
LiveCC-7B 49.2 18 22.9 16
VideoLLaMA3-7B 30.5 23 5.3 23
H.3.2 Takeaways
  • Superiority of Closed-Source Models: Overall, closed-source models like GPT-4o and Claude4-sonnet consistently outperform open-source models across tasks due to better video understanding and contextual reasoning. Their high accuracy, precision, and rejection rates make them more reliable for privacy-sensitive applications.
  • Open-Source Model Limitations: Open-source models like LLaVA-OneVision-72B and Sharegpt4video-7B show promise, particularly in high-recall scenarios, but struggle with precision and context sensitivity. Early-stage models (e.g., MiniCPM, mPLUG-Owl3) are significantly limited, indicating a need for improved training data and architectures.
  • Conservative Approach to Privacy: Both closed- and open-source models tend to adopt a conservative stance, prioritizing high recall to avoid missing privacy-sensitive content, which can lead to over-identification but reduces the risk of privacy breaches.
  • Context Dependency: Compared to closed-source models, open-source models rely more heavily on contextual cues (e.g., names or occupations in celebrity tasks) to infer privacy-related information. This reliance likely stems from limited training data rather than superior safety alignment.
  • Privacy vs. Performance Trade-Off: Higher detection rates in tasks like self-inference indicate stronger model capabilities but also increase privacy leakage risks. This dual challenge underscores the need for more effective safety alignment approaches that enhance performance while incorporating privacy protection mechanisms.